### v1

# RealitySummary: On-Demand Mixed Reality Document Enhancement using Large Language Models

Abstract.
We introduce RealitySummary, a mixed reality reading assistant that can enhance any printed or digital document using on-demand text extraction, summarization, and augmentation. While augmented reading tools promise to enhance physical reading experiences with overlaid digital content, prior systems have typically required pre-processed documents, which limits their generalizability and real-world use cases. In this paper, we explore on-demand document augmentation by leveraging large language models. To understand generalizable techniques for diverse documents, we first conducted an exploratory design study which identified five categories of document enhancements (summarization, augmentation, navigation, comparison, and extraction). Based on this, we developed a proof-of-concept system that can automatically extract and summarize text using Google Cloud OCR and GPT-4, then embed information around documents using a Microsoft Hololens 2 and Apple Vision Pro. We demonstrate real-time examples of six specific document augmentations: 1) summaries, 2) comparison tables, 3) timelines, 4) keyword lists, 5) summary highlighting, and 6) information cards. Results from a usability study (N=12) and in-the-wild study (N=11) highlight the potential benefits of on-demand MR document enhancement and opportunities for future research.
Mixed Reality; Natural Language Processing; Large Language Models; Augmented Reading
journalyear: 2024copyright: acmlicensedconference: ; May, 2024; $c$s: Human-centered computing Mixed / augmented reality
Figure 1. RealitySummary provides an on-demand reading assistant with embedded mixed reality (MR) text summaries and visual aids. Using markerless tracking, OCR, and GPT-4, the system can provide live summaries, keyword lists, timelines, and text highlighting based on the real-time extraction, analysis, and visualization of any physical or digital document. (Simulated high-resolution illustration based on real Unity assets associated with content in the article shown.)
1. Introduction
“We live in a complex world, filled with myriad objects, tools, toys, and people. [..] Yet, for the most part, our computing takes place sitting in front of, and staring at, a single glowing screen. […] How can we escape from the computer screen and bring these two worlds together?” — Pierre Wellner (Wellner, 1991)
In his pioneering work (Wellner, 1991), Pierre Wellner presented the first demonstration of AR-enhanced documents, arguing that the future of computer interfaces should augment our physical environment, rather than confining us to a flat rectangle screen. This fusion harnessed the dynamic capabilities of digital content while retaining the tangible benefits of physical documents. With recent advances in mixed reality technologies like the Microsoft Hololens and Apple Vision Pro, we are getting closer to realizing his vision where we can naturally augment surrounding paper, books, and documents, as well as content on screens, to enhance our reading and learning experiences (Li et al., 2019; Rajaram and Nebeling, 2022; Chen et al., 2020).
However, most prior systems (Billinghurst et al., 2001; Zhao et al., 2014; Li et al., 2019) have typically required documents to be pre-processed with pre-generated content. The lack of on-demand and real-time content generation for non-prepared documents hinders their generalizability and practicality in real-world applications beyond research labs.
This paper introduces RealitySummary, a mixed reality reading assistant that enhances any printed or digital document using on-demand text extraction, summarization, and augmentation. Our on-demand document augmentation leverages optical character recognition (OCR) and a large language model (LLM) to enhance any given document without prior preparation. To design our system, we first conducted a formative design workshop with six participants to identify common document enhancement techniques across diverse document types. We identified 20 unique techniques across five categories of document enhancement: summarization, augmentation, navigation, comparison, and extraction. Insights from the study led us to incorporate the top six techniques into our system: 1) summaries, 2) comparison tables, 3) timelines, 4) keyword lists, 5) summary highlighting, and 6) information cards.
Based on this insight, we developed our system for the Microsoft Hololens 2 and Apple Vision Pro. The system first captures the document, and then extracts text and the document image using Google’s Cloud OCR API and OpenCV. Given this extracted text, the system generates different augmented content and extracts related images through GPT-4 and Google Image Search. The system also tracks the document based on the extracted document image, rendering the generated content around the document as a spatial 2D canvas. Furthermore, we integrated a speech input using Google’s speech-to-text API, enabling users to activate different summarization features or ask questions related to the current content through voice commands.
To evaluate our system, we conducted two user studies: 1) a usability study with twelve participants, and 2) an in-the-wild study with eleven participants. Both studies highlighted the unique benefits of contextual assistant for diverse set of reading materials without prior preparation. Participants valued the always-on feature for enabling implicit interactions and appreciated the mixed reality interface for facilitating spatial and tangible exploration, significantly enhancing user engagement and the applicability of AI-generated content across various scenarios. Based on the study insights, we also discuss potential risks and future research opportunities of AI-powered mixed reality reading assistants.
Finally, our paper contributes:
  1. (1)
    RealitySummary, a mixed reality reading assistant that uses a large language model to provide on-demand content summarization and augmentation.
  2. (2)
    A set of design features and summarization techniques for mixed reality document enhancement derived from a formative design workshop (N=6).
  3. (3)
    Insights and findings from our usability study (N=12) and in-the-wild study (N=11) that highlight the potential benefits of on-demand mixed reality document enhancements.
2. Related Work
2.1. Reading Augmentation
2.1.1. Document Enhancement
Human-computer interaction researchers have explored ways to augment physical documents with AR-enhanced content. Examples include HoloDoc (Li et al., 2019), Replicate and Resuse (Gupta et al., 2020), and PaperTrail (Rajaram and Nebeling, 2022), all of which laid the groundwork for augmenting physical paper by superimposing dynamic virtual content onto printed documents. Prior works have demonstrated these concepts on different platforms. For example, Pacer (Liao et al., 2010), MagPad (Xu et al., 2015), Tecahable Reality (Monteiro et al., 2023), and Dually Noted (Qian et al., 2022) use mobile AR for text augmentation, whereas Replicate and Resuse (Gupta et al., 2020) and Wiki-TUI (Wu et al., 2007) leverage head-mounted displays to overlay information. Other examples such as DigitalDesk (Wellner, 1991), DocuDesk (Everitt et al., 2008), Affinity Lens (Subramonyam et al., 2019), Matulic et al. (Matulic and Norrie, 2013; Matulic et al., 2013), and QOOK (Zhao et al., 2014) use projection mapping for document augmentation. At their core, such augmented reading research (Kirner et al., 2012; Billinghurst et al., 2001; Hidayat et al., 2020; Grasset et al., 2008) aims to improve cognitive load (Cheng, 2017), knowledge accumulation (Dünser et al., 2012; Zhang et al., 2019), and spatial visualization (Shelton and Hedley, 2004), given a long-standing consensus that readers prefer physical paper over screen-based reading  (Tashman and Edwards, 2011a; Sellen and Harper, 2002).
However, none of the existing systems have achieved an on-demand document enhancement. Instead, documents required pre-processing with content manually prepared before being embedded in AR. For example, HoloDoc (Li et al., 2019), PaperTrail (Rajaram and Nebeling, 2022), and Affinity Lens (Subramonyam et al., 2019) did not extract text directly from the given document; instead, content was prepared in advance and loaded based on attached fiducial markers. As a result, these systems are not fully adaptable and deployable to real-world scenarios. Some preliminary works have explored on-demand text analysis using machine learning, such as Dually Noted (Qian et al., 2022) analyzing document structure in real-time, Augmented Math (Chulpongsatorn et al., 2023) extracting graphs and math equations, and SOCRAR (Strecker et al., 2022) using OCR to extract key information. However, none have yet achieved comprehensive and general-purpose document enhancement. In this paper, we contribute to the first demonstration of on-demand document enhancement by leveraging large language models. The goal of this paper is not only to demonstrate its potential but also to uncover its limitations and challenges from a user experience perspective through user evaluations.
2.1.2. Screen-Based Reading Support Tools
Outside of the context of mixed reality, a wide range of projects and systems have explored reading support tools for screen-based interfaces. Examples including LiquidText (Tashman and Edwards, 2011b), texSketch (Subramonyam et al., 2020), ScholarPhi (Head et al., 2021), Chameleon (Masson et al., 2020), and Marvista (Chen et al., 2022) developed on-screen reading support tools. Similar to our work, Marvista (Chen et al., 2022) helps users read online documents by providing reading aids including summaries, contextual prompt questions, and reading metrics for active reading. Other works studied active and close reading support like Metatation (Mehta et al., 2017) for closed reading, XLibris (Price et al., 1998) with free form annotation, Matulic and Norrie (Matulic and Norrie, 2012) to support navigation, and GatherReader (Hinckley et al., 2012) with information gathering. Other works also support literature review and reference search for science papers (Palani et al., 2023; Kang et al., 2023; Chang et al., 2023). Lastly, works like Explorable Explanations (Victor, 2011) and Potluck (Litt et al., 2022) try to enhance static documents with interactive content to improve reading comprehension through exploration. While these past works are limited to screen-based interfaces, this paper explores the augmented reading design for AR-based interfaces.
2.2. Content Summarization
2.2.1. Automatic Summarization
In the field of natural language processing, several works have employed automatic summarization to support readers (El-Kassas et al., 2021; Zhang et al., 2023). These approaches provide various summarization methods, including abstractive and extractive summaries generated from document photos (Benharrak et al., 2022), real-time summarization of user writing for quick text iterations (Dang et al., 2022), enhancing reading comprehension through automatic summarization (Chen et al., 2022), and responsive text summarization for adapting document summaries to screen sizes ranging from large displays to small watch faces (Leiva, 2018). With the recent advances in large language models like GPT-4, such automatic summarization tasks have become more accurate, robust, and accessible. For example, Goyal et al. (Goyal et al., 2022) showed that GPT-3 summaries were preferred by humans over those generated by fine-tuned models. Various types of summarization have been explored using GPT-3, such as opinion summarization (Bhaskar et al., 2022), news summarization (Goyal et al., 2022), and medical dialogue summarization (Chintagunta et al., 2021) while others have built custom frameworks for feeding long texts into GPT-3 to elicit more finely-controlled summarizations (Payne, 2022).
2.2.2. Interactive Summarization.
Instead of automatically summarizing the content, the researchers have also explored interactive summarization based on the context and user’s needs. For example, Hoeve et al. (ter Hoeve et al., 2020) studied Q&A interactions to assist users with a document-centric question-answering paradigm, finding that information-seeking was a common query. Semantic Reactor (Google, [n.d.]), developed by Google, enables users to ask questions about documents in natural language by fine-tuning BERT (Devlin et al., 2019) using the SQuAD dataset (Rajpurkar et al., 2018). VERSE (Vtyurina et al., 2019) is another Q&A system that allows users to ask specific types of questions about a document but is limited by its inability to summarize. Hoeve et al. (ter Hoeve et al., 2020) identified various question types asked by participants regarding documents, with factual/summarization questions being the most frequent. Some question types include document-related, factoid, mechanical (answerable by rule-based methods), factual, yes/no, navigational, and summary questions. Voice assistants like Cortana, Alexa, Siri, VERSE (Vtyurina et al., 2019), Firefox Voice (Cambre et al., 2021), and Pushpak (Holani et al., 2019) offer efficient query modalities, as speech input has been found to be significantly faster than manual input, especially for mixed reality devices (Ruan et al., 2016; Adhikary and Vertanen, 2021). Our goal is to build on top of this prior work, developing and exploring automatic and interactive summarization techniques for on-demand and real-time summaries embedded in documents.
3. Exploring Document Enhancement Techniques for Augmented Reading
To explore the possible design of an AI-assisted augmented reading, we conducted formative design workshops with six participants (4 males, 2 females), who had various design backgrounds as undergraduate and graduate HCI researchers.
3.1. Method
We asked participants to complete a four-part design task where they read and graphically annotated both their own document (Table 1) and a Wikipedia article [^1] Source: https://en.wikipedia.org/wiki/Russian_invasion_of_Ukraine. After reading the documents, participants brainstormed possible ideas for AI-powered augmentations to enhance their readability. In the first design task, participants annotated a physical printout of their own article, sketching ideas for possible enhancements, and then did the same with a printed copy of the Wikipedia article. In the second design task, we gave participants digital copies of the two documents on a virtual whiteboard (Miro) and asked them to design a second set of document enhancements with the assistance of ChatGPT. Each task took approximately 15 minutes, and each session of the entire study took approximately 60 minutes.
Document Types Topic Pages Words
P1 Academic Paper (Horie et al., 2021) HCI 8 2400
P2 Book Chapter (Robinson and Rodrigues, 2022) Religion 6 1800
P3 Academic Paper (Peirson-Smith, 2013) Social Science 4 2000
P4 News Article (NYT, [n.d.]) Non-Fiction 8 1600
P5 Laboratory protocol Engineering 4 1750
P6 Academic Paper (Lin et al., 2018) HCI 4 2200
All Wikipedia Article (wik, [n.d.]) World Affairs 6 4750
Table 1. Documents used by participants in the workshop.
Before commencing the study, we gave participants a verbal walk-through of the study’s purpose and provided a one-page reference containing sample visualizations. We also provided a set of sample prompts for ChatGPT to demonstrate the AI’s summarization and enhancement capabilities. The study followed a talk-aloud protocol and the authors took detailed notes of responses and observations. At the study’s conclusion, the authors collected all sketches authored by the participants.
Upon study completion, all designs elicited from the participants were collated in a single Miro board. Four authors independently categorized and coded those results to identify common themes. We identified twenty distinct types of document enhancements, each with its specific usage, purpose, popularity, and the context in which it was used. Next, all authors discussed the initial design dimensions to refine the consistency and comprehensiveness of the categorization. Finally, four authors collaboratively re-coded the examples to produce a final set of design dimensions.
Figure 2. Document enhancements produced by six participants in our formative design workshop. Highlights Refer to caption indicated the document enhancements implemented in our system.
3.2. Elicited Designs and Findings
We identified twenty unique classes of document enhancements, with each participant generating designs from on average three classes (SD = 1.41). We then identify five high-level categories, which include summarize, compare, augment, extract, and navigate. Figure 2 illustrates the summary of our study results. In this section, we describe each category in more detail.
Category 1. Summarize
One of the most common themes was to expedite the reading process by condensing lengthy text. All participants suggested text-based summaries with different scopes (document-level, page-level, or paragraph-level). Some participants also explored personalized summaries, such as “summarize this for a 5-year-old”, which help understand unfamiliar concepts and fit the unique comprehension levels and interests. Furthermore, many participants recommended the use of text summaries as a reference tool, enabling them to quickly review content from different parts of the document without losing focus on the current page.
Category 2: Compare
The most unique approach was the idea to dynamically change content representation by leveraging AI. For instance, participants envisioned transforming text and numerical data into more visual and spatial formats, such as tables, mind maps, or graphs. These alternative representations enable users to visually compare and analyze elements that would be less apparent in text. Moreover, these representations could be designed to be fluid and adaptive; for instance, tables or graphs could dynamically update, filter, or highlight based on user queries. Another idea was summarizing all of the chronological information into timelines. Such timelines could also automatically update by adding new content as the user reads through the document. This approach represents a unique design for AI-enabled augmentation, which emphasizes reshaping existing content into more comprehensible and comparable formats, offering an alternative method for active reading.
Category 3. Augment
All participants showed interest in enhancing their content with external data, such as maps and images. For instance, one participant envisioned a feature similar to Apple’s keyword lookup for quick reference to unfamiliar words. Such augmented content could come either from external sources (like images and maps), user-generated content (labels and sticky notes), or AI-generated content (converting 2D figures into 3D models). While such ideas have been explored in prior works (Li et al., 2019; Qian et al., 2022), participants also came up with unique AI-driven enhancements beyond simple augmentation. For instance, participants envisioned that by visualizing the importance, relevance, or number of citations for each reference in academic papers, they could quickly determine which paper to read next. Another unique concept involved synthesizing external information with the comparison feature. For example, when looking up the president of Ukraine, participants suggested not just showing a biography image and profile but also visualizing the organizational relationships through a network diagram with other individuals on a page by combining both external information and content on the page.
Category 4. Extract
Participants envisioned interacting with reading materials by extracting elements for persistent references, such as creating keyword lists or extracting figures like scrapbooks. They also considered the possibility of extracting citations from academic articles’ bibliographies or footnotes. For example, one participant envisioned a feature where inline citations could be expanded on paper documents. Furthermore, there was interest in pulling out significant quotes from books over the years, enabling them to accumulate and store these quotes for future use.
Category 5. Navigate
Participants proposed various methods to enhance navigation through dense documents. They suggested the creation of document reading progress indicators to clarify their location within a text. Additionally, there was interest in making a table of contents readily accessible to better comprehend the document’s structure. Another idea involved implementing collapsible headings, allowing users to hide sections they find irrelevant in extensive documents, thereby facilitating a more focused reading experience. One participant suggested making the document searchable in AR. By integrating extraction features, they imagined users verbally dictating a search term and subsequently viewing a list of pages along with a concise summary of that term within the document.
Other Findings Related AI-Driven Augmented Reading
Additional findings include the potential benefits of AI to support idea generation and critical thinking. For example, participants believed that AI-assisted enhancements could deepen understanding by presenting multiple perspectives or providing new insights to enhance critical thinking. On the other hand, participants also identified several possible challenges associated with AI-generated content, including the risk of inaccurate information, instability of output quality, overly lengthy responses, and the possibility of information overload leading to distraction. In summary, while some of the proposed features are beyond the scope of our implementation, these design ideas, along with insights, have informed the development of our system, which we describe next.
4. System and Implementation
Building on insights from our formative study, we developed RealitySummary, a mixed reality reading assistant that utilizes real-time OCR and a large language model.
4.1. System Design
Figure 3 summarizes the overview of our system design. The system first detects and captures the document with the headset’s camera and proceeds to extract text from printed or digital documents using optical character recognition (Google Cloud OCR). Given the extracted text, our system employs the GPT-4 large language model through OpenAI’s API for analyzing and extracting important information from the document. Additionally, we have a middle layer on the server that parses the text using spaCy’s named entity recognition module to extract a list of persons, locations, organizations, and keywords that are used for the web search via Google search API. The processed information is transformed into either a text summary or visual aid with a prepared design template, which we built based on our formative study (Figure 4).
Figure 3. RealitySummary system architecture.
4.1.1. Mixed Reality Headsets and Image Capture
The system then embeds the text and visual output via the mixed reality headset. Since our pipeline is hardware independent, we deployed our system to Microsoft Hololens 2 and Apple Vision Pro. For HoloLens 2, we use an internal camera for image capture, Vuforia for image tracking, and handle content rendering with MRTK3 (Mixed Reality Toolkit 3) SDK [^2] Source:https://github.com/MicrosoftDocs/mixed-reality/blob/docs/mrtk-unity/mrtk3-overview/index.md. For Apple Vision Pro, we implement the same rendering and image tracking features using the RealityKit framework and Swift. Since the Vision Pro does not allow raw camera access, our prototype feeds the camera scene to a MacBook Air via AirPlay, then extracts images via screen capture. Additionally, we offer voice interaction support to modify or create these summary visualizations using Google’s speech-to-text API and a semantic similarity check between the transcript and feature descriptions to detect which feature the user intends to choose.
4.1.2. Text Extraction
We extract text content from physical or digital documents using OCR. We decided to use Google OCR in Google Cloud Vision API for accurate text extraction. OCR output also includes the position of each word, sentence, and paragraph, which helps embed the content for on-document augmentation. However, the OCR processing time is not fast enough for continuous capturing, we only first take a screenshot of the page every 5 seconds, hence, the system only needs to perform OCR once, as long as the page content does not change. We detect whether the user still reads the same page by comparing the new OCR results from the previous one. On top of the repeated capturing, the user can also manually activate the OCR with a voice command.
4.1.3. Document Tracking
We use Vuforia’s markerless image tracking or Vision Pro’s built-in image tracking features to track the document. The captured image is used for both OCR and document image segmentation, which runs in parallel. When the image is captured through the above step, we segment the document boundary with OpenCV rectangle detection. The system then uses the segmented image to generate a dynamic image target in Vuforia and Vision Pro. This allows RealitySummary to track essentially any document with sufficiently distinct visual features without attaching markers in advance. However, the accuracy of the system depends on various factors, such as lighting and tracked image pattern, the user can also manually place and move the content with hands. Recognizing the challenges of image tracking in diverse settings, our Vision Pro implementation also allows viewers to manually anchor content panels to the environment rather than the document. By taking advantage of the platform’s extremely robust spatial tracking, these environmentally-anchored panels (combined with its much higher-resolution displays) dramatically improve the readability of augmented content while reducing instability and jitter caused by moving image targets.
4.1.4. Text Summarization
After text extraction and document localization, the system extracts important information using NLP. We utilize spaCy and the GPT-4 large language model in the server. We use render AR content as a series of 2D canvases. We employ WebSockets communicate between mixed reality headset and the server with JSON format. We also store the OCR results, captured images, and GPT response history in Firebase database. We query multiple prompts to multiple GPT-4 API simultaneously to reduce the generation time. The specific implementation choices for each type of document enhancement are described in the next section.
4.1.5. Speech Input Interaction
The user can enable or disable different summarization panels simply by saying a trigger word like “show me the summary of this page” and “show me the table-of-contents”. Our system leverages Google’s speech-to-text API to convert spoken sentences into text. We leverage spaCy to measure the semantic similarity between the user’s spoken sentence and our predefined list of action descriptions. Beyond a simple summary activation, the user can also ask free-form, open-ended questions about the reading content by simply saying a question, which is directly sent to GPT-4 API along with the OCR results.
4.2. Document Enhancements
In this section, we describe our supported document enhancements. On top of the elicited designs that we derive from our formative study, our design is also inspired by the prior MR- and non-MR-based reading assistant, such as HoloDoc (Li et al., 2019), DuallyNoted (Qian et al., 2022), Marvista (Chen et al., 2022), and SummaryLens (Benharrak et al., 2022). While we acknowledge that individual design elements may not be significantly novel, we believe that our system contributes to the following two aspects: 1) our holistic framework and its evaluation provide a better understanding of document augmentation design space, 2) automated and real-time approach to creating on-demand content opens up a new potential of mixed reality assistant tools.
Each of our document enhancement techniques relies on specific language model queries (more commonly known as prompts, which are documented in our supplemental material), often in combination with additional web services like image search for images and people bio information cards, and mapbox for generating maps. Figure LABEL:fig:documents_user_study shows various reading materials augmented with a variety of information through our system.
Figure 4. RealitySummary document enhancement features. (All images captured live from HoloLens 2.)
Summary
The text summary is extracted directly by using a GPT-4 API call from the Node.js server. We send the entire text of the current page to GPT-4 and get a summary in return, which is then sent to be rendered on HoloLens. We simply prompt GPT-4 asking it to summarize the current text data extracted from the page. This technique is often used in prior summary-based reading assistants like Marvista (Chen et al., 2022). We limit the maximum word length of the summary to less than 150 words, as we learned the user struggles to digest lengthy text through the formative study.
Comparison Table
RealitySummary generates comparison tables based on the user’s prompt. We use GPT-4 to generate a comparison table between the two items along with contextual text from the document similar to generating text (see supplementary material for the prompt). Alternatively, while parsing the document, we prompt GPT-4 beforehand to generate possible comparison tables and store them locally through the session (and query directly from it when comparison action is detected).
Timeline
A timeline is a series of summarised events in a given text. When a user requests a timeline, we prompt GPT-4 to generate a series of events in the form of a timeline based on the user’s query and get a JSON with the relevant data. We then render this using Unity Canvas (Hololens) or HTML Canvas (Vision Pro) in a custom format to make it easy for the user to skim through.
Keyword List
The keyword list is generated using both GPT-4 and spaCy. SpaCy is used for keyword and phrase extraction, which also provides named entity such as date, location, and persons, which will be used for web search. We simply process the text by calling spaCy’s python wrapper using python-shell from our Node.js server and send the result back to the headset for rendering a list of keywords or highlighting them using word/phrase location, similar to keyword-based skimming techniques like ConceptEVA (Zhang et al., 2023).
Summary Highlighting
Summary Highlighting works very similarly to summaries, in that, the user asks to summarize or find more information about a specific phrase, sentence, or paragraph. Once the user wants information about content inside a document, we prompt GPT-4 for a brief summary and use word/phrase, sentence, or paragraph matching to find the location of the content in the document. We then highlight, draw arrows, or annotate based on OCR’s word bounding box results. Other features also leverage word and phrase highlighting to connect relevant information directly to the text in the document.
Information Cards
Inspired by AR-based annotations demonstrated in DigitalDesk (Wellner, 1991) and DuallyNoted (Qian et al., 2022), information cards, as the name suggests, contain information about keywords and people (image with a description), places (a map), and images about common keywords. To reduce the number of requests made to the LLM, we use Google Search API to extract information related to key terms, places, or people. We also use Google Image Search API to source pictures related to terms or phrases on demand. Moreover, we use mapbox API to generate maps. In the rare case that the API fails to provide a response, we query GPT-4 for more information.
4.3. Technical Evaluation
To characterize the technical performance of our prototype, we assessed the reliability of document tracking and overall system latency, as well as the tool’s ability to recognize documents of a variety of different types. To measure the system performances, we also systematically tested RealitySummary with (1) AR tracking, (2) OCR, and (3) summarization relevance/correctness using 20 additional documents (10 physical and 10 on-screen) with greater visual and technical diversity, including posters, newspapers, online articles, patents, magazines, thesis, books, blog posts, etc.
4.3.1. Document Tracking
The AR tracking analysis measured the system’s document tracking reliability with Microsoft Hololens 2 over 60 seconds each. We saw reliable tracking for 14 documents that included graphics (M=92%, SD=5.4% up-time). However, the 6 documents that included only text showed worse performance (M=64%, SD=3.7% up-time) due to their lack of distinct trackable features. AR tracking is not always very reliable, especially with text documents because they have fewer feature points to track. We used system logs to automatically analyze document tracking accuracy throughout the user study. We measured tracking loss whenever the system lost AR rendering or image tracking. We found our system’s tracking to be 72% accurate while tracking documents with only black-and-white text and 92% accurate with tracking documents with some colors and visuals (images, diagrams, charts, etc.). We also found substantial changes in tracking with different lighting conditions (better tracking with better lighting), document image tracking size (large image targets performed much better than smaller documents), and motion (sudden headset movements resulted in temporary tracking loss or flickering).
4.3.2. Applicability of Physical or Digital Content
OCR analysis examined the fraction of words correctly extracted from the documents. Across the 20 documents, we saw extremely high OCR accuracy (M= 97.9%, SD=5.0%). We tested our system’s generalizability and applicability with 20 different document types—newspapers, magazines, academic papers, books, legal documents, and technical documents. Two authors checked each summarization feature on all 20 documents manually to evaluate the applicability and accuracy of the content generated by the large language model. We found our system to be 94% accurate when generalized to other documents in providing correct information based on the context of the document. Summarization analysis verified the relevance and factual correctness of the initial 120 (20 $\times$× 6) summary cards generated by RealitySummary across these 20 documents. Two authors manually examined and verified all documents and cards, and found that 96.77% were both factually correct and relevant.
4.3.3. Latency
In our technical evaluation of latency, we measured the average latency of three different components: AR rendering, network communication, and API responses. Network communication represented our socket.io connection between the Node.js server and the headset client. The system recorded an average latency of 100 ms for rendering AR content in HoloLens and Vision Pro and an average latency of 40 ms for socket.io network connection, which is relatively low and suggests that the rendering and network process is responsive and efficient. On the other hand, API callbacks included two main components, namely external request to the GPT-4 model and pyShell request to our own NLP processing unit built with spaCy. Latency analysis included their API callbacks and wait-time. The average latency for GPT-4 was 2.4 s and NLP Module was 670 ms.
5. Usability Study
We evaluated our system with 12 participants via a think-aloud usability study. Our goal was to qualitatively assess our system’s usability, examine participants’ reactions to our document enhancements, and elicit new opportunities for MR reading assistants.
5.1. Method
We recruited participants (P1-P12, 6 male, 6 female; Ages = 19-35 years) through word-of-mouth and they were compensated with \$15 for their one-hour participation in the study. The study was conducted with Microsoft Hololens 2 in a lab environment. All the participants had verbal and written fluency in English.
Before starting the study, the authors familiarized participants with AR and the Hololens 2, then introduced RealitySummary and each of the document enhancements described in the previous section. After this introduction, we provided three reading tasks using: 1) an A4 size magazine in physical format, 2) a digital document of the participant’s choice, and 3) an A3 size book. Participants’ digital documents included 3 books, 2 blog posts, 2-course assignment PDFs, 2 Wikipedia articles, 2 magazine articles, and 1 news article which they viewed on either a PC (4/12), tablet (5/12), or laptop (3/12). We asked participants to use RealitySummaryto aid with their reading, using whatever combination of enhancements best suited the current document. Throughout this 30-minute period, we encouraged participants to think aloud, documenting their experiences with the system. After the tasks, we conducted semi-structured interviews and asked them to complete a Likert scale questionnaire, including a System Usability Scale (SUS) protocol (Brooke, 1995). Together, the questionnaire and interview took roughly 20 minutes.
Afterward, we conducted a thematic analysis by first transcribing the think-aloud sessions and interviews, then individually familiarizing ourselves with the content and organizing data into rough categories. Two authors then individually searched for broader groups and patterns and clustered these into themes. The writing team then iterated to further refine the identified themes.
5.2. Results
Participants responded positively to RealitySummary (Figure 5), with most participants either agreeing or strongly agreeing that the the prototype was easy to use (10/12), easy to understand (11/12), and well-integrated with the task (10/12). Similarly, nearly all participants reported that the summary approaches provided by RealitySummary helped with comprehension (12/12), were relevant to the document content (11/12), improved comprehension speed (10/12), and augmented their reading experience (10/12). Finally, the System Usability Scale (SUS) returned an overall composite score of 71, suggesting a reasonable level of usability for an early-stage prototype.
Figure 5. Usability study questionnaire responses.
5.3. Findings
Our analysis identified five themes related to participants’ experience comprehending and navigating documents, as well as their sense of focus and control. We also discuss our participants’ comparisons of the visual and text-based summarization techniques and layouts implemented in RealitySummary.
Comprehension
Document enhancements can improve readers’ sense of comprehension, and readers value the ability to see multiple types of enhancements simultaneously. All the users in the study felt that the system improved their reading comprehension. Out of the 12 participants surveyed, 11 reported that they were able to understand the document without actually reading it, solely by relying on the contextual information provided by RealitySummary. P9 even stated, “I didn’t even read the document but I understand what it is about.” Furthermore, multiple participants (P7, P8, P11, P12) argued that the combination of multiple features like summaries, timelines, keywords, images, and people cards provided them with a more comprehensive understanding of the document, allowing them to engage with it from different angles. This allowed them to gain multiple perspectives on the same topic, leading to a deeper comprehension of the material. All participants mentioned that RealitySummary helped them comprehend documents more quickly than their typical approaches (like visually scanning a physical book and manually searching for a related summary or timeline on a phone/PC). For example, after perusing the first magazine, P12 noted that “I’ve not read this magazine thoroughly but, it seems that these features can significantly help me understand what this document is about in less amount of time.”.
Navigation
Both structured overviews and in-context highlights provide useful entry points that can make it easier to navigate documents in a top-down fashion, even within individual pages. Broadly, participants spoke positively about how RealitySummary helped them navigate documents. Four participants (P3, P4, P6, P8) specifically reported using structured navigation elements like timelines to get a better understanding and structural flow of the whole text, with P4 noting that “the timeline helped me organize and navigate my thoughts very quickly while reading the text” (P4). Meanwhile, at least a third of participants (P5, P6, P9, P10) also mentioned using lower-level augmentations to help navigate the document— highlighting how the proximity of annotations and content allowed them to use keyword highlights and information cards as visual anchors. Additionally, several participants found the connection lines between highlighted words and their corresponding features to be a valuable tool for navigating the document—with P8 stating, “the words in the documents are highlighted […] and connected with all the features […] it is very helpful. It is also pretty helpful that I can see the word ‘Yosemite’ and where it is on the map around the text!”.
In fact, ten different participants reported that the system allowed them to focus their reading by taking a “top-down” approach—first examining high-level ideas and terms surfaced by RealitySummary, then gradually transitioning into reading lower-level content where necessary. P8 highlighted that “[the enhancements] give you a foundation before you actually jump inside [the document]”. This allowed participants more leeway to prioritize their reading to “focus on things that you’re more unsure about rather than just kind of trying to find external information” (P1).
Interestingly, several participants also reflected positively on the fact that summaries and keywords surfaced in our prototype were drawn only from the currently visible page, rather than the document as a whole. P5 reported that, “I like page by page summary—having a summary for each page does help make it feel less dense compared to a summary of the whole document”. Similarly P6 stressed that, “I like that it (system) gives you keywords for each page and not as a whole.
Visual vs. Text Enhancements
Both visual and text-based enhancements can benefit readers, with visuals providing supporting high-level overviews and providing context, and text summaries capturing more complex concepts. While our participants responded positively to all of the document enhancements in the prototype, they reported that they preferred text features when trying to gain a deeper understanding of a particular topic or content within the document. Conversely, participants generally felt that visual features improved their comprehension speed, increased engagement, and improved memorability.
All 12 participants ranked the visual summarization techniques—including image cards, people bios, and maps—as helpful, with eight participants specifically highlighting the value of images and bio cards. Many participants (P1, P4, P6, P7, P9, P10, and P11) agreed that the visual augmentations helped them understand the document faster compared to just reading the text, noting that, “seeing the visuals helped me grasp what’s happening in the document right away” (P11). Several participants (P7, P8) also stressed the potential for visuals to increase their engagement with the text, with P7 stating, “I’m not a big fan of reading so having visuals are helping me assist in reading this”. These participants also highlighted the potential for visuals to increase memorability with P8 noting that for historical documents the people cards could “help them put a face to a name”. Similarly, P7 noted that while text-based summaries might contribute to greater understanding, “graphical stuff makes me remember it for a longer time”.
Layouts and Visual Clutter
Placing enhancements in the space around the document using mixed reality can successfully reduce the potential for visual clutter and distraction. Because they can sit alongside and even on top of document content, summaries, information cards, and other document enhancement have the potential to occlude or distract attention from the document. However, reducing the number of enhancements or moving them further from the text may limit their utility. In our case, several participants (P7, P10, P11) specifically highlighted how our around-document placement of the layouts integrated nicely with the documents themselves—calling them natural and intuitive to use. P10 reported that “[the system] is like a natural extension of the book. It is very easy to follow through.”. Similarly, four participants (P1, P2, P5, P11) praised the fact that the augmentations did not disrupt their visual focus on the document, with P5 noting that “it is not cutting into your field of view and it’s not invasive at all”. Moreover, 10 participants also remarked positively on the connecting lines linking document text to cards around around the document—indicating that they were easy to follow but did not impede readability.
6. In-the-Wild User Study
To explore additional real-world use cases, we also conducted an in-the-wild study with eleven participants. In contrast to the previous usability study, which focuses on standardized reading tasks in a lab setting, this study instead focused on examining the versatility and applicability of our system across uncontrolled environments and diverse reading materials. Our goal was to uncover unique advantages and potential use cases of employing mixed reality and AI-generated assistants in everyday settings. For this study, the system was deployed on an Apple Vision Pro, which provides considerably higher rendering quality than the Hololens 2. To accommodate diverse applications and reduce the obtrusiveness of additional content in everyday scenarios, the user interface was refined, limiting augmented content to a text box and an image slot.
Figure 6. In-the-wild study.
6.1. Method
We recruited 11 participants (W1-W11, 7 male and 4 female; Ages 19-32 years) from the local community. They were compensated with \$25 for their participation in a 1.5-hour study, which consisted of a one-hour experiment and a half-an-hour interview. The study was conducted in various locations at a local university, including a library, cafeteria, store, hallway, classroom, and multiple lab settings.
Before the experiments, we collected participants’ demographics and their reading habits, which were input into the system as the prior knowledge of ChatGPT. Participants familiarized themselves with Apple Vision Pro by doing the eye tracking calibration setup. Unlike structured tasks or prepared reading materials, participants were encouraged to utilize our system with any text content they encountered during the experiments, including documents, books, papers, posters, online news on mobile devices, and other materials in their environment. Throughout this one-hour experiment, we also encouraged participants to think aloud about their experiences. After the experiment session, we conducted semi-structured interviews, followed by a Likert scale questionnaire similar to the usability study. Both the interview and questionnaire took approximately 30 minutes. As in the earlier usability study, we then analyzed participants’ questionnaire responses and reflections to identify high-level trends and opportunities for augmented reading tools.
6.2. Results
We gathered feedback from participants on various aspects including the reading experience (immersiveness, understanding, distraction), perceptions of the system (content and query comprehension, trustworthiness, and privacy concerns), and overall usability (content placement and comfort using the MR headset). In general, most participant either agree or strongly agree that RealitySummary deepened their understanding of the reading content (7/11), RealitySummary demonstrated a deep understanding of the content they were reading (7/11). Participants reported feeling more immersed and interested while reading, with an average rating of 3.45 (SD = 1.04), and they perceived the system as effectively interpreting their queries (M = 3.27, SD = 1). Besides, their responses on the content placement within the system were also overall positive (M = 3.36, SD = 1.29).
Participants had varied opinions on whether the system became distracting throughout the reading process (M = 2.55, SD = 1.37), and whether they tended to trust the accuracy and reliability of contents provided by the system (M = 2.82, SD = 1.25). Privacy concerns were also noted by the majority (6/11), and discomfort using the headset was prevalent among the respondents (8/11). These findings are detailed in Figure 7.
Figure 7. In-the-wild study questionnaire responses.
6.3. Findings
Observations from the in-the-wild deployment highlighted a wider range of use cases for augmented reading as well as some of the practical trade-offs associated with real-world use.
Types of Reading Content
The in-the-wild study demonstrated our system’s broad utility, revealing diverse exploratory use cases across various reading materials. Participants engaged with a variety of books, spanning textbooks, encyclopedias, fiction, and non-fiction. The system was also employed for different types of documents, including academic papers, handouts, and class assignments, as well as digital content accessed through smartphones and laptops, like social media posts, online news articles, and e-commerce websites. Beyond conventional reading materials, participants also used our system with everyday objects and texts in the environment, such as restaurant menus, printer instructions, product nutrition labels, furniture assembly guides, posters, and signage on trash bins and advertisements. Notably, some participants even applied our system to their handwritten notes from lectures.
Benefits of Always-On Assistant
The exploration of such diverse reading materials can be attributed to the always-on capability of the mixed reality reading assistant. A key advantage of this always-on feature is its support for implicit interactions, offering a distinct user experience compared to ChatGPT on smartphones or laptops, which require explicit effort to engage. This implicit and automatic content capture greatly expands the range of applications, beyond what participants initially considered possible. The always-on capturing feature eliminated the need for manually supplying context, saving considerable time. For example, W9 experienced a smoother workflow when working on her business course project with RealitySummary compared to conventional screen-based methods. She was able to concentrate on her project, without having to switch between smartphones or laptops to familiarize ChatGPT with the context.
Unique Affordances of Spatial and Tangible Interactions
Beyond the always-on implicit assistance, participants appreciated the unique affordances offered by mixed reality interfaces, such as the opportunities for spatial and tangible exploration. In a library setting, for example, participants could naturally navigate through bookshelves (spatial exploration), physically pick up a book (tangible interaction), and browse through its pages to automatically receive summaries (always-on assistance). W5 and W10 even leveraged our system to provide summaries or recommendations for entire collections of books on a bookshelf, not just content from a single page. Furthermore, W1 demonstrated the capability to synthesize information across multiple books on a table, requesting summaries or connections between various books and pages. Such observations highlight how mixed reality interfaces enable spatial and tangible interactions that go beyond the limitations of traditional screen-based interfaces, which can only capture and summarize visible content.
Unique Ways of Using AI-Generated Content
We observed several unique and unexpected use cases that took advantage of the LLM’s capacity for free-form questioning and answering. For instance, when reading a physics textbook, W8 asked the system to generate ten quizzes to help him understand quantum mechanics, transforming the textbook into a more active and personalized learning tool. Other common usages included translations (W4, W5, W6), learning aids (W4, W8), and personalized content recommendations (W5, W10). For example, applied the system to his handwritten calculus notes to find answers to particular questions. Beyond traditional reading applications, the AI’s capabilities supported practical decision-making; W4 compared nutritional information of products on store shelves and menu items to determine healthier options, while W5 used it for guidance on waste disposal and printer operations. The LLM’s ability to process a broad spectrum of inquiries significantly expanded the system’s applicability and utility.
Proactive Summaries vs On-Demand Question Answering
Our system offers both 1) proactive assistance through automatic summary generation and 2) on-demand assistance where users manually ask questions. Participants shared their insights on these two functionalities by comparing the benefits and limitations. Proactive summaries were appreciated for providing a useful initial overview of unfamiliar topics. Preferences for proactive summaries varied with reading habits. For example, infrequent readers like W6 and W10 found these proactive summaries helpful in deciding whether to read the content or understand it without reading the full text. On the other hand, W10, who is a more frequent reader, found the constant summary updates unnecessary. In contrast, on-demand question answering allowed for a more focused engagement on specific topics. For instance, W1 enjoyed the ”rapid-fire” questioning with a physics textbook, asking more questions than with traditional ChatGPT saying ”You have to describe it properly you have to specifically say what you’re asking about what you’re reading, so that kind of all of this abstracted in this system is like frees up your mind to just focus on the curiosity part Like rapid-fire questions and get rapid-fire answers”. W2 valued how questioning was woven into reading, preserving thought flow without the disruption of toggling between reading and ChatGPT interfaces. Overall, participants favored the on-demand feature due to its applicability to diverse situations, yet many appreciated having both options for different purposes: proactive features for implicit skimming and on-demand for addressing more specific questions.
Trust of AI-Generated Content
Many participants approached the system’s responses with caution, often verifying the answers with the reading material. They highlighted the necessity of consulting additional sources for in-depth or technical readings when depending on the system. There was a general agreement among participants about their reluctance to fully trust an always-on reading assistant in their daily activities. Conversely, we noted some intriguing discussions regarding how participants perceive our system, either as a mere tool or more like a personal companion. While most participants engaged with the system as a conventional chatbot, some regarded it more as a personal companion than a mere tool. Notably, W3 and W5 described the system as a friend for discussing and inquiring about their readings. The system’s ability to provide always-on immediate responses facilitates this personification.
Potential Risks of AI-Driven Augmented Reading
Participants also expressed several concerns about potential risks, notably privacy issues due to the possibility of constant surveillance. Despite this concern, most participants were willing to continue using the system, if they could control when it was active. For example, W1 expressed the desire to capture all text throughout her entire life, allowing for an extensive searchable archive of everything she has encountered. Additionally, participants also mentioned the risk of decreased motivation and skill development. For example, W6 mentioned that relying too heavily on the system for translations could reduce the motivation to learn new languages. Similarly, W11 worried that immediate access to answers for every question might contribute to a broader intellectual complacency. Overall, participants wanted the flexibility to choose when and how the system would be used to mitigate these risks.
Possible Future Improvements
A desired feature among participants was the system’s ability to recognize not only text but also images. For instance, W1 faced limitations when attempting to inquire about a diagram on a page, which the system could not recognize. Similarly, they anticipated broader usage scenarios beyond text recognition, such as generating cooking recipes from fridge ingredients or providing information on real-world objects. We anticipate that enhancing the system with multi-modal LLM capabilities could address such needs. Additionally, there was interest in a “chat history” feature to recall previous interactions with the system, highlighting a need for temporal relevance in user-system interactions. Finally, participants suggested they would be more inclined to use such a system long-term if hardware constraints were overcome.
7. Limitations and Future Work
Our experiences building RealitySummary and the insights from both our formative workshop and study highlight a range of promising research directions for MR document enhancement.
Technical Limitations
We first recognize that our system exhibits numerous technical challenges that need addressing in future enhancements. Concerning AR tracking difficulties, our evaluation found that AR tracking becomes unstable with text-only documents. This instability is exacerbated in low ambient light, diminishing the tracking’s effectiveness and reliability. Such issues raise questions about the AR technology’s adaptability and dependability for daily use. Additionally, OCR bottlenecks and image tracking errors introduced complications related to processing speeds, data transmission efficiency, and reliance on the quality of images. Given the variance in document quality, the performance of OCR and image tracking can fluctuate, leading to possible inaccuracies and delays for different types of documents. Lastly, LLM generated errors show areas for improvement, particularly in situations where the text is ambiguous or highly specialized.
From Page to Documents
Currently, the system cannot display summaries, annotations, or relationships spanning multiple pages. As the system’s inputs are limited to visible pages, content not yet encountered cannot be referenced. Future research might explore a hybrid approach. For example, if a digital version of the document is available online, the system could access that content to improve summarization. These enhancements could include the types of navigation interactions (such as dynamic tables of contents or visual reading histories) highlighted by participants in our formative study. Connecting extracted text with complete digital versions of documents could enable new visual search, preview, and comparison tools that build upon the interactions seen in prior prototypes like HoloDoc (Li et al., 2019). By merging the capabilities of augmented reality with extensive digital information repositories, we believe this gap can be bridged in the future.
Beyond Text-Based Summaries
While our current work has focused on examining text-based summarizations using a large language model, our formative study (Section 3.2) also highlights the potential for visual summarization approaches, including automated mind-mapping and charting tools. Techniques like these could support more efficient search, comparison, and integration, building upon word-scale (Goffin et al., 2014) and micro visualization techniques (Parnow and Dörk, 2015) designed to reveal data in and around existing text. Integrated paragraph-level summary visualizations, such as those introduced in Caragraphs (Masson et al., 2023), are especially promising. They would be well-suited to the kinds of placements around documents that mixed reality reading tools enable.
Multimodal Inputs
The improving multimodal input capabilities of mixed reality devices and LLMs, such as visual and audio input, eye tracking, more precise hand tracking, and richer gestural interactions, create additional opportunities for dynamic reading assistance. For instance, the eye-tracking capabilities of modern MR hardware could be used to estimate gaze and incorporate that information to tailor summaries according to the viewer’s current task or area of interest. If the system also captures verbal communication in addition to written text, like in RealityTalk (Liao et al., 2022), we could use this for conversation and discussion assistance. In this scenario, summaries would reflect elements of the discussion. Combining speech and physical manipulation could allow users to dynamically request, expand, adapt, or dismiss summaries and support interactions in a wider range of contexts and environments (Oviatt, 2007). Finally, the use of multimodal LLMs to receive graphical, environmental, and gestural input have the potential to broaden these systems’ applicability.
Long-Term Evaluation
While it highlights a variety of opportunities for augmented reading tools, our in-the-wild study is still a short-term evaluation. Long-term studies, in which participants interact with these systems for weeks or even months will almost certainly reveal deeper insights and challenges. Long-term studies are challenging given the current limitations of mixed reality hardware like the Hololens 2 and Vision Pro, which are not comfortable enough for everyday use. We anticipate that the continuous improvement of the headset hardware will make long-term studies feasible in the future.
8. Conclusion
RealitySummary is the first example of an on-demand MR reading assistant that uses OCR and an LLM to generate reading aids for any physical or digital document. We hope our study opens up new opportunities for on-demand and natural language-powered reading assistance and inspires the HCI community to continue exploring such systems to realize the full potential of intelligent reading tools.

### v2

@@
-RealitySummary: On-Demand Mixed Reality Document Enhancement using Large Language Models
+RealitySummary: Exploring On-Demand Mixed Reality Text Summarization and Question Answering using Large Language Models
 Abstract.
-We introduce RealitySummary, a mixed reality reading assistant that can enhance any printed or digital document using on-demand text extraction, summarization, and augmentation.
-While augmented reading tools promise to enhance physical reading experiences with overlaid digital content, prior systems have typically required pre-processed documents, which limits their generalizability and real-world use cases.
-In this paper, we explore on-demand document augmentation by leveraging large language models.
-To understand generalizable techniques for diverse documents, we first conducted an exploratory design study which identified five categories of document enhancements (summarization, augmentation, navigation, comparison, and extraction).
-Based on this, we developed a proof-of-concept system that can automatically extract and summarize text using Google Cloud OCR and GPT-4, then embed information around documents using a Microsoft Hololens 2 and Apple Vision Pro.
-We demonstrate real-time examples of six specific document augmentations: 1) summaries, 2) comparison tables, 3) timelines, 4) keyword lists, 5) summary highlighting, and 6) information cards.
-Results from a usability study (N=12) and in-the-wild study (N=11) highlight the potential benefits of on-demand MR document enhancement and opportunities for future research.
-Mixed Reality; Natural Language Processing; Large Language Models; Augmented Reading
-journalyear: 2024copyright: acmlicensedconference: ; May, 2024; $c$s: Human-centered computing Mixed / augmented reality
+Large Language Models (LLMs) are gaining popularity as tools for reading and summarization aids.
+However, little is known about their potential benefits when integrated with mixed reality (MR) interfaces to support everyday reading assistants.
+We developed RealitySummary, an MR reading assistant that seamlessly integrates LLMs with always-on camera access, OCR-based text extraction, and augmented spatial and visual responses in MR interfaces.
+Developed iteratively, RealitySummary evolved across three versions, each shaped by user feedback and reflective analysis: 1) a preliminary user study to understand user perceptions (N=12), 2) an in-the-wild deployment to explore real-world usage (N=11), and 3) a diary study to capture insights from real-world work contexts (N=5).
+Our findings highlight the unique advantages of combining AI and MR, including an always-on implicit assistant, minimal context switching, and spatial affordances, demonstrating significant potential for future LLM-MR interfaces beyond traditional screen-based interactions.
+Mixed Reality; Large Language Models; Augmented Reading; In-the-Wild Study; Diary Study
+journalyear: 2024copyright: acmlicensedconference: arXiv; September, 2024; $c$s: Human-centered computing Mixed / augmented reality
 Figure 1.
-RealitySummary provides an on-demand reading assistant with embedded mixed reality (MR) text summaries and visual aids.
-Using markerless tracking, OCR, and GPT-4, the system can provide live summaries, keyword lists, timelines, and text highlighting based on the real-time extraction, analysis, and visualization of any physical or digital document.
-(Simulated high-resolution illustration based on real Unity assets associated with content in the article shown.)
-1. Introduction
-“We live in a complex world, filled with myriad objects, tools, toys, and people.
-[..] Yet, for the most part, our computing takes place sitting in front of, and staring at, a single glowing screen.
-[…] How can we escape from the computer screen and bring these two worlds together?” — Pierre Wellner (Wellner, 1991)
-In his pioneering work (Wellner, 1991), Pierre Wellner presented the first demonstration of AR-enhanced documents, arguing that the future of computer interfaces should augment our physical environment, rather than confining us to a flat rectangle screen.
-This fusion harnessed the dynamic capabilities of digital content while retaining the tangible benefits of physical documents.
-With recent advances in mixed reality technologies like the Microsoft Hololens and Apple Vision Pro, we are getting closer to realizing his vision where we can naturally augment surrounding paper, books, and documents, as well as content on screens, to enhance our reading and learning experiences (Li et al., 2019; Rajaram and Nebeling, 2022; Chen et al., 2020).
-However, most prior systems (Billinghurst et al., 2001; Zhao et al., 2014; Li et al., 2019) have typically required documents to be pre-processed with pre-generated content.
-The lack of on-demand and real-time content generation for non-prepared documents hinders their generalizability and practicality in real-world applications beyond research labs.
-This paper introduces RealitySummary, a mixed reality reading assistant that enhances any printed or digital document using on-demand text extraction, summarization, and augmentation.
-Our on-demand document augmentation leverages optical character recognition (OCR) and a large language model (LLM) to enhance any given document without prior preparation.
-To design our system, we first conducted a formative design workshop with six participants to identify common document enhancement techniques across diverse document types.
-We identified 20 unique techniques across five categories of document enhancement: summarization, augmentation, navigation, comparison, and extraction.
-Insights from the study led us to incorporate the top six techniques into our system: 1) summaries, 2) comparison tables, 3) timelines, 4) keyword lists, 5) summary highlighting, and 6) information cards.
-Based on this insight, we developed our system for the Microsoft Hololens 2 and Apple Vision Pro.
-The system first captures the document, and then extracts text and the document image using Google’s Cloud OCR API and OpenCV.
-Given this extracted text, the system generates different augmented content and extracts related images through GPT-4 and Google Image Search.
-The system also tracks the document based on the extracted document image, rendering the generated content around the document as a spatial 2D canvas.
-Furthermore, we integrated a speech input using Google’s speech-to-text API, enabling users to activate different summarization features or ask questions related to the current content through voice commands.
-To evaluate our system, we conducted two user studies: 1) a usability study with twelve participants, and 2) an in-the-wild study with eleven participants.
-Both studies highlighted the unique benefits of contextual assistant for diverse set of reading materials without prior preparation.
-Participants valued the always-on feature for enabling implicit interactions and appreciated the mixed reality interface for facilitating spatial and tangible exploration, significantly enhancing user engagement and the applicability of AI-generated content across various scenarios.
-Based on the study insights, we also discuss potential risks and future research opportunities of AI-powered mixed reality reading assistants.
-Finally, our paper contributes:
-  1. (1)
-    RealitySummary, a mixed reality reading assistant that uses a large language model to provide on-demand content summarization and augmentation.
-  2. (2)
-    A set of design features and summarization techniques for mixed reality document enhancement derived from a formative design workshop (N=6).
-  3. (3)
-    Insights and findings from our usability study (N=12) and in-the-wild study (N=11) that highlight the potential benefits of on-demand mixed reality document enhancements.
+We developed RealitySummary, an on-demand mixed reality reading assistant, over three iterations and conducted three different studies.
+Through this evolution, we highlight three key findings: 1) unique benefits of MR and LLMs, 2) emergent use cases of MR reading assistants, and 3) design implications for mixed reality reading assistants.
 1. Related Work
+We summarize prior work on reading augmentation and content summarization with a focus on interactive approaches.
 1.1. Reading Augmentation
-2.1.1. Document Enhancement
+Document Enhancement
 Human-computer interaction researchers have explored ways to augment physical documents with AR-enhanced content.
@@
 However, none have yet achieved comprehensive and general-purpose document enhancement.
-In this paper, we contribute to the first demonstration of on-demand document enhancement by leveraging large language models.
-The goal of this paper is not only to demonstrate its potential but also to uncover its limitations and challenges from a user experience perspective through user evaluations.
-2.1.2. Screen-Based Reading Support Tools
+Screen-Based Reading Support Tools
 Outside of the context of mixed reality, a wide range of projects and systems have explored reading support tools for screen-based interfaces.
@@
 1.2. Content Summarization
-2.2.1. Automatic Summarization
+Automatic Summarization
 In the field of natural language processing, several works have employed automatic summarization to support readers (El-Kassas et al., 2021; Zhang et al., 2023).
@@
 Various types of summarization have been explored using GPT-3, such as opinion summarization (Bhaskar et al., 2022), news summarization (Goyal et al., 2022), and medical dialogue summarization (Chintagunta et al., 2021) while others have built custom frameworks for feeding long texts into GPT-3 to elicit more finely-controlled summarizations (Payne, 2022).
-2.2.2. Interactive Summarization.
-Instead of automatically summarizing the content, the researchers have also explored interactive summarization based on the context and user’s needs.
+Interactive Summarization.
+Researchers have also explored interactive summarization approaches based on context and users’ needs.
 For example, Hoeve et al.
 (ter Hoeve et al., 2020) studied Q&A interactions to assist users with a document-centric question-answering paradigm, finding that information-seeking was a common query.
-Semantic Reactor (Google, [n.d.]), developed by Google, enables users to ask questions about documents in natural language by fine-tuning BERT (Devlin et al., 2019) using the SQuAD dataset (Rajpurkar et al., 2018).
+Semantic Reactor (Google, [n.
+d.]), developed by Google, enables users to ask questions about documents in natural language by fine-tuning BERT (Devlin et al., 2019) using the SQuAD dataset (Rajpurkar et al., 2018).
 VERSE (Vtyurina et al., 2019) is another Q&A system that allows users to ask specific types of questions about a document but is limited by its inability to summarize.
@@
 Our goal is to build on top of this prior work, developing and exploring automatic and interactive summarization techniques for on-demand and real-time summaries embedded in documents.
-3. Exploring Document Enhancement Techniques for Augmented Reading
-To explore the possible design of an AI-assisted augmented reading, we conducted formative design workshops with six participants (4 males, 2 females), who had various design backgrounds as undergraduate and graduate HCI researchers.
+The limitations of prior works such as Holodoc (Li et al., 2019), as explained before, prevented the study of long term and in-the-wild use of MR reading assistants.
+To our knowledge GazePointAR (Lee et al., 2024) is the closent work which explored long-term use of pronoun disambiguation in MR voice assistants.
+Still, it is unclear how can LLM powered MR assistants are used in everyday life to assist users in various types of reading and knowledge work tasks.
+We aim to fill this gap.
+In this paper, we contribute to the first demonstration of on-demand document enhancement by leveraging large language models.
+The goal of this paper is not only to demonstrate its potential but also to uncover its limitations and challenges from a user experience perspective through iterative implementation and multiple user evaluations.
+2. Overview of Our Research Approach
+Figure 2. Overview of our research approach spanning an evolution of our system across three different versions.
+To uncover insights into on-demand reading assistants, we employed a three-phase iterative design process to investigate user interactions and experiences with RealitySummary.
+Our approach consisted of a usability study on the initial prototype (Study 1: RS-Doc), followed by an in-the-wild study on a second prototype (Study 2: RS-Wild), and concluded with an author-conducted diary study using a third prototype (Study 3: RS-Diary).
+Each phase was designed to build upon the insights gained from the previous round, allowing for continuous refinement of the design and functionality.
+We chose this multi-method, iterative approach to capture a comprehensive understanding of user interactions across various contexts.
+The initial usability study provided controlled insights into the basic functionality and user experience of RealitySummary.
+The subsequent in-the-wild study allowed us to observe how users interacted with the improved, more reliable prototype in everyday environments, uncovering real-world usage patterns and challenges.
+Finally, the diary study offered deep, longitudinal insights into the integration of RealitySummary into daily life, allowing for nuanced observations of subtle interaction patterns and evolving user needs.
+This progression from controlled to naturalistic settings, coupled with iterative prototyping, allowed us to holistically explore the complex dynamics of AR-AI interactions while addressing emerging design considerations and ethical implications at each stage.
+3. Study 0: Elicitation Study
+Before implementing our system, we conducted a design elicitation study to explore potential designs for an AI-assisted augmented reading assistant.
+We held formative design workshops with six participants (4 males and 2 females) who had various design backgrounds as undergraduate and graduate HCI researchers.
+The study involved two tasks, each lasting approximately 15-20 minutes, with the entire session taking about 60 minutes.
 3.1. Method
-We asked participants to complete a four-part design task where they read and graphically annotated both their own document (Table 1) and a Wikipedia article [^1] Source: https://en.wikipedia.org/wiki/Russian_invasion_of_Ukraine.
-After reading the documents, participants brainstormed possible ideas for AI-powered augmentations to enhance their readability.
-In the first design task, participants annotated a physical printout of their own article, sketching ideas for possible enhancements, and then did the same with a printed copy of the Wikipedia article.
-In the second design task, we gave participants digital copies of the two documents on a virtual whiteboard (Miro) and asked them to design a second set of document enhancements with the assistance of ChatGPT.
-Each task took approximately 15 minutes, and each session of the entire study took approximately 60 minutes.
-Document Types Topic Pages Words
-P1 Academic Paper (Horie et al., 2021) HCI 8 2400
-P2 Book Chapter (Robinson and Rodrigues, 2022) Religion 6 1800
-P3 Academic Paper (Peirson-Smith, 2013) Social Science 4 2000
-P4 News Article (NYT, [n.d.]) Non-Fiction 8 1600
-P5 Laboratory protocol Engineering 4 1750
-P6 Academic Paper (Lin et al., 2018) HCI 4 2200
-All Wikipedia Article (wik, [n.d.]) World Affairs 6 4750
-Table 1. Documents used by participants in the workshop.
-Before commencing the study, we gave participants a verbal walk-through of the study’s purpose and provided a one-page reference containing sample visualizations.
-We also provided a set of sample prompts for ChatGPT to demonstrate the AI’s summarization and enhancement capabilities.
-The study followed a talk-aloud protocol and the authors took detailed notes of responses and observations.
-At the study’s conclusion, the authors collected all sketches authored by the participants.
-Upon study completion, all designs elicited from the participants were collated in a single Miro board.
-Four authors independently categorized and coded those results to identify common themes.
-We identified twenty distinct types of document enhancements, each with its specific usage, purpose, popularity, and the context in which it was used.
-Next, all authors discussed the initial design dimensions to refine the consistency and comprehensiveness of the categorization.
-Finally, four authors collaboratively re-coded the examples to produce a final set of design dimensions.
-Figure 2. Document enhancements produced by six participants in our formative design workshop. Highlights Refer to caption indicated the document enhancements implemented in our system.
-3.2. Elicited Designs and Findings
-We identified twenty unique classes of document enhancements, with each participant generating designs from on average three classes (SD = 1.41).
-We then identify five high-level categories, which include summarize, compare, augment, extract, and navigate.
-Figure 2 illustrates the summary of our study results.
-In this section, we describe each category in more detail.
-Category 1. Summarize
-One of the most common themes was to expedite the reading process by condensing lengthy text.
-All participants suggested text-based summaries with different scopes (document-level, page-level, or paragraph-level).
-Some participants also explored personalized summaries, such as “summarize this for a 5-year-old”, which help understand unfamiliar concepts and fit the unique comprehension levels and interests.
-Furthermore, many participants recommended the use of text summaries as a reference tool, enabling them to quickly review content from different parts of the document without losing focus on the current page.
-Category 2: Compare
-The most unique approach was the idea to dynamically change content representation by leveraging AI.
-For instance, participants envisioned transforming text and numerical data into more visual and spatial formats, such as tables, mind maps, or graphs.
-These alternative representations enable users to visually compare and analyze elements that would be less apparent in text.
-Moreover, these representations could be designed to be fluid and adaptive; for instance, tables or graphs could dynamically update, filter, or highlight based on user queries.
-Another idea was summarizing all of the chronological information into timelines.
-Such timelines could also automatically update by adding new content as the user reads through the document.
-This approach represents a unique design for AI-enabled augmentation, which emphasizes reshaping existing content into more comprehensible and comparable formats, offering an alternative method for active reading.
-Category 3. Augment
-All participants showed interest in enhancing their content with external data, such as maps and images.
-For instance, one participant envisioned a feature similar to Apple’s keyword lookup for quick reference to unfamiliar words.
-Such augmented content could come either from external sources (like images and maps), user-generated content (labels and sticky notes), or AI-generated content (converting 2D figures into 3D models).
-While such ideas have been explored in prior works (Li et al., 2019; Qian et al., 2022), participants also came up with unique AI-driven enhancements beyond simple augmentation.
-For instance, participants envisioned that by visualizing the importance, relevance, or number of citations for each reference in academic papers, they could quickly determine which paper to read next.
-Another unique concept involved synthesizing external information with the comparison feature.
-For example, when looking up the president of Ukraine, participants suggested not just showing a biography image and profile but also visualizing the organizational relationships through a network diagram with other individuals on a page by combining both external information and content on the page.
-Category 4. Extract
-Participants envisioned interacting with reading materials by extracting elements for persistent references, such as creating keyword lists or extracting figures like scrapbooks.
-They also considered the possibility of extracting citations from academic articles’ bibliographies or footnotes.
-For example, one participant envisioned a feature where inline citations could be expanded on paper documents.
-Furthermore, there was interest in pulling out significant quotes from books over the years, enabling them to accumulate and store these quotes for future use.
-Category 5. Navigate
-Participants proposed various methods to enhance navigation through dense documents.
-They suggested the creation of document reading progress indicators to clarify their location within a text.
-Additionally, there was interest in making a table of contents readily accessible to better comprehend the document’s structure.
-Another idea involved implementing collapsible headings, allowing users to hide sections they find irrelevant in extensive documents, thereby facilitating a more focused reading experience.
-One participant suggested making the document searchable in AR.
-By integrating extraction features, they imagined users verbally dictating a search term and subsequently viewing a list of pages along with a concise summary of that term within the document.
-Other Findings Related AI-Driven Augmented Reading
-Additional findings include the potential benefits of AI to support idea generation and critical thinking.
-For example, participants believed that AI-assisted enhancements could deepen understanding by presenting multiple perspectives or providing new insights to enhance critical thinking.
-On the other hand, participants also identified several possible challenges associated with AI-generated content, including the risk of inaccurate information, instability of output quality, overly lengthy responses, and the possibility of information overload leading to distraction.
-In summary, while some of the proposed features are beyond the scope of our implementation, these design ideas, along with insights, have informed the development of our system, which we describe next.
-4. System and Implementation
-Building on insights from our formative study, we developed RealitySummary, a mixed reality reading assistant that utilizes real-time OCR and a large language model.
-4.1. System Design
-Figure 3 summarizes the overview of our system design.
-The system first detects and captures the document with the headset’s camera and proceeds to extract text from printed or digital documents using optical character recognition (Google Cloud OCR).
-Given the extracted text, our system employs the GPT-4 large language model through OpenAI’s API for analyzing and extracting important information from the document.
-Additionally, we have a middle layer on the server that parses the text using spaCy’s named entity recognition module to extract a list of persons, locations, organizations, and keywords that are used for the web search via Google search API.
-The processed information is transformed into either a text summary or visual aid with a prepared design template, which we built based on our formative study (Figure 4).
-Figure 3. RealitySummary system architecture.
-4.1.1. Mixed Reality Headsets and Image Capture
-The system then embeds the text and visual output via the mixed reality headset.
-Since our pipeline is hardware independent, we deployed our system to Microsoft Hololens 2 and Apple Vision Pro.
-For HoloLens 2, we use an internal camera for image capture, Vuforia for image tracking, and handle content rendering with MRTK3 (Mixed Reality Toolkit 3) SDK [^2] Source:https://github.com/MicrosoftDocs/mixed-reality/blob/docs/mrtk-unity/mrtk3-overview/index.md.
-For Apple Vision Pro, we implement the same rendering and image tracking features using the RealityKit framework and Swift.
-Since the Vision Pro does not allow raw camera access, our prototype feeds the camera scene to a MacBook Air via AirPlay, then extracts images via screen capture.
-Additionally, we offer voice interaction support to modify or create these summary visualizations using Google’s speech-to-text API and a semantic similarity check between the transcript and feature descriptions to detect which feature the user intends to choose.
-4.1.2. Text Extraction
-We extract text content from physical or digital documents using OCR.
-We decided to use Google OCR in Google Cloud Vision API for accurate text extraction.
-OCR output also includes the position of each word, sentence, and paragraph, which helps embed the content for on-document augmentation.
-However, the OCR processing time is not fast enough for continuous capturing, we only first take a screenshot of the page every 5 seconds, hence, the system only needs to perform OCR once, as long as the page content does not change.
-We detect whether the user still reads the same page by comparing the new OCR results from the previous one.
-On top of the repeated capturing, the user can also manually activate the OCR with a voice command.
-4.1.3. Document Tracking
-We use Vuforia’s markerless image tracking or Vision Pro’s built-in image tracking features to track the document.
-The captured image is used for both OCR and document image segmentation, which runs in parallel.
-When the image is captured through the above step, we segment the document boundary with OpenCV rectangle detection.
-The system then uses the segmented image to generate a dynamic image target in Vuforia and Vision Pro.
-This allows RealitySummary to track essentially any document with sufficiently distinct visual features without attaching markers in advance.
-However, the accuracy of the system depends on various factors, such as lighting and tracked image pattern, the user can also manually place and move the content with hands.
-Recognizing the challenges of image tracking in diverse settings, our Vision Pro implementation also allows viewers to manually anchor content panels to the environment rather than the document.
-By taking advantage of the platform’s extremely robust spatial tracking, these environmentally-anchored panels (combined with its much higher-resolution displays) dramatically improve the readability of augmented content while reducing instability and jitter caused by moving image targets.
-4.1.4. Text Summarization
-After text extraction and document localization, the system extracts important information using NLP.
-We utilize spaCy and the GPT-4 large language model in the server.
-We use render AR content as a series of 2D canvases.
-We employ WebSockets communicate between mixed reality headset and the server with JSON format.
-We also store the OCR results, captured images, and GPT response history in Firebase database.
-We query multiple prompts to multiple GPT-4 API simultaneously to reduce the generation time.
-The specific implementation choices for each type of document enhancement are described in the next section.
-4.1.5. Speech Input Interaction
-The user can enable or disable different summarization panels simply by saying a trigger word like “show me the summary of this page” and “show me the table-of-contents”.
-Our system leverages Google’s speech-to-text API to convert spoken sentences into text.
-We leverage spaCy to measure the semantic similarity between the user’s spoken sentence and our predefined list of action descriptions.
-Beyond a simple summary activation, the user can also ask free-form, open-ended questions about the reading content by simply saying a question, which is directly sent to GPT-4 API along with the OCR results.
-4.2. Document Enhancements
-In this section, we describe our supported document enhancements.
-On top of the elicited designs that we derive from our formative study, our design is also inspired by the prior MR- and non-MR-based reading assistant, such as HoloDoc (Li et al., 2019), DuallyNoted (Qian et al., 2022), Marvista (Chen et al., 2022), and SummaryLens (Benharrak et al., 2022).
-While we acknowledge that individual design elements may not be significantly novel, we believe that our system contributes to the following two aspects: 1) our holistic framework and its evaluation provide a better understanding of document augmentation design space, 2) automated and real-time approach to creating on-demand content opens up a new potential of mixed reality assistant tools.
-Each of our document enhancement techniques relies on specific language model queries (more commonly known as prompts, which are documented in our supplemental material), often in combination with additional web services like image search for images and people bio information cards, and mapbox for generating maps.
-Figure LABEL:fig:documents_user_study shows various reading materials augmented with a variety of information through our system.
-Figure 4. RealitySummary document enhancement features. (All images captured live from HoloLens 2.)
-Summary
-The text summary is extracted directly by using a GPT-4 API call from the Node.js server.
-We send the entire text of the current page to GPT-4 and get a summary in return, which is then sent to be rendered on HoloLens.
-We simply prompt GPT-4 asking it to summarize the current text data extracted from the page.
-This technique is often used in prior summary-based reading assistants like Marvista (Chen et al., 2022).
-We limit the maximum word length of the summary to less than 150 words, as we learned the user struggles to digest lengthy text through the formative study.
-Comparison Table
-RealitySummary generates comparison tables based on the user’s prompt.
-We use GPT-4 to generate a comparison table between the two items along with contextual text from the document similar to generating text (see supplementary material for the prompt).
-Alternatively, while parsing the document, we prompt GPT-4 beforehand to generate possible comparison tables and store them locally through the session (and query directly from it when comparison action is detected).
-Timeline
-A timeline is a series of summarised events in a given text.
-When a user requests a timeline, we prompt GPT-4 to generate a series of events in the form of a timeline based on the user’s query and get a JSON with the relevant data.
-We then render this using Unity Canvas (Hololens) or HTML Canvas (Vision Pro) in a custom format to make it easy for the user to skim through.
-Keyword List
-The keyword list is generated using both GPT-4 and spaCy.
-SpaCy is used for keyword and phrase extraction, which also provides named entity such as date, location, and persons, which will be used for web search.
-We simply process the text by calling spaCy’s python wrapper using python-shell from our Node.js server and send the result back to the headset for rendering a list of keywords or highlighting them using word/phrase location, similar to keyword-based skimming techniques like ConceptEVA (Zhang et al., 2023).
-Summary Highlighting
-Summary Highlighting works very similarly to summaries, in that, the user asks to summarize or find more information about a specific phrase, sentence, or paragraph.
-Once the user wants information about content inside a document, we prompt GPT-4 for a brief summary and use word/phrase, sentence, or paragraph matching to find the location of the content in the document.
-We then highlight, draw arrows, or annotate based on OCR’s word bounding box results.
-Other features also leverage word and phrase highlighting to connect relevant information directly to the text in the document.
-Information Cards
-Inspired by AR-based annotations demonstrated in DigitalDesk (Wellner, 1991) and DuallyNoted (Qian et al., 2022), information cards, as the name suggests, contain information about keywords and people (image with a description), places (a map), and images about common keywords.
-To reduce the number of requests made to the LLM, we use Google Search API to extract information related to key terms, places, or people.
-We also use Google Image Search API to source pictures related to terms or phrases on demand.
-Moreover, we use mapbox API to generate maps.
-In the rare case that the API fails to provide a response, we query GPT-4 for more information.
-4.3. Technical Evaluation
-To characterize the technical performance of our prototype, we assessed the reliability of document tracking and overall system latency, as well as the tool’s ability to recognize documents of a variety of different types.
-To measure the system performances, we also systematically tested RealitySummary with (1) AR tracking, (2) OCR, and (3) summarization relevance/correctness using 20 additional documents (10 physical and 10 on-screen) with greater visual and technical diversity, including posters, newspapers, online articles, patents, magazines, thesis, books, blog posts, etc.
-4.3.1. Document Tracking
-The AR tracking analysis measured the system’s document tracking reliability with Microsoft Hololens 2 over 60 seconds each.
-We saw reliable tracking for 14 documents that included graphics (M=92%, SD=5.4% up-time).
-However, the 6 documents that included only text showed worse performance (M=64%, SD=3.7% up-time) due to their lack of distinct trackable features.
-AR tracking is not always very reliable, especially with text documents because they have fewer feature points to track.
-We used system logs to automatically analyze document tracking accuracy throughout the user study.
-We measured tracking loss whenever the system lost AR rendering or image tracking.
-We found our system’s tracking to be 72% accurate while tracking documents with only black-and-white text and 92% accurate with tracking documents with some colors and visuals (images, diagrams, charts, etc.).
-We also found substantial changes in tracking with different lighting conditions (better tracking with better lighting), document image tracking size (large image targets performed much better than smaller documents), and motion (sudden headset movements resulted in temporary tracking loss or flickering).
-4.3.2. Applicability of Physical or Digital Content
-OCR analysis examined the fraction of words correctly extracted from the documents.
-Across the 20 documents, we saw extremely high OCR accuracy (M= 97.9%, SD=5.0%).
-We tested our system’s generalizability and applicability with 20 different document types—newspapers, magazines, academic papers, books, legal documents, and technical documents.
-Two authors checked each summarization feature on all 20 documents manually to evaluate the applicability and accuracy of the content generated by the large language model.
-We found our system to be 94% accurate when generalized to other documents in providing correct information based on the context of the document.
-Summarization analysis verified the relevance and factual correctness of the initial 120 (20 $\times$× 6) summary cards generated by RealitySummary across these 20 documents.
-Two authors manually examined and verified all documents and cards, and found that 96.77% were both factually correct and relevant.
-4.3.3. Latency
-In our technical evaluation of latency, we measured the average latency of three different components: AR rendering, network communication, and API responses.
-Network communication represented our socket.io connection between the Node.js server and the headset client.
-The system recorded an average latency of 100 ms for rendering AR content in HoloLens and Vision Pro and an average latency of 40 ms for socket.io network connection, which is relatively low and suggests that the rendering and network process is responsive and efficient.
-On the other hand, API callbacks included two main components, namely external request to the GPT-4 model and pyShell request to our own NLP processing unit built with spaCy.
-Latency analysis included their API callbacks and wait-time.
-The average latency for GPT-4 was 2.4 s and NLP Module was 670 ms.
-5. Usability Study
-We evaluated our system with 12 participants via a think-aloud usability study.
-Our goal was to qualitatively assess our system’s usability, examine participants’ reactions to our document enhancements, and elicit new opportunities for MR reading assistants.
+We asked participants to complete a four-part design task where they read and graphically annotated both a document of their own and a Wikipedia article [^1] Source: https://en.wikipedia.org/wiki/Russian_invasion_of_Ukraine.
+After reading, participants brainstormed ideas for AI-powered augmentations to improve readability.
+In the first design task, participants annotated a physical printout of their own article, sketching ideas for possible enhancements, and then repeated the task with a printed copy of the Wikipedia article.
+In the second task, participants were provided with digital copies of the two documents on a virtual whiteboard (Miro) and asked to design a second set of document enhancements, this time with assistance from ChatGPT.
+Before the study began, we briefed participants, gave them a reference page of sample visualizations, and demonstrated prompts for ChatGPT that showcased its summarization and enhancement capabilities.
+Using a think-aloud protocol, the authors took detailed notes and collected all participant sketches.
+These designs were later collated on a Miro board, where four authors independently categorized and coded them, identifying 20 distinct types of document enhancements based on usage, purpose, popularity, and context.
+The authors then discussed and refined these categories, re-coding the examples to produce a final set of enhancement types  (Figure 3).
+Figure 3. In our elicitation study, participants identified a variety of potential summary and augmentation types based on sketching and ChatGPT exercises using diverse text documents.
+3.2. Results
+We identified five high-level categories: summarize, compare, augment, extract, and navigate.
+The summarize category focused on condensing text for quick reference, ranging from broad document summaries to personalized summaries for different comprehension levels.
+Summaries were seen as useful reference tools, allowing users to quickly review content without losing focus.
+The compare category emphasized transforming textual and numerical data into visual formats like tables, graphs, and timelines, which could dynamically update based on user interactions, making comparisons more intuitive.
+The augment category explored ways to enhance document content with external or AI-generated data, such as adding maps and images or converting 2D figures into 3D models.
+Advanced ideas included visualizing the importance of academic citations or integrating external information to provide richer context.
+The extract category focused on enabling users to pull out and store key elements, such as keywords, figures, or quotes, for future reference.
+Lastly, the navigate category proposed various methods to improve document navigation, such as progress indicators, accessible tables of contents, and using AR to make documents searchable.
+These insights, alongside considerations of the potential and challenges of AI-driven tools, informed the development of our system.
+4. Study 1 RS-Doc: RealitySummary First Prototype and Usability Study
+We implemented the first version of our system on a Microsoft Hololens 2 and conducted a preliminary user evaluation.
+In this section, we present initial insights from using on-demand, mixed reality document summarization and discuss the associated limitations and design implications.
+4.1. Implementation
+Figure 4. RS-Doc Hololens 2 architecture.
+Figure 5. Examples from our RS-Doc Hololens implementation.
+The system design for the HoloLens version involves capturing documents using the HoloLens 2’s internal camera and extracting text with Google Cloud OCR.
+The extracted text is processed using GPT-4 via OpenAI’s API to identify key information, while spaCy’s named entity recognition (NER) module assists in identifying important entities for further web searches.
+The system employs Vuforia for markerless image tracking and the Mixed Reality Toolkit 3 (MRTK3) SDK for content rendering.
+Users interact with the system through voice commands processed via Google’s speech-to-text API, allowing them to modify or query visualizations.
+The system captures images periodically for OCR, with an option for manual activation, and tracks documents by generating dynamic image targets, ensuring accurate document tracking and augmented content rendering.
+Informed by the elicitation study, the system supports several document enhancements aimed at improving the reading experience.
+These enhancements include text summarization, achieved by sending the current page’s text to GPT-4 for a concise summary, limited to 150 words for easier digestion.
+Comparison tables are generated on demand based on the document parsing, allowing for quick comparisons within the document.
+Timelines of events are created by prompting GPT-4 to generate a sequence of summarized events.
+A keyword list is extracted using GPT-4 and spaCy, with spaCy handling named entities for web searches and keyword highlighting.
+Summary highlighting allows users to obtain brief summaries or additional information on specific phrases, sentences, or paragraphs, which are then visually highlighted.
+Information cards provide further details on keywords, people, places, or images, leveraging Google Search, Image APIs, and Mapbox for generating maps.
+All these features are integrated into the HoloLens environment, offering real-time, interactive content within the mixed reality experience.
 4.2. Method
-We recruited participants (P1-P12, 6 male, 6 female; Ages = 19-35 years) through word-of-mouth and they were compensated with \$15 for their one-hour participation in the study.
-The study was conducted with Microsoft Hololens 2 in a lab environment.
-All the participants had verbal and written fluency in English.
-Before starting the study, the authors familiarized participants with AR and the Hololens 2, then introduced RealitySummary and each of the document enhancements described in the previous section.
-After this introduction, we provided three reading tasks using: 1) an A4 size magazine in physical format, 2) a digital document of the participant’s choice, and 3) an A3 size book.
-Participants’ digital documents included 3 books, 2 blog posts, 2-course assignment PDFs, 2 Wikipedia articles, 2 magazine articles, and 1 news article which they viewed on either a PC (4/12), tablet (5/12), or laptop (3/12).
-We asked participants to use RealitySummaryto aid with their reading, using whatever combination of enhancements best suited the current document.
-Throughout this 30-minute period, we encouraged participants to think aloud, documenting their experiences with the system.
-After the tasks, we conducted semi-structured interviews and asked them to complete a Likert scale questionnaire, including a System Usability Scale (SUS) protocol (Brooke, 1995).
-Together, the questionnaire and interview took roughly 20 minutes.
-Afterward, we conducted a thematic analysis by first transcribing the think-aloud sessions and interviews, then individually familiarizing ourselves with the content and organizing data into rough categories.
-Two authors then individually searched for broader groups and patterns and clustered these into themes.
-The writing team then iterated to further refine the identified themes.
+The usability study involved 12 participants (P1-P12, 6 male and 6 female, aged 19-35) recruited via word-of-mouth and compensated \$15 for their one-hour participation.
+The study aimed to evaluate the system’s usability, explore participants’ preferences for different document enhancement strategies, and identify limitations for future iterations with real-world use in mind.
+Conducted in a lab environment with Microsoft HoloLens 2, participants were first introduced to AR and the HoloLens 2, followed by an overview of the system and its document enhancements.
+They completed three reading tasks using different document formats (physical, digital, and a book) and utilized the system’s enhancements as they preferred.
+The participants’ digital documents included books, blog posts, course assignments, Wikipedia articles, magazine articles, and news articles viewed on various devices.
+Throughout the study, participants were encouraged to think aloud, and their experiences were documented.
+Following the tasks, semi-structured interviews and a Likert scale questionnaire, including a System Usability Scale (SUS), were conducted.
+Thematic analysis was performed on the transcriptions of the think-aloud sessions and interviews, resulting in the identification and refinement of key themes.
 4.3. Results
-Participants responded positively to RealitySummary (Figure 5), with most participants either agreeing or strongly agreeing that the the prototype was easy to use (10/12), easy to understand (11/12), and well-integrated with the task (10/12).
-Similarly, nearly all participants reported that the summary approaches provided by RealitySummary helped with comprehension (12/12), were relevant to the document content (11/12), improved comprehension speed (10/12), and augmented their reading experience (10/12).
+Overall, participants responded positively to RealitySummary, with most participants agreeing that the prototype was easy to use (M=4.3/5, SD=0.98), easy to understand (M=4.5/5, SD=0.67), and well-integrated with the task (M=4.3/5, SD=0.77) with 5-point Likert-scale questionnaire.
+Similarly, nearly all participants reported that the summary approaches provided by RealitySummary helped with comprehension (M=4.6/5, SD=0.514), were relevant to the document content (M=4.7/5, SD=0.65), improved comprehension speed (M=4.4/5, SD=0.79), and augmented their reading experience (M=4.7/5, SD=0.65).
 Finally, the System Usability Scale (SUS) returned an overall composite score of 71, suggesting a reasonable level of usability for an early-stage prototype.
-Figure 5. Usability study questionnaire responses.
-5.3. Findings
-Our analysis identified five themes related to participants’ experience comprehending and navigating documents, as well as their sense of focus and control.
-We also discuss our participants’ comparisons of the visual and text-based summarization techniques and layouts implemented in RealitySummary.
-Comprehension
-Document enhancements can improve readers’ sense of comprehension, and readers value the ability to see multiple types of enhancements simultaneously.
-All the users in the study felt that the system improved their reading comprehension.
-Out of the 12 participants surveyed, 11 reported that they were able to understand the document without actually reading it, solely by relying on the contextual information provided by RealitySummary.
-P9 even stated, “I didn’t even read the document but I understand what it is about.” Furthermore, multiple participants (P7, P8, P11, P12) argued that the combination of multiple features like summaries, timelines, keywords, images, and people cards provided them with a more comprehensive understanding of the document, allowing them to engage with it from different angles.
-This allowed them to gain multiple perspectives on the same topic, leading to a deeper comprehension of the material.
-All participants mentioned that RealitySummary helped them comprehend documents more quickly than their typical approaches (like visually scanning a physical book and manually searching for a related summary or timeline on a phone/PC).
-For example, after perusing the first magazine, P12 noted that “I’ve not read this magazine thoroughly but, it seems that these features can significantly help me understand what this document is about in less amount of time.”.
-Navigation
-Both structured overviews and in-context highlights provide useful entry points that can make it easier to navigate documents in a top-down fashion, even within individual pages.
+4.4. Insights
+Our analysis identified several themes related to participants’ experience in comprehending and navigating documents.
+We also explore participants’ comparisons of the visual and text-based summarization techniques as well as preferences for various augmentation techniques.
+Enable Multi-Angle Content Understanding for the Current Page
+RealitySummary’s document enhancements can significantly improve readers’ comprehension.
+Several participants (P7, P8, P11, P12) highlighted that features such as summaries, timelines, keywords, images, and people cards provided a more comprehensive understanding of the document, allowing them to engage with the content from multiple angles and perspectives.
+All participants mentioned that RealitySummary enabled them to grasp content more quickly than with their usual methods, such as visually scanning a physical book or manually searching for timelines on a phone or PC.
+For instance, after reviewing the first magazine, P12 remarked, I haven’t read this magazine thoroughly, but it seems these features can significantly help me understand what it’s about in less time.” Interestingly, several participants appreciated that the summaries and keywords were drawn only from the currently visible page rather than from the entire document.
+P5 reported, I like the page-by-page summary—having a summary for each page does help make it feel less dense compared to a summary of the whole document.” Similarly, P6 emphasized, “I like that it (the system) gives you keywords for each page and not for the whole document.
+Facilitate Efficient, Top-Down Navigation
+Both structured overviews and in-context highlights provide useful entry points that make it easier to navigate documents from a top-down perspective, even within individual pages.
 Broadly, participants spoke positively about how RealitySummary helped them navigate documents.
-Four participants (P3, P4, P6, P8) specifically reported using structured navigation elements like timelines to get a better understanding and structural flow of the whole text, with P4 noting that “the timeline helped me organize and navigate my thoughts very quickly while reading the text” (P4).
-Meanwhile, at least a third of participants (P5, P6, P9, P10) also mentioned using lower-level augmentations to help navigate the document— highlighting how the proximity of annotations and content allowed them to use keyword highlights and information cards as visual anchors.
-Additionally, several participants found the connection lines between highlighted words and their corresponding features to be a valuable tool for navigating the document—with P8 stating, “the words in the documents are highlighted […] and connected with all the features […] it is very helpful.
-It is also pretty helpful that I can see the word ‘Yosemite’ and where it is on the map around the text!”.
-In fact, ten different participants reported that the system allowed them to focus their reading by taking a “top-down” approach—first examining high-level ideas and terms surfaced by RealitySummary, then gradually transitioning into reading lower-level content where necessary.
-P8 highlighted that “[the enhancements] give you a foundation before you actually jump inside [the document]”.
-This allowed participants more leeway to prioritize their reading to “focus on things that you’re more unsure about rather than just kind of trying to find external information” (P1).
-Interestingly, several participants also reflected positively on the fact that summaries and keywords surfaced in our prototype were drawn only from the currently visible page, rather than the document as a whole.
-P5 reported that, “I like page by page summary—having a summary for each page does help make it feel less dense compared to a summary of the whole document”.
-Similarly P6 stressed that, “I like that it (system) gives you keywords for each page and not as a whole.
-Visual vs. Text Enhancements
-Both visual and text-based enhancements can benefit readers, with visuals providing supporting high-level overviews and providing context, and text summaries capturing more complex concepts.
-While our participants responded positively to all of the document enhancements in the prototype, they reported that they preferred text features when trying to gain a deeper understanding of a particular topic or content within the document.
-Conversely, participants generally felt that visual features improved their comprehension speed, increased engagement, and improved memorability.
-All 12 participants ranked the visual summarization techniques—including image cards, people bios, and maps—as helpful, with eight participants specifically highlighting the value of images and bio cards.
-Many participants (P1, P4, P6, P7, P9, P10, and P11) agreed that the visual augmentations helped them understand the document faster compared to just reading the text, noting that, “seeing the visuals helped me grasp what’s happening in the document right away” (P11).
-Several participants (P7, P8) also stressed the potential for visuals to increase their engagement with the text, with P7 stating, “I’m not a big fan of reading so having visuals are helping me assist in reading this”.
-These participants also highlighted the potential for visuals to increase memorability with P8 noting that for historical documents the people cards could “help them put a face to a name”.
-Similarly, P7 noted that while text-based summaries might contribute to greater understanding, “graphical stuff makes me remember it for a longer time”.
+Four participants (P3, P4, P6, P8) specifically mentioned using structured navigation elements like timelines to better understand the structure and flow of the text, with P4 noting that the timeline helped me organize and navigate my thoughts very quickly while reading the text.” In fact, ten participants reported that the system allowed them to focus their reading by taking a “top-down” approach—first examining the high-level ideas and terms, then transitioning into reading the lower-level content as needed.
+P8 highlighted, “[the enhancements] give you a foundation before you actually jump inside [the document].” This approach enabled participants to prioritize their reading, allowing them to “focus on things that you’re more unsure about” (P1).
+Support Quick Understanding with Visual Aids
+Both visual and text-based enhancements offer distinct benefits to readers—visual aids provide quick, high-level overviews, while text summaries capture more complex details.
+Some participants preferred text-based features for gaining a deeper understanding of specific topics, while others found that visual elements improved their comprehension speed, engagement, and retention.
+All 12 participants rated the visual summarization techniques as helpful, with eight specifically highlighting the usefulness of images and bio cards.
+Several participants (P1, P4, P6, P7, P9, P10, P11) noted that visual augmentations allowed them to grasp the document’s content more quickly compared to reading text alone, with P11 stating, “Seeing the visuals helped me grasp what’s happening in the document right away.” Participants (P7, P8) also emphasized that visuals enhanced their engagement with the material, with P7 commenting, I’m not a big fan of reading, so having visuals helps me engage with the content.” Additionally, participants highlighted the potential of visuals to improve memorability, with P8 noting that for historical documents, people cards helped put a face to a name.” Similarly, P7 observed that while text-based summaries aided understanding, “graphical elements help me remember it for a longer time.”
+4.5. Limitations
+Slow and Unreliable Document Tracking
+Document tracking enables unique features like keyword highlighting and around-document placement.
+However, the primary limitation of the current prototype is the instability of its document tracking.
+Overall, participants reported dissatisfaction with the system’s slow, inaccurate, and unreliable tracking capabilities.
+Unlike the OCR process, document tracking involves a heavier and slower workflow, requiring rectangle detection, image extraction, and image upload for Vuforia processing.
+Even with a perfect image target, tracking performance varies based on several factors, such as lighting conditions, hand occlusion, and viewing angles.
+Furthermore, while the tracking performed better for visually rich documents, it struggled with text-heavy documents, as it does not have any distance features, making consistent tracking difficult over time.
+Even when tracking functioned properly, participants frequently experienced disruptive flickering, which negatively affected the user experience.
+Additionally, the fixed placement of augmented information on documents posed a challenge for long-term use, as participants expressed the need for more flexibility in adjusting the location of the augmentations.
 Layouts and Visual Clutter
-Placing enhancements in the space around the document using mixed reality can successfully reduce the potential for visual clutter and distraction.
-Because they can sit alongside and even on top of document content, summaries, information cards, and other document enhancement have the potential to occlude or distract attention from the document.
-However, reducing the number of enhancements or moving them further from the text may limit their utility.
-In our case, several participants (P7, P10, P11) specifically highlighted how our around-document placement of the layouts integrated nicely with the documents themselves—calling them natural and intuitive to use.
-P10 reported that “[the system] is like a natural extension of the book.
-It is very easy to follow through.”.
-Similarly, four participants (P1, P2, P5, P11) praised the fact that the augmentations did not disrupt their visual focus on the document, with P5 noting that “it is not cutting into your field of view and it’s not invasive at all”.
-Moreover, 10 participants also remarked positively on the connecting lines linking document text to cards around around the document—indicating that they were easy to follow but did not impede readability.
-6. In-the-Wild User Study
-To explore additional real-world use cases, we also conducted an in-the-wild study with eleven participants.
-In contrast to the previous usability study, which focuses on standardized reading tasks in a lab setting, this study instead focused on examining the versatility and applicability of our system across uncontrolled environments and diverse reading materials.
-Our goal was to uncover unique advantages and potential use cases of employing mixed reality and AI-generated assistants in everyday settings.
-For this study, the system was deployed on an Apple Vision Pro, which provides considerably higher rendering quality than the Hololens 2.
-To accommodate diverse applications and reduce the obtrusiveness of additional content in everyday scenarios, the user interface was refined, limiting augmented content to a text box and an image slot.
-Figure 6. In-the-wild study.
+Several participants (P7, P10, P11) noted that around-document augmentation helped them focus on their reading but emphasized that the number of displayed cards and their placements should be adjusted based on the user’s needs and context.
+Displaying too many cards at once can lead to visual clutter and divert attention from the document itself.
+We observed that participants might become overly reliant on the generated content rather than engaging directly with the reading material.
+Whether positive or negative, 11 out of 12 participants reported that they could understand the document without actually reading it, relying entirely on the contextual information provided by RealitySummary.
+P9 even stated, “I didn’t even read the document, but I understand what it’s about.” Currently, the system automatically generates various types of cards when relevant, but allowing users to reduce the number of visible cards based on their preferences could help manage visual clutter and maintain focus.
+Fixed Prompt and Augmentation Types
+Another key issue was the inability of participants to query the generated summaries.
+The fixed prompt approach used for generating the different types of cards limited the system’s flexibility in addressing broader use cases and contexts.
+Several participants (P1, P3, P5) expressed frustration with this constraint, as they had specific questions they wanted to ask but were restricted to the preset summary types provided.
+While these different types of augmentations offered different benefits, those advantages were only realized when the generated content was relevant and accurate.
+Participants also noted that specific aids, like timelines, comparison tables, and keyword lists, were not always helpful or contextually appropriate.
+On the other hand, participants generally appreciated the broader usefulness of the general summaries and visual information cards, which were more adaptable to varying contexts.
+Applicability Beyond Controlled Environments
+Finally, the controlled nature of the study limits its ability to assess the broader applicability and real-world uses.
+While the initial findings are encouraging, they do not fully reflect how the system might perform in diverse, unstructured environments.
+As a result, transitioning to in-the-wild studies will be essential to understanding how document augmentations work in everyday scenarios and to gain insights from prolonged, regular use in real-world settings.
+In conclusion, a more refined version of the system, featuring flexible augmentation placement and query-able summaries, is essential for better evaluating its effectiveness in real-world scenarios.
+5. Study 2 RS-WILD: In-the-Wild Study with the Second Prototype
+Figure 6. Examples from our in-the-wild study.
+To explore real-world use cases, we conducted an in-the-wild study with eleven participants.
+Unlike the previous usability study, which focused on standardized reading tasks in a controlled lab setting, this study aimed to examine the versatility and applicability of our system in uncontrolled environments and with diverse reading materials.
+The goal was to uncover unique advantages and potential use cases for using mixed reality and AI-generated assistants in everyday settings.
+For this study, the system was deployed on an Apple Vision Pro, offering considerably higher rendering quality and a broader field of view compared to the Hololens 2.
+This contributes to better readability, particularly for text summarization, by using a less distractive transparent background.
+Based on the user feedback from the usability study, the user interface was simplified with a single text box and an image slot, which could accommodate any response type given the user’s custom query.
+This feature would better suit diverse applications and minimize the intrusiveness of additional content in daily use.
+5.1. Implementation
+The system design for the Vision Pro follows a similar approach to the Hololens but with key adjustments.
+Since the Vision Pro does not yet allow direct access to the camera view, we implemented a workaround for text extraction.
+First, we screencast the Vision Pro to an external iPad, which is then connected to a MacBook Air to stream the feed via QuickTime Player’s mirroring feature.
+A web-based interface on the Mac captures the live stream from the Vision Pro.
+Using the captured feed, we employed Google OCR to extract text, which was then processed by GPT-4 via the OpenAI API.
+The Vision Pro application, developed in Swift, featured a single movable card with an image/map slot.
+After querying the LLM, data was transmitted from the web-based interfaces to Vision Pro via Firebase’s real-time database.
+Additionally, a separate web interface with speech recognition was run on the user’s smartphone, allowing them to query the system via voice (using Web Speech Recognition) and ask follow-up questions.
+This enabled users to request specific summary types or submit open-ended queries about the text they were reading.
+The web-based interfaces, built with React.js and hosted on Glitch, were accessed by participants through their phones.
 5.2. Method
 We recruited 11 participants (W1-W11, 7 male and 4 female; ages 19-32 years) from the local community.
-They were compensated with \$25 for their participation in a 1.5-hour study, which consisted of a one-hour experiment and a half-an-hour interview.
-The study was conducted in various locations at a local university, including a library, cafeteria, store, hallway, classroom, and multiple lab settings.
-Before the experiments, we collected participants’ demographics and their reading habits, which were input into the system as the prior knowledge of ChatGPT.
-Participants familiarized themselves with Apple Vision Pro by doing the eye tracking calibration setup.
-Unlike structured tasks or prepared reading materials, participants were encouraged to utilize our system with any text content they encountered during the experiments, including documents, books, papers, posters, online news on mobile devices, and other materials in their environment.
-Throughout this one-hour experiment, we also encouraged participants to think aloud about their experiences.
-After the experiment session, we conducted semi-structured interviews, followed by a Likert scale questionnaire similar to the usability study.
+They were compensated \$25 for participating in a 1.5-hour study, which included a one-hour experiment and a 30-minute interview.
+The study took place in various locations at a local university, such as a library, cafeteria, store, hallway, classroom, and multiple labs.
+Before the experiment, we collected participants’ demographic information and reading habits, which were input into the system as prior knowledge for ChatGPT.
+Participants familiarized themselves with the Apple Vision Pro by completing the eye-tracking calibration setup.
+Unlike structured tasks or prepared reading materials, participants were encouraged to use the system with any text they encountered during the study, including documents, books, papers, posters, online news on mobile devices, and other materials in their environment.
+Throughout the one-hour experiment, participants were encouraged to think aloud about their experiences.
+After the experiment, we conducted semi-structured interviews, followed by a Likert scale questionnaire similar to the one used in the earlier usability study.
 Both the interview and questionnaire took approximately 30 minutes.
-As in the earlier usability study, we then analyzed participants’ questionnaire responses and reflections to identify high-level trends and opportunities for augmented reading tools.
+As in the previous usability study, we analyzed participants’ questionnaire responses and reflections to identify high-level trends and opportunities for augmented reading tools.
 5.3. Results
-We gathered feedback from participants on various aspects including the reading experience (immersiveness, understanding, distraction), perceptions of the system (content and query comprehension, trustworthiness, and privacy concerns), and overall usability (content placement and comfort using the MR headset).
-In general, most participant either agree or strongly agree that RealitySummary deepened their understanding of the reading content (7/11), RealitySummary demonstrated a deep understanding of the content they were reading (7/11).
-Participants reported feeling more immersed and interested while reading, with an average rating of 3.45 (SD = 1.04), and they perceived the system as effectively interpreting their queries (M = 3.27, SD = 1).
-Besides, their responses on the content placement within the system were also overall positive (M = 3.36, SD = 1.29).
-Participants had varied opinions on whether the system became distracting throughout the reading process (M = 2.55, SD = 1.37), and whether they tended to trust the accuracy and reliability of contents provided by the system (M = 2.82, SD = 1.25).
+We gathered feedback from participants on various aspects, including the reading experience (immersiveness, understanding, distraction), perceptions of the system (content and query comprehension, trustworthiness, and privacy concerns), and overall usability (content placement and comfort using the MR headset).
+In general, participants reported feeling more immersed and interested while reading (M=3.5/5, SD=1.04), and they perceived the system as effectively interpreting their queries (M=3.3/5, SD=1.00).
+Besides, their responses on the content placement within the system were also overall positive (M=3.4/5, SD=1.29).
+Figure 7. In-the-wild study questionnaire responses.
+Participants had varied opinions on whether the system became distracting throughout the reading process (M=2.6/5, SD=1.37) and whether they tended to trust the accuracy and reliability of contents provided by the system (M=2.8/5, SD=1.25).
 Privacy concerns were also noted by the majority (6/11), and discomfort using the headset was prevalent among the respondents (8/11).
 These findings are detailed in Figure 7.
-Figure 7. In-the-wild study questionnaire responses.
-6.3. Findings
+Figure 8. Snapshots captured from the Vision Pro during our in-the-wild study.
+5.4. Insights
 Observations from the in-the-wild deployment highlighted a wider range of use cases for augmented reading as well as some of the practical trade-offs associated with real-world use.
@@
 Notably, some participants even applied our system to their handwritten notes from lectures.
-Benefits of Always-On Assistant
-The exploration of such diverse reading materials can be attributed to the always-on capability of the mixed reality reading assistant.
-A key advantage of this always-on feature is its support for implicit interactions, offering a distinct user experience compared to ChatGPT on smartphones or laptops, which require explicit effort to engage.
-This implicit and automatic content capture greatly expands the range of applications, beyond what participants initially considered possible.
-The always-on capturing feature eliminated the need for manually supplying context, saving considerable time.
-For example, W9 experienced a smoother workflow when working on her business course project with RealitySummary compared to conventional screen-based methods.
-She was able to concentrate on her project, without having to switch between smartphones or laptops to familiarize ChatGPT with the context.
+Benefits of Always-On AI Assistant
+The exploration of such diverse reading materials can be attributed to the always-on capability of the mixed reality reading assistant, as opposed to screen-based interfaces.
+A key benefit of this feature is its support for implicit inputs, which dramatically lowers the barrier to asking questions of LLMs.
+This automatic, continuous content capture broadens the range of applications beyond what participants initially envisioned.
+For instance, W4 mentioned they had never thought to type a question into ChatGPT when looking at a product label.
+Even for common reading tasks, they appreciated that the summary was automatically generated and updated in the background as they flipped through pages in a library.
+Overall, the always-on, implicit assistant made it effortless to integrate AI into daily tasks, expanding potential use cases and drastically reducing the explicit effort required from users.
 Unique Affordances of Spatial and Tangible Interactions
-Beyond the always-on implicit assistance, participants appreciated the unique affordances offered by mixed reality interfaces, such as the opportunities for spatial and tangible exploration.
+Participants appreciated the unique affordances offered by mixed reality interfaces, such as the opportunities for spatial and tangible exploration.
 In a library setting, for example, participants could naturally navigate through bookshelves (spatial exploration), physically pick up a book (tangible interaction), and browse through its pages to automatically receive summaries (always-on assistance).
@@
 Other common usages included translations (W4, W5, W6), learning aids (W4, W8), and personalized content recommendations (W5, W10).
-For example, applied the system to his handwritten calculus notes to find answers to particular questions.
+For example, W8 applied the system to his handwritten calculus notes to find answers to particular questions.
 Beyond traditional reading applications, the AI’s capabilities supported practical decision-making; W4 compared nutritional information of products on store shelves and menu items to determine healthier options, while W5 used it for guidance on waste disposal and printer operations.
-The LLM’s ability to process a broad spectrum of inquiries significantly expanded the system’s applicability and utility.
+The LLM’s ability to process a broad spectrum of inquiries clearly expanded the system’s applicability and utility.
 Proactive Summaries vs On-Demand Question Answering
@@
 In contrast, on-demand question answering allowed for a more focused engagement on specific topics.
-For instance, W1 enjoyed the ”rapid-fire” questioning with a physics textbook, asking more questions than with traditional ChatGPT saying ”You have to describe it properly you have to specifically say what you’re asking about what you’re reading, so that kind of all of this abstracted in this system is like frees up your mind to just focus on the curiosity part Like rapid-fire questions and get rapid-fire answers”.
+For instance, W1 enjoyed the “rapid-fire” questioning with a physics textbook, asking more questions than with traditional ChatGPT saying ”You have to describe it properly you have to specifically say what you’re asking about what you’re reading, so that kind of all of this abstracted in this system is like frees up your mind to just focus on the curiosity part Like rapid-fire questions and get rapid-fire answers”.
 W2 valued how questioning was woven into reading, preserving thought flow without the disruption of toggling between reading and ChatGPT interfaces.
 Overall, participants favored the on-demand feature due to its applicability to diverse situations, yet many appreciated having both options for different purposes: proactive features for implicit skimming and on-demand for addressing more specific questions.
-Trust of AI-Generated Content
-Many participants approached the system’s responses with caution, often verifying the answers with the reading material.
-They highlighted the necessity of consulting additional sources for in-depth or technical readings when depending on the system.
-There was a general agreement among participants about their reluctance to fully trust an always-on reading assistant in their daily activities.
-Conversely, we noted some intriguing discussions regarding how participants perceive our system, either as a mere tool or more like a personal companion.
-While most participants engaged with the system as a conventional chatbot, some regarded it more as a personal companion than a mere tool.
-Notably, W3 and W5 described the system as a friend for discussing and inquiring about their readings.
-The system’s ability to provide always-on immediate responses facilitates this personification.
 Potential Risks of AI-Driven Augmented Reading
@@
 Overall, participants wanted the flexibility to choose when and how the system would be used to mitigate these risks.
-Possible Future Improvements
-A desired feature among participants was the system’s ability to recognize not only text but also images.
-For instance, W1 faced limitations when attempting to inquire about a diagram on a page, which the system could not recognize.
-Similarly, they anticipated broader usage scenarios beyond text recognition, such as generating cooking recipes from fridge ingredients or providing information on real-world objects.
-We anticipate that enhancing the system with multi-modal LLM capabilities could address such needs.
-Additionally, there was interest in a “chat history” feature to recall previous interactions with the system, highlighting a need for temporal relevance in user-system interactions.
-Finally, participants suggested they would be more inclined to use such a system long-term if hardware constraints were overcome.
-7. Limitations and Future Work
-Our experiences building RealitySummary and the insights from both our formative workshop and study highlight a range of promising research directions for MR document enhancement.
-Technical Limitations
-We first recognize that our system exhibits numerous technical challenges that need addressing in future enhancements.
-Concerning AR tracking difficulties, our evaluation found that AR tracking becomes unstable with text-only documents.
-This instability is exacerbated in low ambient light, diminishing the tracking’s effectiveness and reliability.
-Such issues raise questions about the AR technology’s adaptability and dependability for daily use.
-Additionally, OCR bottlenecks and image tracking errors introduced complications related to processing speeds, data transmission efficiency, and reliance on the quality of images.
-Given the variance in document quality, the performance of OCR and image tracking can fluctuate, leading to possible inaccuracies and delays for different types of documents.
-Lastly, LLM generated errors show areas for improvement, particularly in situations where the text is ambiguous or highly specialized.
-From Page to Documents
-Currently, the system cannot display summaries, annotations, or relationships spanning multiple pages.
-As the system’s inputs are limited to visible pages, content not yet encountered cannot be referenced.
-Future research might explore a hybrid approach.
-For example, if a digital version of the document is available online, the system could access that content to improve summarization.
-These enhancements could include the types of navigation interactions (such as dynamic tables of contents or visual reading histories) highlighted by participants in our formative study.
-Connecting extracted text with complete digital versions of documents could enable new visual search, preview, and comparison tools that build upon the interactions seen in prior prototypes like HoloDoc (Li et al., 2019).
-By merging the capabilities of augmented reality with extensive digital information repositories, we believe this gap can be bridged in the future.
-Beyond Text-Based Summaries
-While our current work has focused on examining text-based summarizations using a large language model, our formative study (Section 3.2) also highlights the potential for visual summarization approaches, including automated mind-mapping and charting tools.
-Techniques like these could support more efficient search, comparison, and integration, building upon word-scale (Goffin et al., 2014) and micro visualization techniques (Parnow and Dörk, 2015) designed to reveal data in and around existing text.
-Integrated paragraph-level summary visualizations, such as those introduced in Caragraphs (Masson et al., 2023), are especially promising.
-They would be well-suited to the kinds of placements around documents that mixed reality reading tools enable.
-Multimodal Inputs
-The improving multimodal input capabilities of mixed reality devices and LLMs, such as visual and audio input, eye tracking, more precise hand tracking, and richer gestural interactions, create additional opportunities for dynamic reading assistance.
-For instance, the eye-tracking capabilities of modern MR hardware could be used to estimate gaze and incorporate that information to tailor summaries according to the viewer’s current task or area of interest.
-If the system also captures verbal communication in addition to written text, like in RealityTalk (Liao et al., 2022), we could use this for conversation and discussion assistance.
-In this scenario, summaries would reflect elements of the discussion.
-Combining speech and physical manipulation could allow users to dynamically request, expand, adapt, or dismiss summaries and support interactions in a wider range of contexts and environments (Oviatt, 2007).
-Finally, the use of multimodal LLMs to receive graphical, environmental, and gestural input have the potential to broaden these systems’ applicability.
-Long-Term Evaluation
-While it highlights a variety of opportunities for augmented reading tools, our in-the-wild study is still a short-term evaluation.
-Long-term studies, in which participants interact with these systems for weeks or even months will almost certainly reveal deeper insights and challenges.
-Long-term studies are challenging given the current limitations of mixed reality hardware like the Hololens 2 and Vision Pro, which are not comfortable enough for everyday use.
-We anticipate that the continuous improvement of the headset hardware will make long-term studies feasible in the future.
+Trust of AI-Generated Content
+Many participants approached the system’s responses with caution, often verifying the answers with the reading material.
+They highlighted the necessity of consulting additional sources for in-depth or technical readings when depending on the system.
+There was a general agreement among participants about their reluctance to fully trust an always-on reading assistant in their daily activities.
+Conversely, we noted some intriguing discussions regarding how participants perceive our system, either as a mere tool or more like a personal companion.
+While most participants engaged with the system as a conventional chatbot, some regarded it more as a personal companion than a mere tool.
+Notably, W3 and W5 described the system as a friend for discussing and inquiring about their readings.
+The system’s ability to provide always-on, immediate responses facilitates this personification.
+5.5. Limitations
+Overall, the system design of the second prototype enhanced its applicability and stability for real-world use.
+Compared to the first prototype on the Hololens 2, the Vision Pro’s superior rendering quality and broader field of view provided a better user experience.
+Additionally, the simple system design greatly improved usability despite the absence of document tracking.
+While the second prototype showed great potential, we also identified several feature-level limitations, along with inherent study-related constraints.
+Lack of History and Recall Support
+While the minimalist design effectively reduces visual clutter and conserves screen space, it also limits the system’s overall utility.
+Although users appreciated the clean and simple interface, many expressed frustration with the lack of options to save or revisit previously generated responses.
+Several participants suggested that a history feature or the ability to store responses for future reference would substantially enhance the user experience.
+Challenges with Screen Viewing for Extended Use
+Many participants raised concerns about the difficulty of clearly viewing phone and computer screens through the Vision Pro without dramatically enlarging the text.
+While this worked for short-term tasks like quickly reviewing an assignment, participants found it frustrating for longer viewing periods, particularly when using laptops.
+Since much of their work is done on computer screens, this limitation poses a challenge for assessing the system’s effectiveness in supporting realistic, work-related tasks.
+Need for a Longer-Term Study
+Our in-the-wild study was inherently limited by its short-term, public use — with each session lasting only one hour per participant.
+As a result, participants did not have the opportunity to fully engage with the system in their personal workflows.
+For instance, due to privacy concerns related to screen recording, participants rarely used our system for more personalized content, such as emails, messages, or calendars.
+Furthermore, since participants were encouraged to explore various scenarios, they tended to experiment broadly rather than focusing on any specific task in depth.
+While the in-the-wild study provided valuable insights into real-world use cases, it only captured short-term interactions in public spaces.
+There is still much to learn about the sustained, regular use of reading assistants in daily life, particularly in how users interact with information over time, revisit previous queries, or support ongoing tasks.
+6. Study 3 RS-Diary: Diary Study with the Third Prototype
+To explore the long-term use cases, we developed a third iteration, incorporating feedback from the in-the-wild study.
+This version introduced a multiple responses feature and added the ability to view or mirror the content on an external device, addressing the readability issues for the long-term use of the Vision Pro.
+Figure 9. Snapshots captured from the Vision Pro during our diary study.
+6.1. Implementation
+The final version of RealitySummary builds on the earlier Vision Pro implementation with two key improvements: 1) Users can now add multiple draggable cards on-demand by either making a new query or clicking an “add” button.
+User can spatially anchor these cards anywhere in their immediate environment and can also re-select cards to update them with new prompts or text from the scene.
+These cards are stored as JSON objects in a real-time database.
+2) The system allows users to mirror their device screen within the Vision Pro for high-resolution viewing.
+This functionality — implemented through a WebRTC video call or via Mac Virtual Display on the Vision Pro — lets users view their external devices at full resolution in the MR environment and also allows RealitySummary to more accurately capture and process on-screen content.
+6.2. Method
+Five authors (A1-A5) used the system over the course of 3 weeks, integrating it into their day-to-day activities and accumulating 30 total hours of use.
+Due to physical fatigue and battery limitations of the Vision Pro, each session lasted less than 2 hours.
+The authors independently documented their use cases, insights, frustrations, thoughts, and reflections while using RealitySummary, which we later thematically coded.
+During the study, each participating author recorded their observations via a phone microphone, following a think-aloud process.
+Additionally, the first author conducted follow-up interviews to gain deeper insights into their experiences.
+6.3. Results
+Throughout the diary study, participants used the system in various settings, including labs, kitchens, outdoor areas, and offices, though the majority of usage took place at desks in work environments.
+They engaged with a wide range of reading materials, such as academic papers, news articles, code, emails, calendars, Google Scholar, classroom assignments, lecture slides, textbooks, and YouTube videos.
+Participants quickly adapted the system to fit a variety of personal, academic, and technical tasks.
+For example, A4 used the assistant not only to summarize email content but also to gather higher-level insights, such as identifying how many emails were sent by a particular sender.
+A2 utilized the system to support programming assignments, where they used the system to break down complex coding problems into manageable steps and provided detailed explanations of specific code snippets.
+A2 also leveraged the system for learning aids, such as generating to-do lists and quiz questions based on problem descriptions.
+The ability to spatially organize tasks and concepts allowed him to effectively manage his workload, maintain focus, and dynamically track his progress.
+Additionally, in his language learning efforts, A2 used the assistant to generate contextual vocabulary lists from Japanese reading materials, which helped him focus on the most relevant and frequently used words.
+A5 employed it to extract a step-by-step sequence of keyboard commands from an article on rebooting a MacBook.
+Overall, the diary study highlights the system’s flexibility in supporting diverse tasks, from email management and coding to technical troubleshooting, language learning, and academic research.
+6.4. Insights
+From these experiences, participants shared various insights, frustrations, and potential future improvements for the current prototype.
+Our diary study also provided insight into the use of an always-on reading assistant in more realistic work-oriented settings.
+Quick and Simple Summary Responses Meet Many Immediate Needs
+Participants valued the system’s ability to deliver quick, concise responses, which were often sufficient for immediate clarifications.
+A2, for instance, appreciated how the system provided brief answers when they needed a rapid comparison of philosophical concepts, like distinguishing between act and rule utilitarianism.
+In many cases, just a high-level summary captured through manual input was enough to satisfy their needs.
+A4 found this method to be a useful shortcut, as pressing a button was faster and less disruptive than speaking and often provided exactly the information they were seeking.
+These simple, proactive summaries were particularly beneficial for tasks that required quick comprehension and minimal interruption.
+Minimal Context Switching Enhances Focus and Flow
+Participants appreciated the integration of LLM into the MR interfaces, in part, because it helped them minimize context switching, allowing them to maintain greater focus.
+The system’s automatic text extraction feature enabled participants to stay in a flow state.
+For example, A2 noted that the interface allowed them to concentrate on their work without the mental disruption of switching between tasks or manually inputting context, even when using a computer screen.
+This seamless workflow kept participants engaged, enhancing productivity and reducing cognitive load.
+A1 highlighted how the system functioned like a “teacher” by offering real-time assistance without requiring them to step away from their tasks.
+Participants also remarked that on-demand support boosted their efficiency during complex tasks like coding or writing, where focus is critical.
+Using Generated Cards as Peripheral References
+Participants highlighted the advantages of peripheral output modalities.
+For instance, A2 used the system as a reading aid while learning Japanese.
+Instead of frequently opening a dictionary, which disrupted the reading flow, they generated vocabulary flashcards around the document.
+This provided the needed language support without interrupting the reading and learning process.
+A4 also noted that they would often cluster and group these generated cards thematically to the sides of the display for quick reference and then delete them once they were no longer needed.
+Similarly, A1 used the system to “pin” generated code versions and key concepts around their workspace, allowing for easy access without breaking focus and enabling them to quickly revisit important points as needed.
+Spatial Organization Improves Recall and Multitasking
+A key finding from the study was how participants leveraged the system’s spatial organization features, similar to arranging post-it notes on a board.
+Both A1 and A2 utilized this capability to group related concepts together.
+For example, A2 clustered cards on CPU architecture, placing related topics like buses and registers nearby, much like organizing post-its by theme.
+This spatial arrangement enabled quick access and improved recall, as users could visually track where specific information was placed.
+Similarly, A1 pinned cards documenting different problem-solving methods across their workspace, ensuring easy access when needed.
+This post-it-like organization supported memory retention and task management by allowing users to arrange information in a way that aligned with their thought processes.
+6.5. Limitations
+Challenges Controlling What to Capture
+Currently, the system captures the user’s entire field of view, but participants expressed a desire for more control over what is captured.
+For example, A1 struggled to get the system to ignore personal emails displayed next to the article he was reading, leading to summaries that inadvertently included private information.
+Similarly, A2 wanted to focus on a specific code snippet to obtain more accurate results while debugging.
+Additionally, the lack of visual feedback left users uncertain about what OCR content or context was being processed in each request.
+To address these issues, the system could offer enhanced visual feedback and allow users to control specific areas of the screen, such as through gaze interaction, which would help mitigate privacy concerns and prevent the inclusion of irrelevant or personal content.
+Cumbersome Re-Prompting Process for Common and Repeated Tasks
+Participants frequently found themselves re-prompting similar queries for routine tasks, such as requesting bullet point summaries or filtering by specific criteria.
+A potential improvement could be the addition of dynamic buttons that save commonly used prompt types, allowing users to easily reuse them later.
+Additionally, participants expressed frustration with testing out lots of different prompts to generate a complex output, such as aggregating the Google Scholar page into a timeline.
+Saving or sharing the same specific instructions for repeated tasks might mitigate these problems.
+Managing Growing Numbers of Summaries
+In the first prototype, participants found it distracting to have multiple summary cards anchored to the image in their workspace.
+However, they did not find this frustrating when they had control over the placement and could easily delete the cards.
+Additionally, participants noted that the system could sometimes generate an overwhelming number of cards.
+For instance, A2 mentioned that while analyzing a poem, the system created a card for each sentence and its meaning, making it difficult to manage information overload.
+This feedback suggests that there is potential value in exploring new interactive and semi-automated methods for efficiently sorting, snapping, organizing, filtering, and removing these cards.
+Such improvements could enhance the usability of the system and help users better manage the content generated.
+7. Lessons Learned from Three Studies
+Drawing on results from across our prototyping and evaluations, we close by showcasing the unique benefits of combining MR and AI tools, enumerating potential use cases, and highlighting opportunities for future research.
+7.1. Unique Benefits of Combining MR and AI
+We identified two unique benefits of combining MR with AI: 1) implicit input and 2) peripheral output.
+Always-On Implicit Input
+The always-on camera in MR enables implicit input, distinguishing it from the explicit input required in desktop interfaces, such as opening an app, typing queries, or copy-pasting contextual text.
+Participants in in-the-wild and diary studies appreciated the system’s ability to capture information without requiring explicit effort.
+This always-on input allows for hands-free exploration in both spatial and temporal contexts, such as scanning entire bookshelves as a query or utilizing the history of flipped pages.
+Implicit input reduces cognitive load and minimizes context switching, as users are not required to actively specify what information should be used for the LLM’s context.
+Peripheral and Background Output
+We found that MR’s spatial output allows information to be presented in a peripheral, background manner.
+In contrast, other output modalities, such as audio or screen-based displays, often require the user’s foreground attention to consume generated content, disrupting focus and interrupting workflow.
+With MR and peripheral outputs, however, AI-generated content becomes less intrusive, as it can be easily ignored if not needed, simply by choosing not to look at it.
+For example, participants in the diary study often appreciated this feature for learning aids, such as understanding textbooks or accessing always-on vocabulary references when learning a foreign language.
+This peripheral output is a unique aspect of the integration of MR and AI.
+We believe this capability would promote more calm human-AI interactions in the future, reducing disruptions to focused attention by combining implicit input with unobtrusive background support.
+7.2. Emergent Use Cases for Mixed Reality Text Augmentation
+Both the in-the-wild and diary studies revealed a variety of interesting emergent use cases.
+To provide a more comprehensive view of how users engage with these systems, we have categorized these into seven groups:
+Summarizing Content
+The primary use case was to summarize content.
+This was the most commonly observed behavior across various contexts.
+Participants summarized a wide range of texts, including research papers, blog posts, news articles, books, and instructional guides.
+Extracting Information
+Participants also utilized the system to extract key information, which can be translated into a more consumable format.
+This aligns with design elicitation results, which suggest tables of contents, keyword lists, timelines, and comparison tables.
+For example, one participant extracted vocabulary flashcards to assist with foreign language reading, while another participant created a table of keyboard shortcut commands from a troubleshooting guide to use as a reference.
+In these examples, the consumable format, such as tables or vocabulary lists, serves as helpful references for their tasks.
+Querying External Sources
+Participants frequently used extracted text as a query to retrieve additional information from external sources.
+Echoing the information types identified in our elicitation study, participants queried specific names or keywords from their reading materials.
+For instance, one participant retrieved details about a player, team, and standings from a sports-related article, while another fact-checked claims on-demand when reading news articles.
+Asking for Suggestions
+Beyond the above use cases identified in the elicitation study, participants frequently asked for recommendations or suggestions based on the current context.
+For instance, users asked for book recommendations when looking at a bookshelf.
+Beyond the reading context, one participant asked for advice on which trash bin can be used based on disposal instructions, while others inquired about healthier menu options or products.
+They also asked for recommendations from external sources.
+For instance, participants asked for suggestions for relevant research papers when reading academic articles.
+Aggregating Information
+An interesting use case involved using the system to aggregate high-level information.
+Unlike simple information extraction, aggregation often involves operations like filtering and counting.
+For example, some participants used the system to count emails based on specific filters, while others filtered and counted first-author papers from a researcher’s Google Scholar page.
+This allowed users to perform quick, approximate queries that would normally require more complex searches.
+Transforming Content
+Participants also used the system to transform their reading content into interactive, actionable outputs.
+For instance, some participants applied our system to generate quizzes from class notes or textbooks, while others decomposed high-level assignments into step-by-step to-do lists.
+One participant asked the system to create reminder cards that remained spatially persistent.
+Participants also use our system to translate from one language to another.
+Unlike information extraction, these tasks involved converting content into actionable, interactive outputs.
+Generating New Content
+Finally, participants generated new content based on existing text.
+In one example, a participant requested multiple variations of a written paragraph, which they then organized spatially.
+Others used the system to create code based on instructions or to debug existing code snippets.
+These emergent use cases highlight novel ways of interacting with AI-generated information, allowing users to engage with individual snippets and convert them into dynamic outputs.
+They also illustrate how participants adapted the assistant to handle different types of content, which suggests a broad space of possible designs for such systems.
+7.3. Future Opportunities and Design Implications
+As both MR and LLM technologies continue to evolve rapidly, considerable opportunities exist for tools that more richly integrate them.
+Based on our three studies, we identify several key opportunities for future research and design.
+Enabling Greater Control Over Capture
+While the system’s implicit input offers significant benefits, it also raises privacy concerns due to the automatic capturing without the user’s constant awareness.
+To address this, future designs should give users control over what data is captured.
+As noted in the diary study, participants expressed a desire for more transparency and control over the captured content.
+One approach could involve providing visual feedback, allowing users to see and select the specific areas or information being captured.
+Future systems should balance the benefits of implicit input with privacy considerations by offering user control and visual feedback.
+Improving Organization and Filtering of Interactive Cards
+Participants expressed the need for more efficient ways to manage the growing number of spatially organized content cards.
+While some diary study participants manually arranged these cards around documents, others suggested more advanced organizational tools.
+Features like sorting, snapping, filtering, or automatically organizing cards could simplify workflows and reduce clutter.
+Such capabilities would be particularly beneficial for long-term tasks, where the accumulation of content can become overwhelming.
+Integrating these organizational tools would enhance usability, supporting users in managing dynamic and task-specific information over time, which enables better workspace management for daily tasks.
+Balancing Between Reusable Prompts and On-Demand Custom Queries
+Our studies highlighted the importance of accommodating both reusable prompts and custom queries.
+Each has its advantages: prepared prompts can generate complex outputs such as timelines or comparison tables, while custom queries offer users flexibility in generating content.
+The diary study participants expressed a desire for dynamic interfaces that could quickly reuse and adapt prompts to different contexts.
+For example, some suggested having customizable buttons that automatically generate summaries or lists in a consistent style.
+Others mentioned the need for a more efficient way to reuse queries across tasks, such as repeatedly generating lists or comparisons for class topics.
+This feedback highlights the importance of designing adaptive interfaces that allow users to seamlessly reuse and modify their prompts, improving efficiency and user satisfaction.
+Facilitating Seamless Transitions Between Screen and Space
+Participants also valued the ability to drag content cards from the system into their computer, such as pasting generated answers directly into coding tasks.
+Some participants also envisioned this feature being useful for writing tasks, allowing for a smooth transfer of generated content.
+Future systems should explore leveraging cross-device interactions between MR and other devices to facilitate seamless transitions, enhancing the workflow between screen-based and spatial interactions.
+Expanding Beyond Text Input to Capture Images and Everyday Objects
+Participants expressed a strong desire for the system to recognize not only text but also images.
+For example, the in-the-wild study revealed the limitations of inquiring about a diagram that the system could not recognize.
+Users also anticipated broader applications beyond text recognition, such as generating recipes based on ingredients in the fridge or providing information on real-world objects.
+Enhancing the system with multi-modal LLM capabilities could address these needs, expanding its utility beyond text-based interactions.
+Personalization Through Always-On History
+Lastly, participants showed interest in an “always-on” history feature that would allow them to recall previous interactions with the system, creating a more personalized experience.
+They speculated that such a system could eventually serve as a personalized assistant, continuously capturing and understanding context through an always-on camera.
+Although current hardware limitations, such as battery life and physical fatigue, prevent continuous long-term use, we believe future MR interfaces will seamlessly integrate with AI to provide both context-aware and personalized assistance based on users’ daily activity histories.
 8. Conclusion
-RealitySummary is the first example of an on-demand MR reading assistant that uses OCR and an LLM to generate reading aids for any physical or digital document.
-We hope our study opens up new opportunities for on-demand and natural language-powered reading assistance and inspires the HCI community to continue exploring such systems to realize the full potential of intelligent reading tools.
+RealitySummary is the first on-demand MR reading assistant that combines OCR and LLM technologies to generate spatially-situated on-demand reading aids for both physical and digital documents.
+Through our empirical studies, we aimed to highlight new possibilities for on-demand, AI-powered reading support.
+This paper presented an iterative development of the system, along with key findings and lessons learned from three studies.
+We hope our work will inspire the HCI community to further explore the potential of integrating MR and AI in future applications.

### v3

@@
 Abstract.
-Large Language Models (LLMs) are gaining popularity as tools for reading and summarization aids.
-However, little is known about their potential benefits when integrated with mixed reality (MR) interfaces to support everyday reading assistants.
-We developed RealitySummary, an MR reading assistant that seamlessly integrates LLMs with always-on camera access, OCR-based text extraction, and augmented spatial and visual responses in MR interfaces.
-Developed iteratively, RealitySummary evolved across three versions, each shaped by user feedback and reflective analysis: 1) a preliminary user study to understand user perceptions (N=12), 2) an in-the-wild deployment to explore real-world usage (N=11), and 3) a diary study to capture insights from real-world work contexts (N=5).
-Our findings highlight the unique advantages of combining AI and MR, including an always-on implicit assistant, minimal context switching, and spatial affordances, demonstrating significant potential for future LLM-MR interfaces beyond traditional screen-based interactions.
+Large Language Models (LLMs) are gaining popularity as reading and summarization aids.
+However, little is known about their potential benefits when integrated with mixed reality (MR) interfaces to support everyday reading.
+In this iterative investigation, we developed RealitySummary  [^1] Project Page: https://adigunturu.github.io/RealitySummary_SUI25/, an MR reading assistant that seamlessly integrates LLMs with always-on camera access, OCR-based text extraction, and augmented spatial and visual responses.
+Developed iteratively, RealitySummary evolved across three versions, each shaped by user feedback and reflective analysis: 1) a preliminary user study to understand reader perceptions (N=12), 2) an in-the-wild deployment to explore real-world usage (N=11), and 3) a diary study to capture insights from real-world work contexts (N=5).
+Our empirical studies’ findings highlight the unique advantages of combining AI and MR, including always-on implicit assistance, long-term temporal history, minimal context switching, and spatial affordances, demonstrating significant potential for future LLM-MR interfaces beyond traditional screen-based interactions.
 Mixed Reality; Large Language Models; Augmented Reading; In-the-Wild Study; Diary Study
-journalyear: 2024copyright: acmlicensedconference: arXiv; September, 2024; $c$s: Human-centered computing Mixed / augmented reality
+journalyear: 2025copyright: ccconference: ACM Symposium on Spatial User Interaction; November 10–11, 2025; Montreal, QC, Canadabooktitle: ACM Symposium on Spatial User Interaction (SUI ’25), November 10–11, 2025, Montreal, QC, Canadadoi: 10.1145/3694907.3765933isbn: 979-8-4007-1259-3/2025/11ccs: Human-centered computing Mixed / augmented reality
 Figure 1.
-We developed RealitySummary, an on-demand mixed reality reading assistant, over three iterations and conducted three different studies.
-Through this evolution, we highlight three key findings: 1) unique benefits of MR and LLMs, 2) emergent use cases of MR reading assistants, and 3) design implications for mixed reality reading assistants.
+RealitySummary is an on-demand mixed reality reading assistant developed over three iterations, with each version extensively shaped by user feedback and reflection.
+Through this evolution, we highlight key insights and two notable findings on integrating LLMs in MR interfaces: 1) the unique benefits of MR + LLMs, and 2) emergent use cases for MR reading assistants.
 1. Related Work
@@
 Examples include HoloDoc (Li et al., 2019), Replicate and Resuse (Gupta et al., 2020), and PaperTrail (Rajaram and Nebeling, 2022), all of which laid the groundwork for augmenting physical paper by superimposing dynamic virtual content onto printed documents.
-Prior works have demonstrated these concepts on different platforms.
-For example, Pacer (Liao et al., 2010), MagPad (Xu et al., 2015), Tecahable Reality (Monteiro et al., 2023), and Dually Noted (Qian et al., 2022) use mobile AR for text augmentation, whereas Replicate and Resuse (Gupta et al., 2020) and Wiki-TUI (Wu et al., 2007) leverage head-mounted displays to overlay information.
-Other examples such as DigitalDesk (Wellner, 1991), DocuDesk (Everitt et al., 2008), Affinity Lens (Subramonyam et al., 2019), Matulic et al.
-(Matulic and Norrie, 2013; Matulic et al., 2013), and QOOK (Zhao et al., 2014) use projection mapping for document augmentation.
+Prior work has demonstrated these concepts on different platforms.
+For example, Pacer (Liao et al., 2010), MagPad (Xu et al., 2015), Teachable Reality (Monteiro et al., 2023), and Dually Noted (Qian et al., 2022) use mobile AR for text augmentation, whereas Replicate and Resuse (Gupta et al., 2020) and Wiki-TUI (Wu et al., 2007) leverage head-mounted displays to overlay information.
+Other examples such as DigitalDesk (Wellner, 1991), DocuDesk (Everitt et al., 2008), Matulic et al.’s pen and touch systems (Matulic and Norrie, 2013; Matulic et al., 2013), and QOOK (Zhao et al., 2014) use projection mapping for document augmentation.
+Finally, cAR (Hincapié-Ramos et al., 2014) leveraged transparent displays for document augmentation to support active reading.
 At their core, such augmented reading research (Kirner et al., 2012; Billinghurst et al., 2001; Hidayat et al., 2020; Grasset et al., 2008) aims to improve cognitive load (Cheng, 2017), knowledge accumulation (Dünser et al., 2012; Zhang et al., 2019), and spatial visualization (Shelton and Hedley, 2004), given a long-standing consensus that readers prefer physical paper over screen-based reading  (Tashman and Edwards, 2011a; Sellen and Harper, 2002).
-However, none of the existing systems have achieved an on-demand document enhancement.
-Instead, documents required pre-processing with content manually prepared before being embedded in AR.
-For example, HoloDoc (Li et al., 2019), PaperTrail (Rajaram and Nebeling, 2022), and Affinity Lens (Subramonyam et al., 2019) did not extract text directly from the given document; instead, content was prepared in advance and loaded based on attached fiducial markers.
+However, none of the existing systems have achieved on-demand document enhancement.
+Here, we define on-demand as the ability to dynamically and interactively enhance documents without requiring prior manual preparation or pre-processing.
+For example, prior work like HoloDoc (Li et al., 2019), PaperTrail (Rajaram and Nebeling, 2022), and Affinity Lens (Subramonyam et al., 2019) did not extract text directly from the given document.
+Instead, content was prepared in advance and loaded based on attached fiducial markers.
 As a result, these systems are not fully adaptable and deployable to real-world scenarios.
-Some preliminary works have explored on-demand text analysis using machine learning, such as Dually Noted (Qian et al., 2022) analyzing document structure in real-time, Augmented Math (Chulpongsatorn et al., 2023) extracting graphs and math equations, and SOCRAR (Strecker et al., 2022) using OCR to extract key information.
+Some preliminary work has explored on-demand text analysis using machine learning, such as Dually Noted (Qian et al., 2022) analyzing document structure in real-time, Augmented Math (Chulpongsatorn et al., 2023) and Augmented Physics (Gunturu et al., 2024) extracting diagrams and math equations, and SOCRAR (Strecker et al., 2022) using OCR to extract key information.
 However, none have yet achieved comprehensive and general-purpose document enhancement.
 Screen-Based Reading Support Tools
-Outside of the context of mixed reality, a wide range of projects and systems have explored reading support tools for screen-based interfaces.
-Examples including LiquidText (Tashman and Edwards, 2011b), texSketch (Subramonyam et al., 2020), ScholarPhi (Head et al., 2021), Chameleon (Masson et al., 2020), and Marvista (Chen et al., 2022) developed on-screen reading support tools.
-Similar to our work, Marvista (Chen et al., 2022) helps users read online documents by providing reading aids including summaries, contextual prompt questions, and reading metrics for active reading.
-Other works studied active and close reading support like Metatation (Mehta et al., 2017) for closed reading, XLibris (Price et al., 1998) with free form annotation, Matulic and Norrie (Matulic and Norrie, 2012) to support navigation, and GatherReader (Hinckley et al., 2012) with information gathering.
-Other works also support literature review and reference search for science papers (Palani et al., 2023; Kang et al., 2023; Chang et al., 2023).
-Lastly, works like Explorable Explanations (Victor, 2011) and Potluck (Litt et al., 2022) try to enhance static documents with interactive content to improve reading comprehension through exploration.
-While these past works are limited to screen-based interfaces, this paper explores the augmented reading design for AR-based interfaces.
+Outside of the context of mixed reality, a wide range of projects and systems have explored reading support via screen-based interfaces.
+Examples include LiquidText (Tashman and Edwards, 2011b), texSketch (Subramonyam et al., 2020), ScholarPhi (Head et al., 2021), Chameleon (Masson et al., 2020), and Marvista (Chen et al., 2022).
+Similar to our work, Marvista (Chen et al., 2022) helps users read online documents by providing reading aids including summaries, contextual prompt questions, and reading metrics.
+Other research has studied active and close reading support, including Metatation (Mehta et al., 2017) for close reading, XLibris (Price et al., 1998) for free form annotation, Matulic and Norrie’s tabletop interfaces (Matulic and Norrie, 2012) to support navigation, and GatherReader (Hinckley et al., 2012) for information gathering.
+Other systems have supported literature review and reference search for scientific papers (Palani et al., 2023; Kang et al., 2023; Chang et al., 2023).
+Lastly, works like Explorable Explanations (Victor, 2011) and Potluck (Litt et al., 2022) try to enhance static documents with interactive content to improve content comprehension through exploration.
+While this past work is limited to screen-based interfaces, our paper explores the augmented reading design for AR-based interfaces.
 1.2. Content Summarization
 Automatic Summarization
-In the field of natural language processing, several works have employed automatic summarization to support readers (El-Kassas et al., 2021; Zhang et al., 2023).
-These approaches provide various summarization methods, including abstractive and extractive summaries generated from document photos (Benharrak et al., 2022), real-time summarization of user writing for quick text iterations (Dang et al., 2022), enhancing reading comprehension through automatic summarization (Chen et al., 2022), and responsive text summarization for adapting document summaries to screen sizes ranging from large displays to small watch faces (Leiva, 2018).
-With the recent advances in large language models like GPT-4, such automatic summarization tasks have become more accurate, robust, and accessible.
+In the field of natural language processing, a variety of tools have employed automatic summarization to support readers (El-Kassas et al., 2021; Zhang et al., 2023).
+These approaches provide various summarization methods, including abstractive and extractive summaries generated from document photos (Benharrak et al., 2022), real-time summarization of user writing for quick text iterations (Dang et al., 2022), enhanced reading comprehension through automatic summarization (Chen et al., 2022), and responsive text summaries for adapting documents to screen sizes ranging from large displays to small watch faces (Leiva, 2018).
+With recent advances in large language models like GPT-4 and GPT-5, such automatic summarization tasks have become more accurate, robust, and accessible.
 For example, Goyal et al.
@@
 Researchers have also explored interactive summarization approaches based on context and users’ needs.
-For example, Hoeve et al.
-(ter Hoeve et al., 2020) studied Q&A interactions to assist users with a document-centric question-answering paradigm, finding that information-seeking was a common query.
-Semantic Reactor (Google, [n.
+For example, Semantic Reactor (Google, [n.
 d.]), developed by Google, enables users to ask questions about documents in natural language by fine-tuning BERT (Devlin et al., 2019) using the SQuAD dataset (Rajpurkar et al., 2018).
@@
 Hoeve et al.
-(ter Hoeve et al., 2020) identified various question types asked by participants regarding documents, with factual/summarization questions being the most frequent.
-Some question types include document-related, factoid, mechanical (answerable by rule-based methods), factual, yes/no, navigational, and summary questions.
-Voice assistants like Cortana, Alexa, Siri, VERSE (Vtyurina et al., 2019), Firefox Voice (Cambre et al., 2021), and Pushpak (Holani et al., 2019) offer efficient query modalities, as speech input has been found to be significantly faster than manual input, especially for mixed reality devices (Ruan et al., 2016; Adhikary and Vertanen, 2021).
-Our goal is to build on top of this prior work, developing and exploring automatic and interactive summarization techniques for on-demand and real-time summaries embedded in documents.
-The limitations of prior works such as Holodoc (Li et al., 2019), as explained before, prevented the study of long term and in-the-wild use of MR reading assistants.
-To our knowledge GazePointAR (Lee et al., 2024) is the closent work which explored long-term use of pronoun disambiguation in MR voice assistants.
-Still, it is unclear how can LLM powered MR assistants are used in everyday life to assist users in various types of reading and knowledge work tasks.
-We aim to fill this gap.
-In this paper, we contribute to the first demonstration of on-demand document enhancement by leveraging large language models.
-The goal of this paper is not only to demonstrate its potential but also to uncover its limitations and challenges from a user experience perspective through iterative implementation and multiple user evaluations.
+(ter Hoeve et al., 2020), meanwhile, identified a diverse questions asked by participants while reading documents, with factual/summarization questions being the most frequent.
+Other types include document-related, factoid, mechanical (answerable by rule-based methods), factual, yes/no, navigational, and summary questions.
+Voice assistants like Cortana, Alexa, Siri, VERSE (Vtyurina et al., 2019), Firefox Voice (Cambre et al., 2021), and Pushpak (Holani et al., 2019) offer efficient query modalities, as speech input has been found to be faster than manual input, especially for mixed reality devices (Ruan et al., 2016; Adhikary and Vertanen, 2021).
+Reality-Based Information Retrieval
+Other recent research has examined visual reality-based information queries (Büschel et al., 2018).
+For example, GazePointAR (Lee et al., 2024) explored long term use of pronoun disambiguation in visual question answering in mixed-reality.
+G-Voila (Wang et al., 2024) explored gaze as a medium for question answering in everyday scenarios.
+While prior work has either focused on document augmentation with manual preparation (Li et al., 2019) or on general mixed-reality LLM-based QnA (Lee et al., 2024), our goal is to build on this foundation by developing and exploring on-demand and real-time reading augmentation within everyday Mixed-Reality settings.
+To our knowledge, RealitySummary is the first LLM powered MR reading system to be studied in the wild over an extended period, enabling the collection of rich, longitudinal insights into real-world reading behaviors.
+This long-term, situated deployment serves as the foundation for our contribution, upon which we highlight implications of on-demand, context-aware mixed reality augmentations.
+Figure 2. Overview of our research approach spanning an evolution of our system across three different versions.
 2. Overview of Our Research Approach
-Figure 2. Overview of our research approach spanning an evolution of our system across three different versions.
-To uncover insights into on-demand reading assistants, we employed a three-phase iterative design process to investigate user interactions and experiences with RealitySummary.
-Our approach consisted of a usability study on the initial prototype (Study 1: RS-Doc), followed by an in-the-wild study on a second prototype (Study 2: RS-Wild), and concluded with an author-conducted diary study using a third prototype (Study 3: RS-Diary).
+To uncover insights into on-demand reading assistants, we employed a three-phase iterative design process to investigate reader interactions and experiences with MR reading assistants.
+These explorations included a usability study using the initial prototype (Study 1: RS-Doc), followed by an in-the-wild study with a second prototype (Study 2: RS-Wild), and an author-conducted diary study with a third prototype (Study 3: RS-Diary).
 Each phase was designed to build upon the insights gained from the previous round, allowing for continuous refinement of the design and functionality.
-We chose this multi-method, iterative approach to capture a comprehensive understanding of user interactions across various contexts.
-The initial usability study provided controlled insights into the basic functionality and user experience of RealitySummary.
-The subsequent in-the-wild study allowed us to observe how users interacted with the improved, more reliable prototype in everyday environments, uncovering real-world usage patterns and challenges.
-Finally, the diary study offered deep, longitudinal insights into the integration of RealitySummary into daily life, allowing for nuanced observations of subtle interaction patterns and evolving user needs.
-This progression from controlled to naturalistic settings, coupled with iterative prototyping, allowed us to holistically explore the complex dynamics of AR-AI interactions while addressing emerging design considerations and ethical implications at each stage.
-3. Study 0: Elicitation Study
-Before implementing our system, we conducted a design elicitation study to explore potential designs for an AI-assisted augmented reading assistant.
-We held formative design workshops with six participants (4 males and 2 females) who had various design backgrounds as undergraduate and graduate HCI researchers.
-The study involved two tasks, each lasting approximately 15-20 minutes, with the entire session taking about 60 minutes.
-3.1. Method
-We asked participants to complete a four-part design task where they read and graphically annotated both a document of their own and a Wikipedia article [^1] Source: https://en.wikipedia.org/wiki/Russian_invasion_of_Ukraine.
-After reading, participants brainstormed ideas for AI-powered augmentations to improve readability.
-In the first design task, participants annotated a physical printout of their own article, sketching ideas for possible enhancements, and then repeated the task with a printed copy of the Wikipedia article.
-In the second task, participants were provided with digital copies of the two documents on a virtual whiteboard (Miro) and asked to design a second set of document enhancements, this time with assistance from ChatGPT.
-Before the study began, we briefed participants, gave them a reference page of sample visualizations, and demonstrated prompts for ChatGPT that showcased its summarization and enhancement capabilities.
-Using a think-aloud protocol, the authors took detailed notes and collected all participant sketches.
-These designs were later collated on a Miro board, where four authors independently categorized and coded them, identifying 20 distinct types of document enhancements based on usage, purpose, popularity, and context.
-The authors then discussed and refined these categories, re-coding the examples to produce a final set of enhancement types  (Figure 3).
-Figure 3. In our elicitation study, participants identified a variety of potential summary and augmentation types based on sketching and ChatGPT exercises using diverse text documents.
+Insights from each study motivated the design of the subsequent system, and the three studies are presented together in this paper as part of a single continuous investigation, providing a complete picture of the iterative design process and cross-study findings.
+All studies were conducted under the approval of the University of Calgary ethics board (IRB Approval Number: REB21-2065).
+Participants provided informed consent prior to participation and could withdraw at any time without penalty.
+3. RS-Doc: Design Exploration and Usability Study with the First Prototype
+We implemented the first version of our system (RS-Doc), which focused on in-place summarization for physical and digital documents, on a Microsoft HoloLens 2.
+In this section, we present initial observations from our experiences developing and evaluating RS-Doc as well as new design opportunities highlighted by those explorations.
+3.1. Implementation
+System Design
+RS-Doc captures documents using the HoloLens 2’s internal camera and extracts text with Google Cloud OCR.
+The extracted text is processed using GPT-4 via OpenAI’s API to identify key information, while spaCy’s named entity recognition (NER) module assists in identifying important entities.
+The system employs Vuforia for markerless image tracking, and the Mixed Reality Toolkit 3 (MRTK3) SDK for content rendering.
+Readers interact with the system through voice commands processed via Google’s speech-to-text API, allowing them to modify or query visualizations.
+The system captures images periodically for OCR, with an option for manual activation, and tracks documents by generating dynamic image targets, ensuring accurate document tracking and augmented content rendering.
+Enhancement Types
+To design and implement our system, we identified five key categories of AI-generated content enhancements: summarize, compare, augment, extract, and navigate—based on an early elicitation study (see appendix) as well as prior work, such as HoloDoc (Li et al., 2019), PaperTrail (Rajaram and Nebeling, 2022), DuallyNoted (Qian et al., 2022), QOOK (Zhao et al., 2014), and Digital Desk (Li et al., 2019).
+These categories informed both the conceptual framework and the system features.
+Summarize condenses text for quick reference, ranging from broad overviews to personalized summaries.
+This is implemented by sending the current page’s text to GPT-4 for concise summaries, limited to 150 words.
+Compare transforms data into visual formats by parsing documents.
+Our system provides the compare feature by generating comparison tables for documents discussing multiple related themes, entities, or concepts, as well as timelines for documents describing sequences of events.
+Augment enhances content with external data.
+For instance, HoloDoc (Li et al., 2019) and PaperTrail (Rajaram and Nebeling, 2022) augment papers with external content, such as references, videos, and figures.
+Similarly, our system augments documents by providing information cards about people and places using Google Search, Image APIs, and Mapbox, based on keywords identified by spaCy.
+Extract enables readers to pull and store elements like keywords and quotes.
+For example, DigitalDesk (Wellner, 1991) allows users to extract content from documents for copy-and-paste functionality.
+RS-Doc implements this feature by generating keyword lists that highlight the most central named entities extracted.
+Finally, Navigate improves document navigation through features like a table of contents or bookmarks.
+For instance, QOOK (Zhao et al., 2014) allows users to add bookmarks to specific sections.
+RS-Doc supports navigation by generating a table of contents, created by the LLM based on the document.
+Interactions
+After launching RS-Doc, a reader can trigger summaries for an in-view document by saying “give me summaries”.
+The system then captures an image of the page currently in view, creates a dynamic image target, and generates summary cards of all types relevant to that page.
+Summaries are displayed around the page using a fixed layout (for example, timeline summaries always appear at the top right of the page and information cards always appear to the left) and remain locked to the page if either the document or reader moves.
+For example, if a user is reading a magazine about aerial drones, the system automatically detects important keywords such as people, places, and phrases from the page.
+The system then generates an image, a comparison table, a summary, and information about key people and keywords.
+Readers can also show or hide specific cards using voice commands (“hide keywords”, “show timeline”, etc.).
+Note that the system only supports one card of each type at a time.
+The user can query for a card multiple times, which will re-prompt the LLM (each card type has its own fixed prompt which the user cannot change).
+The decision to lock content to the document was motivated by the limitations noted in prior work like Holodoc (Li et al., 2019), where the participants expressed frustration and preferred the content to follow their views.
+Additionally, prior research, such as Dually Noted (Qian et al., 2022), indicates that the participants preferred not to obstruct the view of the page itself.
+Our system displays the cards around the page and tracks them in MR with its physical movement.
+Document Tracking and Content Generation Performance
+To characterize the technical performance of our prototype, we assessed document tracking reliability and system latency.
+To test tracking, we collected 60 seconds of tracking data each for 20 diverse documents (10 physical, 10 on-screen) including posters, newspapers, patents, and books—then recording what fraction of time the document was correctly registered.
+Tracking was largely reliable for documents with visuals (M=92%, SD=5.4%) but less so for text-only documents (M=64%, SD=3.7%) which had fewer distinct features.
+Document tracking was more reliable when visuals (e.g., diagrams or figures) were present, but less reliable for text-only documents.
+Tracking performed poorly as compared to text detection because Vuforia’s image tracking depends on detecting distinctive feature points, which are more abundant in images and figures than in uniform text regions.
+Text-only pages often contain repetitive patterns and large blank areas harder to capture with the HoloLens camera that provide fewer robust keypoints, reducing tracking stability.
+Performance varied with lighting (better in well-lit conditions), target size (larger images performed better), and motion (sudden movements caused temporary tracking loss).
+We also measured the average latency of three components: AR rendering, network communication, and API responses.
+AR rendering averaged 100 ms on HoloLens, while network communication via socket.io averaged 40 ms, indicating responsive performance.
+API callbacks included GPT-4 requests (2.4 s) and our NLP module using spaCy (670 ms).
+Since we limited responses to a short response (150 words for a summary), GPT-4 latency remains relatively low.
+These latencies remained largely consistent across our two later prototypes.
+3.2. Study Method
+Using the RS-Doc prototype, we conducted an initial usability study involving 12 participants (P1-P12, 6 male and 6 female, aged 19-35) recruited via snowball sampling and university-wide advertisements.
+We compensated participants \$15 CAD for the one-hour session.
+All participants reported to be proficient in English and accustomed to reading academic or professional materials.
+As all were enrolled university students, they met our inclusion criterion of being able to read documents of varied comprehension levels.
+No additional exclusion criteria were applied.
+The study aimed to evaluate the system’s usability, explore participants’ preferences for different document enhancement strategies, and identify limitations for future iterations, with real-world use in mind.
+We conducted the study in a lab environment, where we first introduced participants to the HoloLens 2 and then provided an overview of the system and the set of document enhancements.
+Participants used our system to explore a variety of materials and media: 1) an A4-sized magazine and an A3-sized book in physical format, and 2) a digital document of their choice.
+The physical magazine and book contained the same reading content for all participants, while the digital documents were selected by participants to reflect their diverse preferences and included blog posts, course assignments, news articles, Wikipedia entries, and digital books.
+The physical magazine and book each contained fixed content, and the same materials were used across all participants.
+This ensured that differences in responses could be compared consistently within each medium, without variability introduced by differing texts.
+This study aimed to investigate participants’ interactions and preferences across different reading materials and formats, rather than focusing on standardized tasks; therefore, we did not explicitly measure or control for difficulty levels.
+We tasked the participants with reading and understanding the documents provided to them, using RealitySummary in whichever way they saw fit and using whatever combination of enhancements best suited the document.
+Throughout the study, we encouraged participants to think aloud and documented their experiences.
+Following the tasks, we also conducted semi-structured interviews and a Likert scale questionnaire, including a System Usability Scale (SUS) (Brooke et al., 1996).
+We then performed a reflexive thematic analysis (Clarke and Braun, 2014) using transcripts of both the think-aloud sessions and interviews, from which we identified and refined key themes.
 3.3. Results
-We identified five high-level categories: summarize, compare, augment, extract, and navigate.
-The summarize category focused on condensing text for quick reference, ranging from broad document summaries to personalized summaries for different comprehension levels.
-Summaries were seen as useful reference tools, allowing users to quickly review content without losing focus.
-The compare category emphasized transforming textual and numerical data into visual formats like tables, graphs, and timelines, which could dynamically update based on user interactions, making comparisons more intuitive.
-The augment category explored ways to enhance document content with external or AI-generated data, such as adding maps and images or converting 2D figures into 3D models.
-Advanced ideas included visualizing the importance of academic citations or integrating external information to provide richer context.
-The extract category focused on enabling users to pull out and store key elements, such as keywords, figures, or quotes, for future reference.
-Lastly, the navigate category proposed various methods to improve document navigation, such as progress indicators, accessible tables of contents, and using AR to make documents searchable.
-These insights, alongside considerations of the potential and challenges of AI-driven tools, informed the development of our system.
-4. Study 1 RS-Doc: RealitySummary First Prototype and Usability Study
-We implemented the first version of our system on a Microsoft Hololens 2 and conducted a preliminary user evaluation.
-In this section, we present initial insights from using on-demand, mixed reality document summarization and discuss the associated limitations and design implications.
-4.1. Implementation
-Figure 4. RS-Doc Hololens 2 architecture.
-Figure 5. Examples from our RS-Doc Hololens implementation.
-The system design for the HoloLens version involves capturing documents using the HoloLens 2’s internal camera and extracting text with Google Cloud OCR.
-The extracted text is processed using GPT-4 via OpenAI’s API to identify key information, while spaCy’s named entity recognition (NER) module assists in identifying important entities for further web searches.
-The system employs Vuforia for markerless image tracking and the Mixed Reality Toolkit 3 (MRTK3) SDK for content rendering.
-Users interact with the system through voice commands processed via Google’s speech-to-text API, allowing them to modify or query visualizations.
-The system captures images periodically for OCR, with an option for manual activation, and tracks documents by generating dynamic image targets, ensuring accurate document tracking and augmented content rendering.
-Informed by the elicitation study, the system supports several document enhancements aimed at improving the reading experience.
-These enhancements include text summarization, achieved by sending the current page’s text to GPT-4 for a concise summary, limited to 150 words for easier digestion.
-Comparison tables are generated on demand based on the document parsing, allowing for quick comparisons within the document.
-Timelines of events are created by prompting GPT-4 to generate a sequence of summarized events.
-A keyword list is extracted using GPT-4 and spaCy, with spaCy handling named entities for web searches and keyword highlighting.
-Summary highlighting allows users to obtain brief summaries or additional information on specific phrases, sentences, or paragraphs, which are then visually highlighted.
-Information cards provide further details on keywords, people, places, or images, leveraging Google Search, Image APIs, and Mapbox for generating maps.
-All these features are integrated into the HoloLens environment, offering real-time, interactive content within the mixed reality experience.
-4.2. Method
-The usability study involved 12 participants (P1-P12, 6 male and 6 female, aged 19-35) recruited via word-of-mouth and compensated \$15 for their one-hour participation.
-The study aimed to evaluate the system’s usability, explore participants’ preferences for different document enhancement strategies, and identify limitations for future iterations with real-world use in mind.
-Conducted in a lab environment with Microsoft HoloLens 2, participants were first introduced to AR and the HoloLens 2, followed by an overview of the system and its document enhancements.
-They completed three reading tasks using different document formats (physical, digital, and a book) and utilized the system’s enhancements as they preferred.
-The participants’ digital documents included books, blog posts, course assignments, Wikipedia articles, magazine articles, and news articles viewed on various devices.
-Throughout the study, participants were encouraged to think aloud, and their experiences were documented.
-Following the tasks, semi-structured interviews and a Likert scale questionnaire, including a System Usability Scale (SUS), were conducted.
-Thematic analysis was performed on the transcriptions of the think-aloud sessions and interviews, resulting in the identification and refinement of key themes.
-4.3. Results
-Overall, participants responded positively to RealitySummary, with most participants agreeing that the prototype was easy to use (M=4.3/5, SD=0.98), easy to understand (M=4.5/5, SD=0.67), and well-integrated with the task (M=4.3/5, SD=0.77) with 5-point Likert-scale questionnaire.
-Similarly, nearly all participants reported that the summary approaches provided by RealitySummary helped with comprehension (M=4.6/5, SD=0.514), were relevant to the document content (M=4.7/5, SD=0.65), improved comprehension speed (M=4.4/5, SD=0.79), and augmented their reading experience (M=4.7/5, SD=0.65).
+Overall, participants responded positively to RealitySummary, with most participants agreeing that the prototype was easy to use (M=4.3/5, SD=0.98), easy to understand (M=4.5/5, SD=0.67), and well-integrated with the task (M=4.3/5, SD=0.77) on a 5-point Likert-scale questionnaire.
+Similarly, nearly all participants reported that the summary approaches provided by RealitySummary helped with perceived comprehension (M=4.6/5, SD=0.514), were relevant to the document content (M=4.7/5, SD=0.65), perceived improvements in comprehension speed of the document (M=4.4/5, SD=0.79), and augmented their reading experience (M=4.7/5, SD=0.65).
 Finally, the System Usability Scale (SUS) returned an overall composite score of 71, suggesting a reasonable level of usability for an early-stage prototype.
-4.4. Insights
-Our analysis identified several themes related to participants’ experience in comprehending and navigating documents.
-We also explore participants’ comparisons of the visual and text-based summarization techniques as well as preferences for various augmentation techniques.
-Enable Multi-Angle Content Understanding for the Current Page
-RealitySummary’s document enhancements can significantly improve readers’ comprehension.
-Several participants (P7, P8, P11, P12) highlighted that features such as summaries, timelines, keywords, images, and people cards provided a more comprehensive understanding of the document, allowing them to engage with the content from multiple angles and perspectives.
-All participants mentioned that RealitySummary enabled them to grasp content more quickly than with their usual methods, such as visually scanning a physical book or manually searching for timelines on a phone or PC.
-For instance, after reviewing the first magazine, P12 remarked, I haven’t read this magazine thoroughly, but it seems these features can significantly help me understand what it’s about in less time.” Interestingly, several participants appreciated that the summaries and keywords were drawn only from the currently visible page rather than from the entire document.
-P5 reported, I like the page-by-page summary—having a summary for each page does help make it feel less dense compared to a summary of the whole document.” Similarly, P6 emphasized, “I like that it (the system) gives you keywords for each page and not for the whole document.
-Facilitate Efficient, Top-Down Navigation
-Both structured overviews and in-context highlights provide useful entry points that make it easier to navigate documents from a top-down perspective, even within individual pages.
-Broadly, participants spoke positively about how RealitySummary helped them navigate documents.
-Four participants (P3, P4, P6, P8) specifically mentioned using structured navigation elements like timelines to better understand the structure and flow of the text, with P4 noting that the timeline helped me organize and navigate my thoughts very quickly while reading the text.” In fact, ten participants reported that the system allowed them to focus their reading by taking a “top-down” approach—first examining the high-level ideas and terms, then transitioning into reading the lower-level content as needed.
+3.4. RS-Doc Observations
+Our analysis identified several themes related to participants’ experience comprehending and navigating documents.
+We also explore participants’ comparisons of the visual and text-based summarization techniques as well as their preferences for various augmentation techniques.
+Page-Level Summaries and Keyword Lists Support Quick Comprehension.
+Since our RS-Doc system can only capture the currently visible page, our summaries are inherently focused on page-level content rather than the entire book.
+Interestingly, when asked about their experiences, several participants expressed appreciation for this feature.
+P5 reported, “I like the page-by-page summary—having a summary for each page does help make it feel less dense compared to a summary of the whole document.” Similarly, P6 emphasized, “I like that it (the system) gives you keywords for each page and not for the whole document.
+Structured Overviews can Facilitate Efficient Document Navigation.
+Both structured overviews and in-context highlights provide useful entry points that make it easy to navigate documents from a top-down perspective, even within individual pages.
+Four participants (P3, P4, P6, P8) specifically mentioned using structured navigation elements like timelines to better understand the structure and flow of the text, with P4 noting that “the timeline helped me organize and navigate my thoughts very quickly while reading the text.” In fact, ten participants reported that the system allowed them to focus their reading by taking a “top-down” approach—first examining the high-level ideas and terms, then transitioning into reading the lower-level content as needed.
 P8 highlighted, “[the enhancements] give you a foundation before you actually jump inside [the document].” This approach enabled participants to prioritize their reading, allowing them to “focus on things that you’re more unsure about” (P1).
-Support Quick Understanding with Visual Aids
+Visual aids improve engagement and memorability.
 Both visual and text-based enhancements offer distinct benefits to readers—visual aids provide quick, high-level overviews, while text summaries capture more complex details.
@@
 All 12 participants rated the visual summarization techniques as helpful, with eight specifically highlighting the usefulness of images and bio cards.
-Several participants (P1, P4, P6, P7, P9, P10, P11) noted that visual augmentations allowed them to grasp the document’s content more quickly compared to reading text alone, with P11 stating, “Seeing the visuals helped me grasp what’s happening in the document right away.” Participants (P7, P8) also emphasized that visuals enhanced their engagement with the material, with P7 commenting, I’m not a big fan of reading, so having visuals helps me engage with the content.” Additionally, participants highlighted the potential of visuals to improve memorability, with P8 noting that for historical documents, people cards helped put a face to a name.” Similarly, P7 observed that while text-based summaries aided understanding, “graphical elements help me remember it for a longer time.”
-4.5. Limitations
-Slow and Unreliable Document Tracking
-Document tracking enables unique features like keyword highlighting and around-document placement.
-However, the primary limitation of the current prototype is the instability of its document tracking.
-Overall, participants reported dissatisfaction with the system’s slow, inaccurate, and unreliable tracking capabilities.
-Unlike the OCR process, document tracking involves a heavier and slower workflow, requiring rectangle detection, image extraction, and image upload for Vuforia processing.
-Even with a perfect image target, tracking performance varies based on several factors, such as lighting conditions, hand occlusion, and viewing angles.
-Furthermore, while the tracking performed better for visually rich documents, it struggled with text-heavy documents, as it does not have any distance features, making consistent tracking difficult over time.
-Even when tracking functioned properly, participants frequently experienced disruptive flickering, which negatively affected the user experience.
-Additionally, the fixed placement of augmented information on documents posed a challenge for long-term use, as participants expressed the need for more flexibility in adjusting the location of the augmentations.
+Several participants (P1, P4, P6, P7, P9, P10, P11) noted that visual augmentations allowed them to grasp the document’s content more quickly compared to reading text alone, with P11 stating, “Seeing the visuals helped me grasp what’s happening in the document right away.” Participants (P7, P8) also emphasized that visuals enhanced their engagement with the material, with P7 commenting, I’m not a big fan of reading, so having visuals helps me engage with the content.” Additionally, participants highlighted the potential of visuals to improve memorability, with P8 noting that for historical documents, people cards helped “put a face to a name.” Similarly, P7 observed that while text-based summaries aided understanding, “graphical elements help me remember it for a longer time.”
+3.5. Opportunities for the Second Iteration
+Building on these observations, we identified a number of opportunities to explore in our second iteration.
+Reader-Controllable Placement Rather Than Document Tracking
+An intriguing finding that might contradict previous work is the role of document tracking.
+For on-demand document enhancement, where dedicated markers or image targets cannot be prepared in advance, unreliable or slightly jittery document tracking often leads to frustration.
+For example, while HoloDoc (Li et al., 2019) and Dually Noted (Qian et al., 2022) relied on dedicated markers or pre-defined image targets, enabling much more reliable document tracking, our workflow cannot rely on these methods.
+Instead, we need to rely on complex and unreliable image extraction workflows.
+Moreover, document tracking performed well only with visually rich documents.
+Participants expressed frustration when dealing with text-heavy documents, particularly under challenging conditions such as poor lighting, hand occlusion, or extreme viewing angles.
+Participants also expressed the need for flexibility in the placement.
+Therefore, while we acknowledge the benefits of document tracking, it is worth exploring reader-controlled placement as a less frustrating alternative.
 Layouts and Visual Clutter
@@
 P9 even stated, “I didn’t even read the document, but I understand what it’s about.” Currently, the system automatically generates various types of cards when relevant, but allowing users to reduce the number of visible cards based on their preferences could help manage visual clutter and maintain focus.
-Fixed Prompt and Augmentation Types
-Another key issue was the inability of participants to query the generated summaries.
-The fixed prompt approach used for generating the different types of cards limited the system’s flexibility in addressing broader use cases and contexts.
-Several participants (P1, P3, P5) expressed frustration with this constraint, as they had specific questions they wanted to ask but were restricted to the preset summary types provided.
-While these different types of augmentations offered different benefits, those advantages were only realized when the generated content was relevant and accurate.
-Participants also noted that specific aids, like timelines, comparison tables, and keyword lists, were not always helpful or contextually appropriate.
-On the other hand, participants generally appreciated the broader usefulness of the general summaries and visual information cards, which were more adaptable to varying contexts.
-Applicability Beyond Controlled Environments
-Finally, the controlled nature of the study limits its ability to assess the broader applicability and real-world uses.
-While the initial findings are encouraging, they do not fully reflect how the system might perform in diverse, unstructured environments.
-As a result, transitioning to in-the-wild studies will be essential to understanding how document augmentations work in everyday scenarios and to gain insights from prolonged, regular use in real-world settings.
-In conclusion, a more refined version of the system, featuring flexible augmentation placement and query-able summaries, is essential for better evaluating its effectiveness in real-world scenarios.
-5. Study 2 RS-WILD: In-the-Wild Study with the Second Prototype
-Figure 6. Examples from our in-the-wild study.
-To explore real-world use cases, we conducted an in-the-wild study with eleven participants.
-Unlike the previous usability study, which focused on standardized reading tasks in a controlled lab setting, this study aimed to examine the versatility and applicability of our system in uncontrolled environments and with diverse reading materials.
-The goal was to uncover unique advantages and potential use cases for using mixed reality and AI-generated assistants in everyday settings.
-For this study, the system was deployed on an Apple Vision Pro, offering considerably higher rendering quality and a broader field of view compared to the Hololens 2.
-This contributes to better readability, particularly for text summarization, by using a less distractive transparent background.
-Based on the user feedback from the usability study, the user interface was simplified with a single text box and an image slot, which could accommodate any response type given the user’s custom query.
-This feature would better suit diverse applications and minimize the intrusiveness of additional content in daily use.
+Querying and Contextual Flexibility
+Participants highlighted the potential for enhancing the system by enabling queries on the generated summaries, which could expand its adaptability to broader use cases and contexts.
+The current fixed prompt approach for generating different types of cards suffices for general reading but could be further refined to address more specific reader questions.
+Several participants (P1, P3, P5) envisioned a more conversational system where they could ask targeted questions beyond the preset summary types.
+While participants found general summaries and visual information cards broadly useful and adaptable to varying contexts, they also identified opportunities to enhance aids with specific queries for timelines, comparison tables, and keyword lists that better align with their content.
+Overall, the controlled nature of the study limited its ability to assess broader applicability and real-world uses.
+While the initial findings are encouraging, they do not fully reflect how AR summarization and question-answering tools might perform in diverse, unstructured environments.
+As a result, transitioning to in-the-wild studies is essential to understanding how these augmentations might work in everyday scenarios.
+4. RS-Wild: In-the-Wild Study with the Second Prototype
+To explore these kinds of real-world use cases, we developed a second prototype (RS-Wild) that allowed us to study summarization and question-answering in-the-wild.
+To address performance issues with the HoloLens 2 and support more free-ranging use, we instead implemented our RS-Wild prototype for the Apple Vision Pro.
+This provided higher rendering quality for cards, as well as a broader field of view.
+Building on observations from our prior study, we also simplified the interface—replacing the suite of document-anchored cards from RS-Doc with a single reader-positionable card that readers could dynamically update using spoken queries.
+Figure 3. Snapshots captured from the Vision Pro during our in-the-wild study.
 4.1. Implementation
-The system design for the Vision Pro follows a similar approach to the Hololens but with key adjustments.
-Since the Vision Pro does not yet allow direct access to the camera view, we implemented a workaround for text extraction.
+Our RS-Wild implementation for the Vision Pro follows a similar approach to the HoloLens but with a few key adjustments.
+Since the Vision Pro does not yet allow direct access to the camera feed, we implemented a workaround for text extraction.
 First, we screencast the Vision Pro to an external iPad, which is then connected to a MacBook Air to stream the feed via QuickTime Player’s mirroring feature.
-A web-based interface on the Mac captures the live stream from the Vision Pro.
+A web-based interface on the Mac captures the live stream from this mirror and by proxy, from the Vision Pro.
 Using the captured feed, we employed Google OCR to extract text, which was then processed by GPT-4 via the OpenAI API.
-The Vision Pro application, developed in Swift, featured a single movable card with an image/map slot.
-After querying the LLM, data was transmitted from the web-based interfaces to Vision Pro via Firebase’s real-time database.
-Additionally, a separate web interface with speech recognition was run on the user’s smartphone, allowing them to query the system via voice (using Web Speech Recognition) and ask follow-up questions.
-This enabled users to request specific summary types or submit open-ended queries about the text they were reading.
-The web-based interfaces, built with React.js and hosted on Glitch, were accessed by participants through their phones.
-5.2. Method
-We recruited 11 participants (W1-W11, 7 male and 4 female; ages 19-32 years) from the local community.
-They were compensated \$25 for participating in a 1.5-hour study, which included a one-hour experiment and a 30-minute interview.
-The study took place in various locations at a local university, such as a library, cafeteria, store, hallway, classroom, and multiple labs.
-Before the experiment, we collected participants’ demographic information and reading habits, which were input into the system as prior knowledge for ChatGPT.
+After querying the LLM, data was transmitted from the web-based interfaces to the Vision Pro via Firebase’s real-time database.
+Voice commands were handled via a separate interface (using the Web Speech API and React.js and hosted as a progressive web app) loaded on the reader’s smartphone.
+Interactions
+RS-Wild allows readers to request summaries and pose free-form questions about any text visible in their environment by clicking on a mic button on the smartphone interface and asking a question or clicking a “summaries” button.
+This then triggers the OCR and LLM pipeline.
+Responses to the most recent request are displayed on a single reader-positioned card which can be placed anywhere in the environment.
+This added flexibility allowed viewers to recreate any of the individual cards available in RS-Doc as well as pose new unanticipated questions and request new summary types.
+Using a single manually-positioned card with a transparent background also allowed readers to position that response or summary in more useful locations in their environment—placing it both close to and apart from documents as appropriate for their current task.
+Moreover, the user can choose to keep a history of all the text that has been captured in their session, allowing them to pose questions with a little context.
+Additionally, the card follows the user’s position in MR space, so they always have access to it even while walking or changing locations.
+4.2. Study Method
+To explore question-answering and summarization in-the-wild we recruited a new cohort of 11 participants (W1-W11, 7 male and 4 female; ages 19-32 years) from our university through snowball sampling and university advertisements.
+All participants were university students accustomed to reading in academic environments, which met our inclusion criterion of being able to read documents of varied comprehension levels.
+No additional exclusion criteria were applied.
+We compensated participants \$25 CAD for the 1.5-hour study, which included a one-hour experiment and a 30-minute interview.
+The participants were first called to our lab, briefed on the procedure, and given the opportunity to pick one or multiple places around the university campus (such as a library, cafeteria, store, hallway, classroom, or various labs) to explore with the system.
+They were allowed to freely explore any place in the campus with the headset and the study conductor would follow them.
+Before the experiment, we collected participants’ demographic information and reading habits, which we then input into the system as prior knowledge for GPT-4.
 Participants familiarized themselves with the Apple Vision Pro by completing the eye-tracking calibration setup.
-Unlike structured tasks or prepared reading materials, participants were encouraged to use the system with any text they encountered during the study, including documents, books, papers, posters, online news on mobile devices, and other materials in their environment.
-Throughout the one-hour experiment, participants were encouraged to think aloud about their experiences.
-After the experiment, we conducted semi-structured interviews, followed by a Likert scale questionnaire similar to the one used in the earlier usability study.
-Both the interview and questionnaire took approximately 30 minutes.
-As in the previous usability study, we analyzed participants’ questionnaire responses and reflections to identify high-level trends and opportunities for augmented reading tools.
+Instead of structured tasks or prepared reading materials, we encouraged participants to use the system with any text they encountered during the study, including documents, books, papers, posters, online news on mobile devices, and other materials in their environment.
+We also directed participants to explore various locations at a local university, rather than using spaces prearranged by the experimenters.
+Throughout the one-hour experiment, an experimenter accompanied the participants, observing their usage and encouraging them to think aloud about their experiences.
+After the experiment, we conducted semi-structured interviews, followed by a Likert scale questionnaire similar to the one used in Study 1.
+As in the previous usability study, we analyzed participants’ questionnaire responses and transcripts to identify high-level trends and opportunities for augmented reading tools.
+We performed a reflexive thematic analysis (Clarke and Braun, 2014) using transcripts of both the think-aloud sessions and interviews, from which we identified and refined key themes.
 4.3. Results
@@
 In general, participants reported feeling more immersed and interested while reading (M=3.5/5, SD=1.04), and they perceived the system as effectively interpreting their queries (M=3.3/5, SD=1.00).
-Besides, their responses on the content placement within the system were also overall positive (M=3.4/5, SD=1.29).
-Figure 7. In-the-wild study questionnaire responses.
+Their responses on the content placement within the system were also positive overall (M=3.4/5, SD=1.29).
 Participants had varied opinions on whether the system became distracting throughout the reading process (M=2.6/5, SD=1.37) and whether they tended to trust the accuracy and reliability of contents provided by the system (M=2.8/5, SD=1.25).
-Privacy concerns were also noted by the majority (6/11), and discomfort using the headset was prevalent among the respondents (8/11).
-These findings are detailed in Figure 7.
-Figure 8. Snapshots captured from the Vision Pro during our in-the-wild study.
-5.4. Insights
+Privacy concerns were also noted by the majority (6/11), and discomfort using the Vision Pro headset was prevalent among the respondents (8/11).
+4.4. RS-Wild Observations
 Observations from the in-the-wild deployment highlighted a wider range of use cases for augmented reading as well as some of the practical trade-offs associated with real-world use.
-Types of Reading Content
-The in-the-wild study demonstrated our system’s broad utility, revealing diverse exploratory use cases across various reading materials.
-Participants engaged with a variety of books, spanning textbooks, encyclopedias, fiction, and non-fiction.
-The system was also employed for different types of documents, including academic papers, handouts, and class assignments, as well as digital content accessed through smartphones and laptops, like social media posts, online news articles, and e-commerce websites.
-Beyond conventional reading materials, participants also used our system with everyday objects and texts in the environment, such as restaurant menus, printer instructions, product nutrition labels, furniture assembly guides, posters, and signage on trash bins and advertisements.
+Diverse Reading Content
+Participants engaged with a variety of books (spanning textbooks, encyclopedias, fiction, and non-fiction) as well as other types of documents (including academic papers, handouts, and class assignments) and diverse digital content accessed through smartphones and laptops (like social media posts, online news articles, and e-commerce websites).
+Beyond conventional reading materials, participants also used our system with everyday objects and text in the environment, such as restaurant menus, printer instructions, product nutrition labels, furniture assembly guides, posters, signage on trash bins, and advertisements.
 Notably, some participants even applied our system to their handwritten notes from lectures.
-Benefits of Always-On AI Assistant
-The exploration of such diverse reading materials can be attributed to the always-on capability of the mixed reality reading assistant, as opposed to screen-based interfaces.
-A key benefit of this feature is its support for implicit inputs, which dramatically lowers the barrier to asking questions of LLMs.
+Incorporating Spatial and Tangible Interactions
+Participants appreciated the unique affordances offered by mixed reality interfaces, such as the opportunities for spatial and tangible exploration.
+In a library setting, for example, participants could naturally navigate through bookshelves (spatial exploration), physically pick up a book (tangible interaction), and browse through its pages to automatically receive summaries.
+W5 and W10 even leveraged our system to provide summaries for entire collections of books on a bookshelf, not just content from a single page.
+Furthermore, W1 demonstrated the ability to synthesize information across multiple books on a table, requesting summaries or connections between various books and pages.
+Such observations highlight how mixed reality interfaces enable spatial and tangible interactions that go beyond the limitations of traditional screen-based interfaces, which can only capture and summarize visible content.
+Creative Uses of AI-Generated Responses
+We observed several unique and unexpected use cases that took advantage of the LLM’s capacity for free-form questioning and answering.
+For instance, when reading a physics textbook, W8 asked the system to generate ten quizzes to help him understand quantum mechanics, transforming the textbook into a more active and personalized learning tool.
+Other common uses included translation (W4, W5, W6), creating learning aids (W4, W8), and generating personalized content recommendations (W5, W10).
+For example, W8 applied the system to his handwritten calculus notes to find answers to particular questions.
+Beyond traditional reading applications, the AI’s capabilities supported practical decision-making.
+For example, W4 compared nutritional information of products on store shelves and menu items to determine healthier options, while W5 used it for guidance on waste disposal and printer operations.
+The LLM’s ability to process a broad spectrum of inquiries clearly expanded the system’s applicability and utility.
+Proactive Summaries vs Reader-Driven Question Answering
+Our system offers both 1) proactive assistance through automatic summary generation and 2) reader-driven assistance where readers manually ask questions.
+Participants shared their insights on these two functionalities by comparing their benefits and limitations.
+Readers appreciated how proactive summaries could often provide a useful initial overview of unfamiliar content.
+Preferences for proactive summaries varied with reading habits.
+For example, infrequent readers like W6 and W10 found these proactive summaries helpful in deciding whether to read the content at all.
+On the other hand, W10, who is a more frequent reader, found the constant summary updates unnecessary.
+In contrast, reader-driven question answering allowed for a more focused engagement on specific topics.
+For instance, W1 enjoyed “rapid-fire” questioning with a physics textbook, noting that it encouraged them to ask more questions than they would with traditional ChatGPT.
+W2 valued how questioning was woven into reading, preserving thought flow without the disruption of toggling between reading and chat or search interfaces.
+Overall, participants favored the reader-driven mode due to its applicability to diverse situations, yet many appreciated having both options for different purposes—proactive features for implicit skimming and reader-driven for addressing more specific questions.
+Figure 4. Snapshots captured from the Vision Pro during our diary study.
+4.5. Opportunities for the Third Iteration
+Overall, the design of the second prototype enhanced its capacity for real-world use.
+Compared to the earlier HoloLens 2 prototype, the Vision Pro’s superior rendering quality and broader field of view provided a better reader experience.
+Additionally, the simple system design greatly improved usability despite the absence of document tracking.
+While RS-Wild showed great potential, we also identified several feature-level opportunities for improvement.
+Sustained History and Recall
+Participants reported that the minimalist design effectively reduced visual clutter.
+However, it also limited the system’s overall utility.
+Although readers appreciated the simple interface, many envisioned options to save or revisit previously-generated responses.
+Several participants suggested that a history feature or the ability to store cards for future reference would substantially enhance the reader experience.
+Enhanced Screen Viewing for Extended Use
+Participants highlighted the potential for improving the experience of viewing phone and computer screens through the Vision Pro, particularly for longer viewing periods.
+While the Vision Pro’s video passthrough worked well for short-term tasks like quickly reviewing an assignment, participants envisioned improvements that could better support extended work sessions, especially when interacting with computing devices.
+Mirroring the reader’s screen(s) visually in high resolution through the headset would further align the system with the demands of realistic, work-related tasks in everyday contexts.
+Our in-the-wild study was also inherently limited by its short-term, public use — with each session lasting only one hour per participant.
+As a result, participants did not have the opportunity to fully engage with the system in their personal workflows.
+For instance, due to privacy concerns related to screen recording, participants rarely used our system for more personalized content, such as emails, messages, or calendars.
+Furthermore, since we encouraged participants to explore multiple scenarios, they tended to experiment broadly rather than focusing on any specific task in depth.
+While the in-the-wild study provided valuable insights into real-world use cases, it only captured short-term interactions in public spaces.
+There is still much to learn about the sustained, regular use of reading assistants in daily life, particularly in how readers interact with information over time, revisit previous queries, or support ongoing tasks.
+5. RS-Diary: Diary Study with the Third Prototype
+To explore more realistic longer-term use cases, we developed a third prototype variant (RS-Diary) incorporating feedback from the in-the-wild study.
+This version (Figure 4) introduced the ability to create, retain, and organize multiple information cards, along with support for mirroring content from external devices directly to the Vision Pro.
+5.1. Implementation
+RS-Diary builds on the earlier Vision Pro implementation with two key improvements: 1) Readers can now add multiple draggable cards on-demand by either making a new query or clicking an “add” button (which spawns a new summary card based on the document content currently in front of them).
+They can also spatially anchor these cards anywhere in their immediate environment and can re-select existing cards to update them with new prompts or text content.
+These cards are stored as JSON objects in a real-time database.
+2) The system also allows readers to mirror their device screen to the Vision Pro for high-resolution viewing.
+This functionality—implemented through a WebRTC video call or via Mac Virtual Display on the Vision Pro—lets readers view their computer or tablet screens at full resolution in the MR environment and also allows RS-Diary to more accurately capture and process on-screen content.
+This increased fidelity allows readers to perform a wide variety of everyday reading, coding, browsing, and watching activities for extended periods without the legibility and eye strain experienced with video passthrough.
+5.2. Study Method
+Four authors and collaborators (AA, AM, AJ, AW, and CT  [^2] The first, fourth, fifth, and sixth authors of this submission and one external acknowledged collaborator.
+The second letter of each code corresponds to the author/collaborator’s first initial.) used the system over the course of 3 weeks, integrating it into their day-to-day activities and accumulating 30 total hours of use.
+Due to physical fatigue and battery limitations of the Vision Pro, each session lasted less than 2 hours.
+The participants independently documented their use cases, insights, frustrations, thoughts, and reflections while using RS-Diary.
+During the study, each participant recorded their observations via a smartphone audio recorder and a personal reflections diary for notes, following a think-aloud process.
+Additionally, the first author conducted reflective debrief conversations with the other four participants at the end of the study to gain deeper insights into their experiences.
+We then conducted a reflexive thematic analysis (Clarke and Braun, 2014) using transcripts of these discussions, from which we identified and refined key observations.
+5.3. Results
+Throughout the diary study, participants used the system in various settings, including labs, kitchens, outdoor areas, and offices, though the majority of usage took place at desks in office environments.
+They engaged with a wide range of reading materials, such as academic papers, news articles, code, emails, calendars, Google Scholar, classroom assignments, lecture slides, textbooks, and YouTube videos.
+Participants quickly adapted the system to fit a variety of personal, academic, and technical tasks.
+For example, AW used the assistant not only to summarize email content but also to gather higher-level insights, such as identifying how many emails were sent by a particular sender.
+AM utilized the system to support programming assignments, where they used the system to break down complex coding problems into manageable steps and provided detailed explanations of specific code snippets.
+AM also leveraged the system to create learning aids, such as generating to-do lists and quiz questions based on problem descriptions.
+The ability to spatially organize tasks and concepts allowed him to balance his workload, maintain focus, and dynamically track his progress.
+Additionally, in his language learning efforts, AM used the assistant to generate contextual vocabulary lists from Japanese reading materials, which helped him focus on the most relevant and frequently used words.
+CT employed it to extract a step-by-step sequence of keyboard commands from an article on rebooting a MacBook.
+Overall, the diary study highlighted the system’s flexibility in supporting diverse tasks, from email management and coding to technical troubleshooting, language learning, and academic research.
+5.4. RS-Diary Observations
+From these experiences, participants shared various thoughts, frustrations, and potential future improvements.
+Our diary study also provided insight into the use of an always-on reading assistant in more realistic work-oriented settings.
+Quick Summary Responses Meet Many Immediate Needs
+Participants valued the system’s ability to deliver quick, concise responses, which were often sufficient for immediate clarifications.
+AM, for instance, appreciated how the system provided brief answers when they needed a rapid comparison of philosophical concepts, like distinguishing between act and rule utilitarianism.
+In many cases, just a high-level summary of the page in view was enough to satisfy their needs.
+AW found this method to be a useful shortcut, as pressing the virtual “add” button was faster and less disruptive than speaking and often provided exactly the information they were seeking.
+These simple, proactive summaries were particularly beneficial for tasks that required quick comprehension and minimal interruption.
+Using Generated Cards as Peripheral References
+Participants highlighted the advantages of cards that could exist in the periphery of their current workspace.
+For instance, AM used the system as a reading aid while learning Japanese.
+Instead of frequently opening a dictionary, which disrupted the reading flow, they generated vocabulary flashcards around the document.
+This provided language support without interrupting their reading and learning process.
+AW also noted that they would often cluster and group these generated cards thematically to the sides of the display for quick reference and then delete them once they were no longer needed.
+Similarly, AA used the system to “pin” generated code versions and key concepts around their workspace, allowing for easy access without breaking focus and helping them quickly revisit important points as needed.
+Spatial Organization for Recall and Multitasking
+Perhaps the biggest emergent pattern in the study was how all participants leveraged the system’s spatial organization features, similar to arranging post-it notes on a board.
+Both AA and AM utilized this capability to group related results and summaries together.
+For example, AM clustered cards on CPU architecture, placing related topics like buses and registers nearby, much like organizing post-its by theme.
+This spatial arrangement enabled quick access and improved recall, as readers could visually track where they had placed specific information.
+Similarly, AA pinned cards documenting different problem-solving methods across their workspace, ensuring easy access when needed.
+These organizational structures supported memory retention and task management by allowing readers to arrange information in a way that aligned with their thought processes.
+6. Lessons Learned from Three Studies
+Drawing on results from across our prototyping and evaluations, we close by showcasing the unique benefits of combining MR and AI tools, enumerating potential use cases, and highlighting opportunities for future research.
+6.1. Key Insights from the Three Studies
+Minimal Context Switching Enhances Focus and Flow
+Participants across all three studies appreciated the integration of LLM into the MR interfaces, in part, because it helped them minimize context switching, allowing them to maintain greater focus.
+Several participants reported that the system’s automatic text extraction feature helped them stay in a flow state.
+For example, AM noted that the interface allowed them to concentrate on their work without the mental disruption of switching between tasks or manually inputting context, even when using a computer screen.
+In particular, members of the author team in the RS-Diary study remarked on how the seamless workflow kept them engaged, enhancing productivity and reducing cognitive load.
+AA highlighted how the system functioned like a “teacher” by offering real-time assistance without requiring them to step away from their tasks.
+Participants also remarked that on-demand support boosted their efficiency during complex tasks like coding or writing, where focus is critical.
+A key benefit of this workflow is its support for implicit inputs, which dramatically lower the barrier to asking questions of LLMs.
 This automatic, continuous content capture broadens the range of applications beyond what participants initially envisioned.
@@
 Even for common reading tasks, they appreciated that the summary was automatically generated and updated in the background as they flipped through pages in a library.
-Overall, the always-on, implicit assistant made it effortless to integrate AI into daily tasks, expanding potential use cases and drastically reducing the explicit effort required from users.
-Unique Affordances of Spatial and Tangible Interactions
-Participants appreciated the unique affordances offered by mixed reality interfaces, such as the opportunities for spatial and tangible exploration.
-In a library setting, for example, participants could naturally navigate through bookshelves (spatial exploration), physically pick up a book (tangible interaction), and browse through its pages to automatically receive summaries (always-on assistance).
-W5 and W10 even leveraged our system to provide summaries or recommendations for entire collections of books on a bookshelf, not just content from a single page.
-Furthermore, W1 demonstrated the capability to synthesize information across multiple books on a table, requesting summaries or connections between various books and pages.
-Such observations highlight how mixed reality interfaces enable spatial and tangible interactions that go beyond the limitations of traditional screen-based interfaces, which can only capture and summarize visible content.
-Unique Ways of Using AI-Generated Content
-We observed several unique and unexpected use cases that took advantage of the LLM’s capacity for free-form questioning and answering.
-For instance, when reading a physics textbook, W8 asked the system to generate ten quizzes to help him understand quantum mechanics, transforming the textbook into a more active and personalized learning tool.
-Other common usages included translations (W4, W5, W6), learning aids (W4, W8), and personalized content recommendations (W5, W10).
-For example, W8 applied the system to his handwritten calculus notes to find answers to particular questions.
-Beyond traditional reading applications, the AI’s capabilities supported practical decision-making; W4 compared nutritional information of products on store shelves and menu items to determine healthier options, while W5 used it for guidance on waste disposal and printer operations.
-The LLM’s ability to process a broad spectrum of inquiries clearly expanded the system’s applicability and utility.
-Proactive Summaries vs On-Demand Question Answering
-Our system offers both 1) proactive assistance through automatic summary generation and 2) on-demand assistance where users manually ask questions.
-Participants shared their insights on these two functionalities by comparing the benefits and limitations.
-Proactive summaries were appreciated for providing a useful initial overview of unfamiliar topics.
-Preferences for proactive summaries varied with reading habits.
-For example, infrequent readers like W6 and W10 found these proactive summaries helpful in deciding whether to read the content or understand it without reading the full text.
-On the other hand, W10, who is a more frequent reader, found the constant summary updates unnecessary.
-In contrast, on-demand question answering allowed for a more focused engagement on specific topics.
-For instance, W1 enjoyed the “rapid-fire” questioning with a physics textbook, asking more questions than with traditional ChatGPT saying ”You have to describe it properly you have to specifically say what you’re asking about what you’re reading, so that kind of all of this abstracted in this system is like frees up your mind to just focus on the curiosity part Like rapid-fire questions and get rapid-fire answers”.
-W2 valued how questioning was woven into reading, preserving thought flow without the disruption of toggling between reading and ChatGPT interfaces.
-Overall, participants favored the on-demand feature due to its applicability to diverse situations, yet many appreciated having both options for different purposes: proactive features for implicit skimming and on-demand for addressing more specific questions.
+Overall, the always-on, implicit assistant made it effortless to switch context and maintain focus.
 Potential Risks of AI-Driven Augmented Reading
 Participants also expressed several concerns about potential risks, notably privacy issues due to the possibility of constant surveillance.
-Despite this concern, most participants were willing to continue using the system, if they could control when it was active.
+Despite this concern, most participants expressed interest in continue to use these kinds of systems, provided they could control when it was active.
 For example, W1 expressed the desire to capture all text throughout her entire life, allowing for an extensive searchable archive of everything she has encountered.
-Additionally, participants also mentioned the risk of decreased motivation and skill development.
+Such use cases, however, also pose social privacy questions—particularly when systems might record other individuals or their content.
+Additionally, participants mentioned the risk of decreased motivation and skill development.
 For example, W6 mentioned that relying too heavily on the system for translations could reduce the motivation to learn new languages.
@@
 Trust of AI-Generated Content
-Many participants approached the system’s responses with caution, often verifying the answers with the reading material.
-They highlighted the necessity of consulting additional sources for in-depth or technical readings when depending on the system.
-There was a general agreement among participants about their reluctance to fully trust an always-on reading assistant in their daily activities.
-Conversely, we noted some intriguing discussions regarding how participants perceive our system, either as a mere tool or more like a personal companion.
+Recognizing the known shortcomings or current LLMs, participants also approached the system’s responses with caution, often verifying the answers by reading the source materials and highlighting the necessity of consulting additional sources.
+There was a general agreement among participants in all studies that they would not fully trust the system’s responses to be complete or correct.
+Conversely, we noted some intriguing discussions regarding how participants perceived our system, either as a mere tool or more like a personal companion.
 While most participants engaged with the system as a conventional chatbot, some regarded it more as a personal companion than a mere tool.
@@
 The system’s ability to provide always-on, immediate responses facilitates this personification.
-5.5. Limitations
-Overall, the system design of the second prototype enhanced its applicability and stability for real-world use.
-Compared to the first prototype on the Hololens 2, the Vision Pro’s superior rendering quality and broader field of view provided a better user experience.
-Additionally, the simple system design greatly improved usability despite the absence of document tracking.
-While the second prototype showed great potential, we also identified several feature-level limitations, along with inherent study-related constraints.
-Lack of History and Recall Support
-While the minimalist design effectively reduces visual clutter and conserves screen space, it also limits the system’s overall utility.
-Although users appreciated the clean and simple interface, many expressed frustration with the lack of options to save or revisit previously generated responses.
-Several participants suggested that a history feature or the ability to store responses for future reference would substantially enhance the user experience.
-Challenges with Screen Viewing for Extended Use
-Many participants raised concerns about the difficulty of clearly viewing phone and computer screens through the Vision Pro without dramatically enlarging the text.
-While this worked for short-term tasks like quickly reviewing an assignment, participants found it frustrating for longer viewing periods, particularly when using laptops.
-Since much of their work is done on computer screens, this limitation poses a challenge for assessing the system’s effectiveness in supporting realistic, work-related tasks.
-Need for a Longer-Term Study
-Our in-the-wild study was inherently limited by its short-term, public use — with each session lasting only one hour per participant.
-As a result, participants did not have the opportunity to fully engage with the system in their personal workflows.
-For instance, due to privacy concerns related to screen recording, participants rarely used our system for more personalized content, such as emails, messages, or calendars.
-Furthermore, since participants were encouraged to explore various scenarios, they tended to experiment broadly rather than focusing on any specific task in depth.
-While the in-the-wild study provided valuable insights into real-world use cases, it only captured short-term interactions in public spaces.
-There is still much to learn about the sustained, regular use of reading assistants in daily life, particularly in how users interact with information over time, revisit previous queries, or support ongoing tasks.
-6. Study 3 RS-Diary: Diary Study with the Third Prototype
-To explore the long-term use cases, we developed a third iteration, incorporating feedback from the in-the-wild study.
-This version introduced a multiple responses feature and added the ability to view or mirror the content on an external device, addressing the readability issues for the long-term use of the Vision Pro.
-Figure 9. Snapshots captured from the Vision Pro during our diary study.
-6.1. Implementation
-The final version of RealitySummary builds on the earlier Vision Pro implementation with two key improvements: 1) Users can now add multiple draggable cards on-demand by either making a new query or clicking an “add” button.
-User can spatially anchor these cards anywhere in their immediate environment and can also re-select cards to update them with new prompts or text from the scene.
-These cards are stored as JSON objects in a real-time database.
-2) The system allows users to mirror their device screen within the Vision Pro for high-resolution viewing.
-This functionality — implemented through a WebRTC video call or via Mac Virtual Display on the Vision Pro — lets users view their external devices at full resolution in the MR environment and also allows RealitySummary to more accurately capture and process on-screen content.
-6.2. Method
-Five authors (A1-A5) used the system over the course of 3 weeks, integrating it into their day-to-day activities and accumulating 30 total hours of use.
-Due to physical fatigue and battery limitations of the Vision Pro, each session lasted less than 2 hours.
-The authors independently documented their use cases, insights, frustrations, thoughts, and reflections while using RealitySummary, which we later thematically coded.
-During the study, each participating author recorded their observations via a phone microphone, following a think-aloud process.
-Additionally, the first author conducted follow-up interviews to gain deeper insights into their experiences.
-6.3. Results
-Throughout the diary study, participants used the system in various settings, including labs, kitchens, outdoor areas, and offices, though the majority of usage took place at desks in work environments.
-They engaged with a wide range of reading materials, such as academic papers, news articles, code, emails, calendars, Google Scholar, classroom assignments, lecture slides, textbooks, and YouTube videos.
-Participants quickly adapted the system to fit a variety of personal, academic, and technical tasks.
-For example, A4 used the assistant not only to summarize email content but also to gather higher-level insights, such as identifying how many emails were sent by a particular sender.
-A2 utilized the system to support programming assignments, where they used the system to break down complex coding problems into manageable steps and provided detailed explanations of specific code snippets.
-A2 also leveraged the system for learning aids, such as generating to-do lists and quiz questions based on problem descriptions.
-The ability to spatially organize tasks and concepts allowed him to effectively manage his workload, maintain focus, and dynamically track his progress.
-Additionally, in his language learning efforts, A2 used the assistant to generate contextual vocabulary lists from Japanese reading materials, which helped him focus on the most relevant and frequently used words.
-A5 employed it to extract a step-by-step sequence of keyboard commands from an article on rebooting a MacBook.
-Overall, the diary study highlights the system’s flexibility in supporting diverse tasks, from email management and coding to technical troubleshooting, language learning, and academic research.
-6.4. Insights
-From these experiences, participants shared various insights, frustrations, and potential future improvements for the current prototype.
-Our diary study also provided insight into the use of an always-on reading assistant in more realistic work-oriented settings.
-Quick and Simple Summary Responses Meet Many Immediate Needs
-Participants valued the system’s ability to deliver quick, concise responses, which were often sufficient for immediate clarifications.
-A2, for instance, appreciated how the system provided brief answers when they needed a rapid comparison of philosophical concepts, like distinguishing between act and rule utilitarianism.
-In many cases, just a high-level summary captured through manual input was enough to satisfy their needs.
-A4 found this method to be a useful shortcut, as pressing a button was faster and less disruptive than speaking and often provided exactly the information they were seeking.
-These simple, proactive summaries were particularly beneficial for tasks that required quick comprehension and minimal interruption.
-Minimal Context Switching Enhances Focus and Flow
-Participants appreciated the integration of LLM into the MR interfaces, in part, because it helped them minimize context switching, allowing them to maintain greater focus.
-The system’s automatic text extraction feature enabled participants to stay in a flow state.
-For example, A2 noted that the interface allowed them to concentrate on their work without the mental disruption of switching between tasks or manually inputting context, even when using a computer screen.
-This seamless workflow kept participants engaged, enhancing productivity and reducing cognitive load.
-A1 highlighted how the system functioned like a “teacher” by offering real-time assistance without requiring them to step away from their tasks.
-Participants also remarked that on-demand support boosted their efficiency during complex tasks like coding or writing, where focus is critical.
-Using Generated Cards as Peripheral References
-Participants highlighted the advantages of peripheral output modalities.
-For instance, A2 used the system as a reading aid while learning Japanese.
-Instead of frequently opening a dictionary, which disrupted the reading flow, they generated vocabulary flashcards around the document.
-This provided the needed language support without interrupting the reading and learning process.
-A4 also noted that they would often cluster and group these generated cards thematically to the sides of the display for quick reference and then delete them once they were no longer needed.
-Similarly, A1 used the system to “pin” generated code versions and key concepts around their workspace, allowing for easy access without breaking focus and enabling them to quickly revisit important points as needed.
-Spatial Organization Improves Recall and Multitasking
-A key finding from the study was how participants leveraged the system’s spatial organization features, similar to arranging post-it notes on a board.
-Both A1 and A2 utilized this capability to group related concepts together.
-For example, A2 clustered cards on CPU architecture, placing related topics like buses and registers nearby, much like organizing post-its by theme.
-This spatial arrangement enabled quick access and improved recall, as users could visually track where specific information was placed.
-Similarly, A1 pinned cards documenting different problem-solving methods across their workspace, ensuring easy access when needed.
-This post-it-like organization supported memory retention and task management by allowing users to arrange information in a way that aligned with their thought processes.
-6.5. Limitations
-Challenges Controlling What to Capture
-Currently, the system captures the user’s entire field of view, but participants expressed a desire for more control over what is captured.
-For example, A1 struggled to get the system to ignore personal emails displayed next to the article he was reading, leading to summaries that inadvertently included private information.
-Similarly, A2 wanted to focus on a specific code snippet to obtain more accurate results while debugging.
-Additionally, the lack of visual feedback left users uncertain about what OCR content or context was being processed in each request.
-To address these issues, the system could offer enhanced visual feedback and allow users to control specific areas of the screen, such as through gaze interaction, which would help mitigate privacy concerns and prevent the inclusion of irrelevant or personal content.
-Cumbersome Re-Prompting Process for Common and Repeated Tasks
-Participants frequently found themselves re-prompting similar queries for routine tasks, such as requesting bullet point summaries or filtering by specific criteria.
-A potential improvement could be the addition of dynamic buttons that save commonly used prompt types, allowing users to easily reuse them later.
-Additionally, participants expressed frustration with testing out lots of different prompts to generate a complex output, such as aggregating the Google Scholar page into a timeline.
-Saving or sharing the same specific instructions for repeated tasks might mitigate these problems.
-Managing Growing Numbers of Summaries
-In the first prototype, participants found it distracting to have multiple summary cards anchored to the image in their workspace.
-However, they did not find this frustrating when they had control over the placement and could easily delete the cards.
-Additionally, participants noted that the system could sometimes generate an overwhelming number of cards.
-For instance, A2 mentioned that while analyzing a poem, the system created a card for each sentence and its meaning, making it difficult to manage information overload.
-This feedback suggests that there is potential value in exploring new interactive and semi-automated methods for efficiently sorting, snapping, organizing, filtering, and removing these cards.
-Such improvements could enhance the usability of the system and help users better manage the content generated.
-7. Lessons Learned from Three Studies
-Drawing on results from across our prototyping and evaluations, we close by showcasing the unique benefits of combining MR and AI tools, enumerating potential use cases, and highlighting opportunities for future research.
 6.2. Unique Benefits of Combining MR and AI
@@
 This always-on input allows for hands-free exploration in both spatial and temporal contexts, such as scanning entire bookshelves as a query or utilizing the history of flipped pages.
-Implicit input reduces cognitive load and minimizes context switching, as users are not required to actively specify what information should be used for the LLM’s context.
+Implicit input reduces cognitive load and minimizes context switching, as readers are not required to actively specify what information should be used for the LLM’s context.
 Peripheral and Background Output
-We found that MR’s spatial output allows information to be presented in a peripheral, background manner.
-In contrast, other output modalities, such as audio or screen-based displays, often require the user’s foreground attention to consume generated content, disrupting focus and interrupting workflow.
+We found that MR’s spatial output allowed information to be presented in a peripheral, background manner.
+In contrast, other output modalities, such as audio or screen-based visuals, often require the reader’s foreground attention to consume generated content, disrupting focus and interrupting workflows.
 With MR and peripheral outputs, however, AI-generated content becomes less intrusive, as it can be easily ignored if not needed, simply by choosing not to look at it.
@@
 We believe this capability would promote more calm human-AI interactions in the future, reducing disruptions to focused attention by combining implicit input with unobtrusive background support.
-7.2. Emergent Use Cases for Mixed Reality Text Augmentation
-Both the in-the-wild and diary studies revealed a variety of interesting emergent use cases.
-To provide a more comprehensive view of how users engage with these systems, we have categorized these into seven groups:
+6.3. Emergent Use Cases
+Both the in-the-wild and diary studies revealed a variety of interesting emergent use cases for MR text augmentation.
+To give a more comprehensive view of how readers might use these systems, we identify seven distinct approaches:
 Summarizing Content
@@
 Extracting Information
-Participants also utilized the system to extract key information, which can be translated into a more consumable format.
-This aligns with design elicitation results, which suggest tables of contents, keyword lists, timelines, and comparison tables.
+Participants also utilized the system to extract key information, which can be translated into more consumable formats like tables of contents, keyword lists, timelines, and comparison tables.
 For example, one participant extracted vocabulary flashcards to assist with foreign language reading, while another participant created a table of keyboard shortcut commands from a troubleshooting guide to use as a reference.
-In these examples, the consumable format, such as tables or vocabulary lists, serves as helpful references for their tasks.
+In these examples, the consumable format, such as a table or vocabulary list, serves as helpful references for their tasks.
 Querying External Sources
 Participants frequently used extracted text as a query to retrieve additional information from external sources.
-Echoing the information types identified in our elicitation study, participants queried specific names or keywords from their reading materials.
-For instance, one participant retrieved details about a player, team, and standings from a sports-related article, while another fact-checked claims on-demand when reading news articles.
+Participants often queried specific names or keywords from their reading materials.
+For instance, one participant retrieved details about players, teams, and standings from a sports-related article, while another fact-checked claims on-demand when reading news articles.
 Asking for Suggestions
-Beyond the above use cases identified in the elicitation study, participants frequently asked for recommendations or suggestions based on the current context.
-For instance, users asked for book recommendations when looking at a bookshelf.
-Beyond the reading context, one participant asked for advice on which trash bin can be used based on disposal instructions, while others inquired about healthier menu options or products.
-They also asked for recommendations from external sources.
-For instance, participants asked for suggestions for relevant research papers when reading academic articles.
+Participants frequently asked for recommendations or suggestions based on the current context—for instance, asking for book recommendations when looking at a bookshelf.
+Outside the context of reading, one participant asked for advice on which trash bin to use based on the disposal instructions on an item, while others inquired about healthier menu options or products.
+They also asked for recommendations from external sources, including asking for related research papers when reading academic articles.
 Aggregating Information
@@
 For example, some participants used the system to count emails based on specific filters, while others filtered and counted first-author papers from a researcher’s Google Scholar page.
-This allowed users to perform quick, approximate queries that would normally require more complex searches.
+This allowed readers to perform quick, approximate queries that would normally require more complex searches.
 Transforming Content
@@
 One participant asked the system to create reminder cards that remained spatially persistent.
-Participants also use our system to translate from one language to another.
+Participants also used our system to translate from one language to another.
 Unlike information extraction, these tasks involved converting content into actionable, interactive outputs.
@@
 Others used the system to create code based on instructions or to debug existing code snippets.
-These emergent use cases highlight novel ways of interacting with AI-generated information, allowing users to engage with individual snippets and convert them into dynamic outputs.
-They also illustrate how participants adapted the assistant to handle different types of content, which suggests a broad space of possible designs for such systems.
-7.3. Future Opportunities and Design Implications
-As both MR and LLM technologies continue to evolve rapidly, considerable opportunities exist for tools that more richly integrate them.
-Based on our three studies, we identify several key opportunities for future research and design.
-Enabling Greater Control Over Capture
-While the system’s implicit input offers significant benefits, it also raises privacy concerns due to the automatic capturing without the user’s constant awareness.
-To address this, future designs should give users control over what data is captured.
-As noted in the diary study, participants expressed a desire for more transparency and control over the captured content.
-One approach could involve providing visual feedback, allowing users to see and select the specific areas or information being captured.
-Future systems should balance the benefits of implicit input with privacy considerations by offering user control and visual feedback.
-Improving Organization and Filtering of Interactive Cards
-Participants expressed the need for more efficient ways to manage the growing number of spatially organized content cards.
-While some diary study participants manually arranged these cards around documents, others suggested more advanced organizational tools.
-Features like sorting, snapping, filtering, or automatically organizing cards could simplify workflows and reduce clutter.
-Such capabilities would be particularly beneficial for long-term tasks, where the accumulation of content can become overwhelming.
-Integrating these organizational tools would enhance usability, supporting users in managing dynamic and task-specific information over time, which enables better workspace management for daily tasks.
-Balancing Between Reusable Prompts and On-Demand Custom Queries
-Our studies highlighted the importance of accommodating both reusable prompts and custom queries.
-Each has its advantages: prepared prompts can generate complex outputs such as timelines or comparison tables, while custom queries offer users flexibility in generating content.
-The diary study participants expressed a desire for dynamic interfaces that could quickly reuse and adapt prompts to different contexts.
-For example, some suggested having customizable buttons that automatically generate summaries or lists in a consistent style.
-Others mentioned the need for a more efficient way to reuse queries across tasks, such as repeatedly generating lists or comparisons for class topics.
-This feedback highlights the importance of designing adaptive interfaces that allow users to seamlessly reuse and modify their prompts, improving efficiency and user satisfaction.
-Facilitating Seamless Transitions Between Screen and Space
-Participants also valued the ability to drag content cards from the system into their computer, such as pasting generated answers directly into coding tasks.
-Some participants also envisioned this feature being useful for writing tasks, allowing for a smooth transfer of generated content.
-Future systems should explore leveraging cross-device interactions between MR and other devices to facilitate seamless transitions, enhancing the workflow between screen-based and spatial interactions.
-Expanding Beyond Text Input to Capture Images and Everyday Objects
-Participants expressed a strong desire for the system to recognize not only text but also images.
-For example, the in-the-wild study revealed the limitations of inquiring about a diagram that the system could not recognize.
-Users also anticipated broader applications beyond text recognition, such as generating recipes based on ingredients in the fridge or providing information on real-world objects.
-Enhancing the system with multi-modal LLM capabilities could address these needs, expanding its utility beyond text-based interactions.
-Personalization Through Always-On History
-Lastly, participants showed interest in an “always-on” history feature that would allow them to recall previous interactions with the system, creating a more personalized experience.
-They speculated that such a system could eventually serve as a personalized assistant, continuously capturing and understanding context through an always-on camera.
-Although current hardware limitations, such as battery life and physical fatigue, prevent continuous long-term use, we believe future MR interfaces will seamlessly integrate with AI to provide both context-aware and personalized assistance based on users’ daily activity histories.
 7. Conclusion
-RealitySummary is the first on-demand MR reading assistant that combines OCR and LLM technologies to generate spatially-situated on-demand reading aids for both physical and digital documents.
-Through our empirical studies, we aimed to highlight new possibilities for on-demand, AI-powered reading support.
-This paper presented an iterative development of the system, along with key findings and lessons learned from three studies.
-We hope our work will inspire the HCI community to further explore the potential of integrating MR and AI in future applications.
+RealitySummary is the first exploration of on-demand MR reading assistants that combines OCR and LLM to generate spatially-situated on-demand reading aids for both physical and digital documents.
+Through our empirical studies, we highlighted new possibilities for on-demand, AI-powered reading support.
+This paper presented an iterative development of the system, along with key findings and lessons learned from three iterations and studies: Study 1 surfaced usability concerns and reader perceptions in controlled settings, Study 2 revealed challenges and opportunities of in-the-wild deployment, and Study 3 shed light on longer-term adaptation through a diary study.
+Together, these studies provide a more complete picture of how MR reading assistants can complement everyday reading.
+At the same time, our work faces limitations.
+The participant pool was small, with participants being university students.
+The in-the-wild deployment was relatively short in duration, and system performance was constrained by the practical limitations of the Apple Vision Pro, including its form factor and battery life.
+Additionally, Study 3 relied on author-participants, which may introduce bias; however, our reflective diary study approach provided valuable long-term insights that would have been difficult to obtain otherwise.
+These limitations point to opportunities for future work: evaluating the system with larger and more diverse populations, comparing always-on versus on-demand support, and extending deployments to longer-term contexts.
+We hope our work will inspire the HCI community to further explore the potential of integrating MR and AI in everyday reading applications.
+Acknowledgements.
+This research was partially funded by the NSERC Discovery Grant RGPIN-2021-02857, JST PRESTO Grant Number JPMJPR23I5, the Canada Research Chairs Program, and an Adobe Collaborative Research Gift.
+We also thank all of our study participants.
+We would like to acknowledge Tafreed Ahmad for helping with the studies and giving us constructive feedback.